### v6

# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

Abstract
Training logistic regression over encrypted data has been a compelling approach in addressing security concerns for several years. In this paper, we introduce a more efficient gradient variant, called quadratic gradient, for privacy-preserving logistic regression training. We enhance Nesterov’s Accelerated Gradient (NAG) and Adaptive Gradient Algorithm (Adagrad) by incorporating their quadratic gradients and evaluate these improved algorithms on various datasets. Experiments show that the enhanced algorithms have state-of-the-art performance in terms of convergence speed when compared to their raw first-order gradient methods. Also, we employ the enhanced NAG method to implement homomorphic logistic regression training, achieving comparable results in only 4 iterations. There is a promising possibility that quadratic gradient could combine first-order gradient descent/ascent algorithms with the second-order Newton-Raphson method, making it applicable to a wide range of numerical optimization problems.
1 Introduction
1.1 Background
Given a person’s healthcare data related to a certain disease, we can train a logistic regression (LR) model capable of telling whether or not this person is likely to develop this disease. However, such personal health information is highly private to individuals. The privacy concern, therefore, becomes a major obstacle for individuals to share their biomedical data. The most secure solution is to encrypt the data into ciphertexts first by Homomorphic Encryption (HE) and then securely outsource the ciphertexts to the cloud, without allowing the cloud to access the data directly. iDASH is an annual competition that aims to call for implementing interesting cryptographic schemes in a biological context. Since 2014, iDASH has included the theme of genomics and biomedical privacy. The third track of the 2017 iDASH competition and the second track of the 2018 iDASH competition were both to develop homomorphic-encryption-based solutions for building an LR model over encrypted data.
1.2 Related work
Several studies on logistic regression models are based on homomorphic encryption. Kim et al. [14] discussed the problem of performing LR training in an encrypted environment. They used the full batch gradient descent in the training process and the least-squares method to get the approximation of the sigmoid function. In the iDASH 2017 competition, Bonte and Vercauteren [4], Kim et al. [11], Chen et al. [5], and Crawford et al. [7] all investigated the same problem that Kim et al. [14] studied. In the iDASH competition of 2018, Kim et al. [13] and Blatt et al. [3] further worked on it for an efficient packing and semi-parallel algorithm. There are other related works [12; 2] focusing on various aspects but the papers most relevant to this work are [4] and [11]. Bonte and Vercauteren [4] developed a practical algorithm called the simplified fixed Hessian (SFH) method. Our study extends their work and adopts the ciphertext packing technique proposed by Kim et al. [11] for efficient homomorphic computation.
1.3 Contributions
Our specific contributions in this paper are as follows:
  1. 1.
    We propose a new gradient variant, quadratic gradient, which can combine the first-order gradient algorithms and the second-order Newton-Raphson method (aka Newton’s method) as one.
  2. 2.
    We develop two enhanced gradient algorithms by equipping the original ones with quadratic gradient. The resulting algorithms exhibit state-of-the-art performance in terms of convergence speed.
  3. 3.
    We implement privacy-preserving logistic regression training using the enhanced NAG method, to our best knowledge, which seems to be a good choice without compromising much on computation and storage.
2 Preliminaries
We use the square brackets “[][\ ][ ]” to denote the index of a vector or matrix element in what follows. For example, for a vector $\bm{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\bm{v}[i]$ or $\bm{v}_{[i]}$ means the $i$-th element of vector $\bm{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
2.1 Fully Homomorphic Encryption
Fully Homomorphic Encryption (FHE) is a type of cryptographic scheme that can be used to compute an arbitrary number of additions and multiplications directly on the encrypted data. It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation [8]. FHE schemes themselves are computationally time-consuming; the choice of dataset encoding matters likewise to the efficiency. In addition to these two limits, how to manage the magnitude of plaintext [10] also contributes to the slowdown. Cheon et al. [6] proposed a method to construct an HE scheme with a rescaling procedure which could eliminate this technical bottleneck effectively. We adopt their open-source implementation HEAAN while implementing our homomorphic LR algorithms. In addition, it is inevitable to pack a vector of multiple plaintexts into a single ciphertext for yielding a better amortized time of homomorphic computation. HEAAN supports a parallel technique (aka SIMD) to pack multiple complex numbers in a single polynomial and provides rotation operation on plaintext slots. The underlying HE scheme in HEAAN is well described in [11; 14; 9].
2.2 Database Encoding Method
Kim et al. [11] proposed an efficient and promising database-encoding method by using SIMD technique, which could make full use of the computation and storage resources. Suppose that a database has a training dataset consisting of $n$ samples with $(1+d)$ covariates, they packed the training dataset $Z$ into a single ciphertext in a row-by-row manner.
When employing this encoding scheme, we can manipulate the data matrix $Z$ through HE operations on the ciphertext $Enc[Z]$, utilizing only three HE operations - rotation, addition, and multiplication. For instance, if we wish to isolate the first column of $Enc[Z]$ and exclude the other columns, we can create a constant matrix $F$ with ones in the first column and zeros elsewhere. Multiplying $Enc[Z]$ by $F$ will yield the desired ciphertext.
Han et al. [9] mentioned several basic but important operations used by Kim et al. [11] in their implementation, such as a procedure named “SumColVec” to compute the summation of the columns of a matrix. With these fundamental operations, more intricate computations, like computing gradients in logistic regression models, become achievable.
2.3 Logistic Regression Model
Logistic regression is widely used in binary classification tasks to infer whether a binary-valued variable belongs to a certain class or not. LR can be generalized from linear regression [15] by mapping the whole real line $(\bm{\beta}^{\intercal}\mathbf{x})$ to $(0,1)$ via the sigmoid function $\sigma(z)=1/(1+\exp(-z))$, where the vector $\bm{\beta}\in\mathbb{R}^{(1+d)}$ is the main parameter of LR and the vector $\mathbf{x}=(1,x_{1},\ldots,x_{d})\in\mathbb{R}^{(1+d)}$ the input covariate. Thus logistic regression can be formulated with the class label $y\in\{\pm 1\}$ as follows:
$\Pr(y=+1|\mathbf{x},\bm{\beta})$ $=\sigma(\bm{\beta}^{\intercal}\mathbf{x})$ $=\frac{1}{1+e^{-\bm{\beta}^{\intercal}\mathbf{x}}},$
$\Pr(y=-1|\mathbf{x},\bm{\beta})$ $=1-\sigma(\bm{\beta}^{\intercal}\mathbf{x})$ $=\frac{1}{1+e^{+\bm{\beta}^{\intercal}\mathbf{x}}}.$
LR sets a threshold (usually $0.5$) and compares its output with it to decide the resulting class label.
The logistic regression problem can be transformed into an optimization problem that seeks a parameter $\bm{\beta}$ to maximize $L(\bm{\beta})=\prod_{i=1}^{n}\Pr(y_{i}|\mathbf{x}_{i},\bm{\beta})$ or its log-likelihood function $l(\bm{\beta})$ for convenience in the calculation:
$l(\bm{\beta})=\ln L(\bm{\beta})=-\sum_{i=1}^{n}\ln(1+e^{-y_{i}\bm{\beta}^{\intercal}\mathbf{x}_{i}}),$
where $n$ is the number of examples in the training dataset. LR does not have a closed form of maximizing $l(\bm{\beta})$ and two main methods are adopted to estimate the parameters of an LR model: (a) gradient descent method via the gradient; and (b) Newton’s method by the Hessian matrix. The gradient and Hessian of the log-likelihood function $l(\bm{\beta})$ are given by, respectively:
$\nabla_{\bm{\beta}}l(\bm{\beta})$ $=\sum_{i}(1-\sigma(y_{i}\bm{\beta}^{\intercal}\mathbf{x}_{i}))y_{i}\mathbf{x}_{i},$
$\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$ $=\sum_{i}(y_{i}\mathbf{x}_{i})(\sigma(y_{i}\bm{\beta}^{\intercal}\mathbf{x}_{i})-1)\sigma(y_{i}\bm{\beta}^{\intercal}\mathbf{x}_{i})(y_{i}\mathbf{x}_{i})$
$=X^{\intercal}SX,$
where $S$ is a diagonal matrix with entries $S_{ii}=(\sigma(y_{i}\bm{\beta}^{\intercal}\mathbf{x}_{i})-1)\sigma(y_{i}\bm{\beta}^{\intercal}\mathbf{x}_{i})$ and $X$ the dataset.
The log-likelihood function $l(\bm{\beta})$ of LR has at most a unique global maximum [1], where its gradient is zero. Newton’s method is a second-order technique to numerically find the roots of a real-valued differentiable function, and thus can be used to solve the $\bm{\beta}$ in $\nabla_{\bm{\beta}}l(\bm{\beta})=0$ for LR.
3 Technical Details
It is quite time-consuming to compute the Hessian matrix and its inverse in Newton’s method for each iteration. One way to limit this downside is to replace the varying Hessian with a fixed matrix $\bar{H}$. This novel technique is called the fixed Hessian Newton’s method. Böhning and Lindsay [bohning1988monotonicity] have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\bm{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})-\bar{H}$ is non-negative definite. With such a fixed Hessian matrix $\bar{H}$, the iteration for Newton’s method can be simplified to:
$\bm{\beta}_{t+1}=\bm{\beta}_{t}-\bar{H}^{-1}\nabla_{\bm{\beta}}l(\bm{\beta}).$
Böhning and Lindsay also suggest the fixed matrix $\bar{H}=-\frac{1}{4}X^{\intercal}X$ is a good lower bound for the Hessian of the log-likelihood function $l(\bm{\beta})$ in LR.
3.1 Simplified Fixed Hessian
Bonte and Vercauteren [4] simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain. They replace the matrix $\bar{H}$ with a diagonal matrix $B$ whose diagonal elements are simply the sums of each row in $\bar{H}$. They also suggest a specific order of calculation to optimize the computation of $B$ more efficiently. Their new approximation $B$ of the fixed Hessian is:
$B=\left[\begin{array}[]{cccc}\sum_{i=0}^{d}\bar{h}_{0i}&0&\ldots&0\\ 0&\sum_{i=0}^{d}\bar{h}_{1i}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&\sum_{i=0}^{d}\bar{h}_{di}\\ \end{array}\right],$
where $\bar{h}_{ki}$ is the element of $\bar{H}$. This diagonal matrix $B$ is in a very simple form and can be obtained from $\bar{H}$ without much difficulty. The inverse of $B$ can even be approximated in the encrypted form by computing the inverse of each diagonal element of $B$ using an iterative Newton’s method with an appropriate initial value. Their simplified fixed Hessian method can be formulated as follows:
$\bm{\beta}_{t+1}$ $=\bm{\beta}_{t}-B^{-1}\cdot\nabla_{\bm{\beta}}l(\bm{\beta}),$
$=\bm{\beta}_{t}-\left[\begin{array}[]{cccc}b_{00}&0&\ldots&0\\ 0&b_{11}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&b_{dd}\\ \end{array}\right]\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]$
$=\bm{\beta}_{t}-\left[\begin{array}[]{c}b_{00}\cdot\nabla_{0}\\ b_{11}\cdot\nabla_{1}\\ \vdots\\ b_{dd}\cdot\nabla_{d}\\ \end{array}\right],$
where $b_{ii}$ is the reciprocal of $\sum_{i=0}^{d}\bar{h}_{0i}$ and $\nabla_{i}$ is the element of $\nabla_{\bm{\beta}}l(\bm{\beta})$.
Consider a special situation: if all the elements $b_{00},\ldots,b_{dd}$ had the same value $-\eta$ with $\eta>0$, the iterative formula of the SFH method could be given as:
$\bm{\beta}_{t+1}$ $=\bm{\beta}_{t}-(-\eta)\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\bm{\beta}_{t}+\eta\cdot\nabla_{\bm{\beta}}l(\bm{\beta}),$
which is the same as the formula of the naive gradient $ascent$ method. Such a coincidence not only helps generate the idea behind this work but also leads us to believe that there is a connection between the Hessian matrix and the learning rate of the gradient (descent) method.
We regard $B^{-1}\cdot\nabla_{i}$ as a novel enhanced gradient variant and allocate a distinct learning rate to it. As long as we ensure that this new learning rate decreases from a positive floating-point number greater than 1 (such as 2) to 1 in a bounded number of iteration steps, the fixed Hessian Newton’s method guarantees the algorithm will converge eventually.
The SFH method proposed by Bonte and Vercauteren [4] has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive. For machine learning applications, datasets are typically normalized in advance to the range [0,1], satisfying the convergence condition of the SFH method. However, in other cases, such as numerical optimization, this condition may not always hold; and (b) the simplified fixed Hessian matrix $B$, as well as the fixed Hessian matrix $\bar{H}=-\frac{1}{4}X^{\intercal}X$ , can still be singular, especially when the dataset is a high-dimensional sparse matrix, such as the MNIST datasets. We extend their work by removing these limitations so as to generalize this simplified fixed Hessian to be invertible in any case and propose a faster gradient variant, which we term quadratic gradient.
3.2 Quadratic Gradient Definition
Suppose that a differentiable scalar-valued function $F(\mathbf{x})$ has its gradient $\bm{g}$ and Hessian matrix $H$, with any matrix $\bar{H}\leq H$ in the Loewner ordering for a maximization problem as follows:
$\bm{g}=\left[\begin{array}[]{c}g_{0}\\ g_{1}\\ \vdots\\ g_{d}\\ \end{array}\right],$ $H=\left[\begin{array}[]{cccc}\nabla_{00}^{2}&\nabla_{01}^{2}&\ldots&\nabla_{0d}^{2}\\ \nabla_{10}^{2}&\nabla_{11}^{2}&\ldots&\nabla_{1d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \nabla_{d0}^{2}&\nabla_{d1}^{2}&\ldots&\nabla_{dd}^{2}\\ \end{array}\right],\quad$
$\bar{H}$ $=\left[\begin{array}[]{cccc}\bar{h}_{00}&\bar{h}_{01}&\ldots&\bar{h}_{0d}\\ \bar{h}_{10}&\bar{h}_{11}&\ldots&\bar{h}_{1d}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{h}_{d0}&\bar{h}_{d1}&\ldots&\bar{h}_{dd}\\ \end{array}\right],$
where $\nabla_{ij}^{2}=\nabla_{ji}^{2}=\frac{\partial^{2}F}{\partial x_{i}\partial x_{j}}$. We construct a new diagnoal Hessian matrix $\tilde{B}$ with each diagnoal element $\tilde{B}_{kk}$ being $-\epsilon-\sum_{i=0}^{d}|\bar{h}_{ki}|$,
$\tilde{B}=\left[\begin{array}[]{ccc}-\epsilon-\sum_{i=0}^{d}|\bar{h}_{0i}|&0&\ldots\\ 0&-\epsilon-\sum_{i=0}^{d}|\bar{h}_{1i}|&\ldots\\ \vdots&\vdots&\ddots\\ 0&0&\ldots\\ \end{array}\right],$
where $\epsilon$ is a small positive constant to avoid division by zero (usually set to $81e-$ 8).
As long as $\tilde{B}$ satisfies the convergence condition of the above fixed Hessian method, $\tilde{B}\leq H$, we can use this approximation $\tilde{B}$ of the Hessian matrix as a lower bound. Since we already assume that $\bar{H}\leq H$, it will suffice to show that $\tilde{B}\leq\bar{H}$. We prove $\tilde{B}\leq\bar{H}$ in a similar way that [4] did.
Lemma 1.
Let $A\in\mathbb{R}^{n\times n}$ be a symmetric matrix, and let $B$ be the diagonal matrix whose diagonal entries $B_{kk}=-\epsilon-\sum_{i}|A_{ki}|$ for $k=1,\ldots,n$, then $B\leq A$.
Proof.
By definition of the Loewner ordering, we have to prove the difference matrix $C=A-B$ is non-negative definite, which means that all the eigenvalues of $C$ need to be non-negative. By construction of $C$ we have that $C_{ij}=A_{ij}+\epsilon+\sum_{k=1}^{n}|A_{ik}|$ for $i=j$ and $C_{ij}=A_{ij}$ for $i\neq j$. By means of Gerschgorin’s circle theorem, we can bound every eigenvalue $\lambda$ of $C$ in the sense that $|\lambda-C_{ii}|\leq\sum_{i\neq j}|C_{ij}|$ for some index $i\in\{1,2,\ldots,n\}$. We conclude that $\lambda\geq A_{ii}+\epsilon+|A_{ii}|\geq\epsilon>0$ for all eigenvalues $\lambda$ and thus that $B\leq A$. ∎
Definition 3.1 (Quadratic Gradient).
Given such a $\tilde{B}$ above, we define the quadratic gradient as $G=\bar{B}\cdot\bm{g}$ with a new learning rate $\eta$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$, and $\eta$ should be always no less than 1 and decrease to 1 in a limited number of iteration steps. Note that $G$ is still a column vector of the same size as the gradient $\bm{g}$. To maximize the function $F(\mathbf{x})$, we can use the iterative formulas: $\mathbf{x}_{t+1}=\mathbf{x}_{t}+\eta\cdot G$, just like the naive gradient.
To minimize the function $F(x)$ is the same as to just maximize the function $-F(x)$, in which case we need to construct the $\tilde{B}$ by any good lower bound $\bar{H}$ of the Hessian $-H$ of $-F(x)$ or any good upper bound $\bar{H}$ of the Hessian $H$ of $F(x)$. We point out here that $\bar{H}$ could be the Hessian matrix $H$ itself.
Algorithm 1 The Enhanced Nesterov’s Accelerated Gradient Algorithm
[…]
2:the parameter vector $V\in\mathbb{R}^{(1+d)}$
3:Set $\bar{H}\leftarrow-\frac{1}{4}X^{\intercal}X$triangleright $\bar{H}\in\mathbb{R}^{(1+d)\times(1+d)}$
4:Set $V\leftarrow\bm{0}$, $W\leftarrow\bm{0}$, $\bar{B}\leftarrow\bm{0}$triangleright $V\in\mathbb{R}^{(1+d)}$, $W\in\mathbb{R}^{(1+d)}$, $\bar{B}\in\mathbb{R}^{(1+d)\times(1+d)}$
5:for $i:=0$ to $d$ do
6:     $\bar{B}[i][i]\leftarrow\epsilon$triangleright $\epsilon$ is a small positive constant such as $81e-$ 8
7:     for $j:=0$ to $d$ do
8:         $\bar{B}[i][i]\leftarrow\bar{B}[i][i]+|\bar{H}[i][j]|$
9:     end for
10:     $\bar{B}[i][i]\leftarrow 1/\bar{B}[i][i]$
11:end for
12:Set $alpha_{0}\leftarrow 0.01$
13:Set $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
14:for $count:=1$ to $\kappa$ do
15:     Set $Z\leftarrow\bm{0}$triangleright $Z\in\mathbb{R}^{n}$ will store the inputs for Sigmoid function
16:     for $i:=1$ to $n$ do
17:         for $j:=0$ to $d$ do
18:              $Z[i]\leftarrow Z[i]+Y[i]\times V[j]\times X[i][j]$
19:         end for
20:     end for
21:     Set $\bm{\sigma}\leftarrow\bm{0}$triangleright $\bm{\sigma}\in\mathbb{R}^{n}$ will store the outputs of Sigmoid function
22:     for $i:=1$ to $n$ do
23:         $\bm{\sigma}[i]\leftarrow 1/(1+\exp(-Z[i]))$
24:     end for
25:     Set $\bm{g}\leftarrow\bm{0}$
26:     for $j:=0$ to $d$ do
27:         for $i:=1$ to $n$ do
28:              $\bm{g}[j]\leftarrow\bm{g}[j]+(1-\bm{\sigma}[i])\times Y[i]\times X[i][j]$
29:         end for
30:     end for
31:     Set $G\leftarrow\bm{0}$
32:     for $j:=0$ to $d$ do
33:         $G[j]\leftarrow\bar{B}[j][j]\times\bm{g}[j]$
34:     end for
35:     Set $\eta\leftarrow(1-alpha_{0})/alpha_{1}$
36:     Set $\gamma\leftarrow lr/(n\times count)$triangleright $n$ is the size of training data; $lr$ is set to 10.0 in this work
37:     for $j:=0$ to $d$ do
38:         $w_{temp}\leftarrow V[j]+(1+\gamma)\times G[j]$
39:         $V[j]\leftarrow(1-\eta)\times w_{temp}+\eta\times W[j]$
40:         $W[j]\leftarrow w_{temp}$
41:     end for
42:     $alpha_{0}\leftarrow alpha_{1}$
43:     $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
44:end for
45:return $V$
3.3 Quadratic Gradient Algorithms
This gradient variant, Quadratic Gradient, can be used to enhance various first-order gradient algorithms, such as NAG and Adagrad.
NAG is a different variant of the momentum method to give the momentum term much more prescience. The iterative formulas of the gradient $ascent$ method for NAG are as follows:
$V_{t+1}$ $=\bm{\beta}_{t}+\alpha_{t}\cdot\nabla J(\bm{\beta}_{t}),$ (3)
$\bm{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
where $V_{t+1}$ is the intermediate variable used for updating the final weight $\bm{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position [11]. The enhanced NAG is to replace (3) with $V_{t+1}=\bm{\beta}_{t}+(1+\alpha_{t})\cdot G$. Our enhanced NAG method is described in Algorithm 1 .
Adagrad is a gradient-based algorithm suitable for dealing with sparse data. The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\bm{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
$\bm{\beta}_{[i]}^{(t+1)}$ $=\bm{\beta}_{[i]}^{(t)}-\frac{\eta}{\epsilon+\sqrt{\sum_{k=1}^{t}\bm{g}_{[i]}^{(t)}\cdot\bm{g}_{[i]}^{(t)}}}\cdot\bm{g}_{[i]}^{(t)},$
$\bm{\beta}_{[i]}^{(t+1)}$ $=\bm{\beta}_{[i]}^{(t)}-\frac{1+\eta}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
(a) The iDASH dataset
(b) The Edinburgh dataset
(c) The lbw dataset
(d) The nhanes3 dataset
(e) The pcs dataset
(f) The uis dataset
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 1: The training results of Adagrad and Enhanced Adagrad in the clear setting
Performance Evaluation We assess the performance of various algorithms in a clear setting using the Python programming language on a desktop computer with an Intel Core CPU G640 at 1.60 GHz and 7.3 GB RAM. Since our focus is on how fast the algorithms converge in the training phase, the loss function, maximum likelihood estimation (MLE), is selected as the only indicator. We evaluate four algorithms, NAG, Adagrad, and their quadratic-gradient versions (denoted as Enhanced NAG and Enhanced Adagrad, respectively) on the datasets that Kim et al. [11] adopted: the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (Edinburgh), Low Birth weight Study (lbw), Nhanes III (nhanes3), Prostate Cancer study (pcs), and Umaru Impact Study datasets (uis). The genomic dataset is provided by the third task in the iDASH competition of 2017, which consists of 1579 records. Each record has 103 binary genotypes and a binary phenotype indicating if the patient has cancer. The other five datasets all have a single binary dependent variable. We also evaluate these algorithms on two large datasets from Han et al. [9]: a real financial data consisting of 422,108 samples over 200 features and the restructured public MNIST dataset that consists of 11,982 samples of the training dataset with 196 features. For a fair comparison with the baseline work [11], we use the learning rate $1+10/(1+t)$ since Kim et al. [11] select $10/(1+t)$ as their learning rate. In our experiments, we always use $\bar{H}$ = $-\frac{1}{4}X^{\intercal}X$ to construct our $\tilde{B}$ for the binary LR model.
Figures 1 and 2 show that except for the enhanced Adagrad method on the iDASH genomic dataset our enhanced methods all converge faster than their original ones in other cases. In all the Python experiments, the time to calculate the $\bar{B}$ in quadratic gradient $G$ before running the iterations and the time to run each iteration for various algorithms are not important and can be considered negligible. (few seconds).
An important observation is that the enhanced algorithms have a better performance when using learning rates between $1$ and $2$ than using learning rate otherelse. When using a learning rate greater than $3$, the algorithm is almost certain not to converge. When employing quadratic gradient algorithms, we recommend adopting the learning rate setting $1+A\cdot\gamma^{t}$ as the optimal choice, where $t$ represents the iteration number, $A$ is a positive float number typically set to no less than $1$, and $\gamma$ is usually a positive number less than 1, controlling the rate of decay. We strongly recommend that $A$ is set to $1.0$ and $\gamma$ is tuned to search for optimal results, or that one can set $A$ to a small number less than $1.0$ and $\gamma$ to a larger number greater than $0.9$.
(a) The iDASH dataset
(b) The Edinburgh dataset
(c) The lbw dataset
(d) The nhanes3 dataset
(e) The pcs dataset
(f) The uis dataset
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 2: The training results of NAG and Enhanced NAG in the clear setting
Analysis In Figure 1(a), the enhanced Adagrad algorithm failed to outperform the original Adagrad algorithm. The possible reason for that might be related to the limitations of the raw Adagrad method. Without a doubt, Adagrad is a novel algorithm initiated to accelerate each element of the gradient with different learning rates. However, Adagrad tends to converge to a suboptimal solution due to its aggressive, monotonically decreasing learning rates. This would lead to its main limitation that in the later training phase every learning rate for different components of the gradient is too close to zero due to keeping adding positive additional terms to the denominator, stopping the algorithm from learning anything.
On the other hand, the original Adagrad method has another little-noticed limitation: the learning rate in the first few iterations tends to be large. While this limitation does not affect the performance of the original Adagrad method to some extent, the enhanced Adagrad method exacerbates this phenomenon significantly, leading to the Learning-Rate Explosion . Therefore, the enhanced Adagrad cannot be applied to general optimization problems such as Rosenbrock’s function. The exploding learning rate would be too large for the algorithm to survive the first several iterations, finally leading the optimization function to some point where its output cannot be represented by the computer system. This might explain why the performance of this algorithm in all cases, not just on the iDASH genome dataset, seems to be meaningless and unstable in the first few iterations.
Several improved algorithms upon the Adagrad method, such as RMSProp, have been proposed in order to address these issues existed, via using an exponential moving average of historical gradients rather than just the sum of all squared gradients from the beginning of training. We might be able to overcome the problems existing in the enhanced Adagrad method by adopting the enhanced Adagrad-like variants, like the enhanced Adadelta method and the enhanced RMSProp method.
The reasons why the enhanced algorithms have a better performance might be that quadratic gradient can incorporate curves into first-order gradient algorithms.
(a) The raw first-order gradient ascent algorithm with iterative formulas: $\bm{\beta}_{t+1}=\bm{\beta}_{t}+\mathrm{lr}\cdot\bm{g}$
(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\bm{\beta}_{t+1}=\bm{\beta}_{t}+\mathrm{LR}\cdot G$
Figure 3: The training outcomes for both the raw gradient algorithm and the raw quadratic gradient ascent algorithm in the clear setting, conducted on the lbw dataset
Gradient And Quadratic Gradient We executed the raw gradient ascent algorithm and raw quadratic gradient algorithm on the lbw dataset using various learning rates. Figure 3 shows the detailed results of this experiment. It is precisely these results that inspired the present authors to come up with the idea of quadratic gradient. Similar to the gradient, the quadratic gradient carrying the curve also exhibits the characteristic of smooth progression. That is, when the learning rate is gradually increased, both the gradient and the quadratic gradient will show corresponding gradual changes in performance rather than abrupt shifts.
4 Homomorphic Training
Adagrad method is not a practical solution for homomorphic LR due to its frequent inversion operations. It seems plausible that the enhanced NAG is likely the reasonable choice for homomorphic LR training. We adopt the enhanced NAG method to implement secure logistic regression training based on HE. The difficulty in applying the quadratic gradient is to invert the diagonal matrix $\tilde{B}$ in order to obtain $\bar{B}$. We leave the computation of matrix $\bar{B}$ to data owner and let the data owner upload the ciphertext encrypting the $\bar{B}$ to the cloud. Since data owner has to prepare the dataset and normalize it, it would also be practicable for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model. A common solution is to replace the sigmoid function with a polynomial approximation by using the widely adopted least-squares method. We can call a function named “ polyfit($\cdot$) ” in the Python package Numpy to fit the polynomial in a least-square sense. We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al. [11], utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al. [11] used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$. The weight vector $\beta^{(0)}$ consisting of zeros and the diagnoal elements of $\bar{B}$ are copied $n$ times to form two matrices. The data owner then encrypt the two matrices into two ciphertexts $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$, respectively. The ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ are as follows:
X $=\left[\begin{array}[]{cccc}1&x_{11}&\ldots&x_{1d}\\ 1&x_{21}&\ldots&x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n1}&\ldots&x_{nd}\\ \end{array}\right],\text{Y}=\left[\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\\ \end{array}\right],$
$\text{ct}_{Z}$ $=Enc\left[\begin{array}[]{cccc}y_{1}&y_{1}x_{11}&\ldots&y_{1}x_{1d}\\ y_{2}&y_{2}x_{21}&\ldots&y_{2}x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ y_{n}&y_{n}x_{n1}&\ldots&y_{n}x_{nd}\\ \end{array}\right],$
$\text{ct}_{\beta}^{(0)}$ $=Enc\left[\begin{array}[]{cccc}\beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \vdots&\vdots&\ddots&\vdots\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \end{array}\right],$
$\text{ct}_{\bar{B}}$ $=Enc\left[\begin{array}[]{cccc}\bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \end{array}\right],$
where $\bar{B}_{[i][i]}$ is the diagonal element of $\bar{B}$ that is built from $-\frac{1}{4}X^{\intercal}X$.
The pulbic cloud takes the three ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ and evaluates the enhanced NAG algorithm to find a decent weight vector by updating the vector $\text{ct}_{\beta}^{(0)}$. Refer to [11] for a detailed description about how to calculate the gradient by HE programming.
Limitations
In a privacy-preserving setting, when compared to the NAG method, the primary limitation of the Enhanced NAG method is that it requires one additional ciphertext multiplication to construct the quadratic gradient. In addition, the data owner needs to upload one more ciphertext encrypting the matrix $\bar{B}$. However, the enhanced algorithm converges faster, and we believe it can compensate for the mentioned drawbacks.
Table 1: Implementation Results for iDASH datasets with 10-fold CV
Dataset
Sample
Num
Feature
Num
Method
Iter
Num
Storage
Learn
Time
Accuracy
AUC
iDASH 1579 18 [11] 7 0.04 6.07 62.87 0.689
Ours 4 0.08 4.43 61.46 0.696
Table 2: Implementation Results for other datasets with 5-fold CV
Dataset
Sample
Num
Feature
Num
Method
Iter
Num
Storage
Learn
Time
Accuracy
AUC
Edinburgh 1253 9 [11] 7 0.02 3.6 91.04 0.958
Ours 4 0.04 0.6 89.52 0.943
lbw 189 9 [11] 7 0.02 3.3 69.19 0.689
Ours 4 0.04 0.6 71.35 0.667
nhanes3 15649 15 [11] 7 0.16 7.3 79.22 0.717
Ours 4 0.31 4.5 79.23 0.637
pcs 379 9 [11] 7 0.02 3.5 68.27 0.740
Ours 4 0.04 0.6 63.20 0.733
uis 575 8 [11] 7 0.02 3.5 74.44 0.603
Ours 4 0.04 0.6 74.43 0.597
5 Experiments
Implementation
We implement the enhanced NAG based on HE with the library HEAAN. The C++ source code is publicly available at https://anonymous.4open.science/r/IDASH2017-245B . All the experiments on the ciphertexts were conducted on a public cloud with 32 vCPUs and 64 GB RAM.
For a fair comparison with [11], we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets. Like [11], We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators. Tables 1 and 2 display the results of the two experiments, respectively. The two tables also provide the average evaluation running time for each iteration and the storage (encrypted dataset for the baseline work and encrypted dataset and $\bar{B}$ for our method). We adopt the same packing method that Kim et al. [11] proposed and hence our solution has similar storage of ciphertexts to [11] with some extra ciphertexts to encrypt the $\bar{B}$. We chose $1+0.9^{t}$ as our learning rate configuration.
The parameters of HEAAN we set are same to [11]: $logN=16$, $logQ=1200$, $logp=30$, $slots=32768$, which ensure the security level $\lambda=80$. We use a larger $logp=40$ to encrypt the matrix ${\bar{B}}$ for preserving the precision of ${\bar{B}}$. Refer to [11] for details on these parameters. Since our enhanced NAG method need one more ciphertext multiplication than the baseline work, consuming more modulus, our solution thus can only perform $4$ iterations of the enhanced NAG method. Yet despite only $4$ iterations, our enhanced NAG method still produces a comparable result.
6 Conclusion
In this paper, we proposed a faster gradient variant called quadratic gradient, and implemented the quadratic-gradient version of NAG in the encrypted domain to train the logistic regression model.
The quadratic gradient presented in this work can be constructed from the Hessian matrix directly, and thus somehow combines the first-order Newton’s method and the second-order gradient (descent) method together. There is a promising chance that quadratic gradient could accelerate other gradient methods, which is an open future work.
Also, quadratic gradient might substitute and supersede the line-search method, for example when using enhanced Adagrad-like methods, and could use gradient descent methods to accelerate Newton’s method, resulting in super-quadratic algorithms.

### v7

@@
 There is a promising possibility that quadratic gradient could combine first-order gradient descent/ascent algorithms with the second-order Newton-Raphson method, making it applicable to a wide range of numerical optimization problems.
+Homomorphic Encryption, Logistic Regression, Quadratic Gradient, Simplified Fixed Hessian, Nesterov’s Accelerated Gradient, Gradient Ascent Algorithm
 1 Introduction
-1.1 Background
 Given a person’s healthcare data related to a certain disease, we can train a logistic regression (LR) model capable of telling whether or not this person is likely to develop this disease.
@@
 The third track of the 2017 iDASH competition and the second track of the 2018 iDASH competition were both to develop homomorphic-encryption-based solutions for building an LR model over encrypted data.
-1.2 Related work
 Several studies on logistic regression models are based on homomorphic encryption.
 Kim et al.
-[14] discussed the problem of performing LR training in an encrypted environment.
+(2018b) discussed the problem of performing LR training in an encrypted environment.
 They used the full batch gradient descent in the training process and the least-squares method to get the approximation of the sigmoid function.
-In the iDASH 2017 competition, Bonte and Vercauteren [4], Kim et al.
-[11], Chen et al.
-[5], and Crawford et al.
-[7] all investigated the same problem that Kim et al.
-[14] studied.
+In the iDASH 2017 competition, Bonte and Vercauteren (2018), Kim et al.
+(2018a), Chen et al.
+(2018), and Crawford et al.
+(2018) all investigated the same problem that Kim et al.
+(2018b) studied.
 In the iDASH competition of 2018, Kim et al.
-[13] and Blatt et al.
-[3] further worked on it for an efficient packing and semi-parallel algorithm.
-There are other related works [12; 2] focusing on various aspects but the papers most relevant to this work are [4] and [11].
-Bonte and Vercauteren [4] developed a practical algorithm called the simplified fixed Hessian (SFH) method.
+(2019b) and Blatt et al.
+(2019) further worked on it for an efficient packing and semi-parallel algorithm.
+There are other related works (Kim et al., 2019a; Bergamaschi et al., 2019) focusing on various aspects but the papers most relevant to this work are (Bonte & Vercauteren, 2018) and (Kim et al., 2018a).
+Bonte and Vercauteren (2018) developed a practical algorithm called the simplified fixed Hessian (SFH) method.
 Our study extends their work and adopts the ciphertext packing technique proposed by Kim et al.
-[11] for efficient homomorphic computation.
-1.3 Contributions
+(2018a) for efficient homomorphic computation.
 Our specific contributions in this paper are as follows:
@@
 We use the square brackets “[][\ ][ ]” to denote the index of a vector or matrix element in what follows.
-For example, for a vector $\bm{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\bm{v}[i]$ or $\bm{v}_{[i]}$ means the $i$-th element of vector $\bm{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
+For example, for a vector $\boldsymbol{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\boldsymbol{v}[i]$ or $\boldsymbol{v}_{[i]}$ means the $i$-th element of vector $\boldsymbol{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
 2.1 Fully Homomorphic Encryption
 Fully Homomorphic Encryption (FHE) is a type of cryptographic scheme that can be used to compute an arbitrary number of additions and multiplications directly on the encrypted data.
-It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation [8].
+It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation (2009).
 FHE schemes themselves are computationally time-consuming; the choice of dataset encoding matters likewise to the efficiency.
-In addition to these two limits, how to manage the magnitude of plaintext [10] also contributes to the slowdown.
+In addition to these two limits, how to manage the magnitude of plaintext (Jäschke & Armknecht, 2016) also contributes to the slowdown.
 Cheon et al.
-[6] proposed a method to construct an HE scheme with a rescaling procedure which could eliminate this technical bottleneck effectively.
+(2017) proposed a method to construct an HE scheme with a rescaling procedure which could eliminate this technical bottleneck effectively.
 We adopt their open-source implementation HEAAN while implementing our homomorphic LR algorithms.
@@
 HEAAN supports a parallel technique (aka SIMD) to pack multiple complex numbers in a single polynomial and provides rotation operation on plaintext slots.
-The underlying HE scheme in HEAAN is well described in [11; 14; 9].
+The underlying HE scheme in HEAAN is well described in (Kim et al., 2018a, b; Han et al., 2019).
 2.2 Database Encoding Method
 Kim et al.
-[11] proposed an efficient and promising database-encoding method by using SIMD technique, which could make full use of the computation and storage resources.
+(2018a) proposed an efficient and promising database-encoding method by using SIMD technique, which could make full use of the computation and storage resources.
 Suppose that a database has a training dataset consisting of $n$ samples with $(1+d)$ covariates, they packed the training dataset $Z$ into a single ciphertext in a row-by-row manner.
@@
 Han et al.
-[9] mentioned several basic but important operations used by Kim et al.
-[11] in their implementation, such as a procedure named “SumColVec” to compute the summation of the columns of a matrix.
+(2019) mentioned several basic but important operations used by Kim et al.
+(2018a) in their implementation, such as a procedure named “SumColVec” to compute the summation of the columns of a matrix.
 With these fundamental operations, more intricate computations, like computing gradients in logistic regression models, become achievable.
@@
 Logistic regression is widely used in binary classification tasks to infer whether a binary-valued variable belongs to a certain class or not.
-LR can be generalized from linear regression [15] by mapping the whole real line $(\bm{\beta}^{\intercal}\mathbf{x})$ to $(0,1)$ via the sigmoid function $\sigma(z)=1/(1+\exp(-z))$, where the vector $\bm{\beta}\in\mathbb{R}^{(1+d)}$ is the main parameter of LR and the vector $\mathbf{x}=(1,x_{1},\ldots,x_{d})\in\mathbb{R}^{(1+d)}$ the input covariate.
+LR can be generalized from linear regression (Murphy, 2012) by mapping the whole real line $(\boldsymbol{\beta}^{\top}\mathbf{x})$ to $(0,1)$ via the sigmoid function $\sigma(z)=1/(1+\exp(-z))$, where the vector $\boldsymbol{\beta}\in\mathbb{R}^{(1+d)}$ is the main parameter of LR and the vector $\mathbf{x}=(1,x_{1},\ldots,x_{d})\in\mathbb{R}^{(1+d)}$ the input covariate.
 Thus logistic regression can be formulated with the class label $y\in\{\pm 1\}$ as follows:
-$\Pr(y=+1|\mathbf{x},\bm{\beta})$ $=\sigma(\bm{\beta}^{\intercal}\mathbf{x})$ $=\frac{1}{1+e^{-\bm{\beta}^{\intercal}\mathbf{x}}},$
-$\Pr(y=-1|\mathbf{x},\bm{\beta})$ $=1-\sigma(\bm{\beta}^{\intercal}\mathbf{x})$ $=\frac{1}{1+e^{+\bm{\beta}^{\intercal}\mathbf{x}}}.$
+$\Pr(y=+1|\mathbf{x},\boldsymbol{\beta})$ $=\sigma(\boldsymbol{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{-\boldsymbol{\beta}^{\top}\mathbf{x}}},$
+$\Pr(y=-1|\mathbf{x},\boldsymbol{\beta})$ $=1-\sigma(\boldsymbol{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{+\boldsymbol{\beta}^{\top}\mathbf{x}}}.$
 LR sets a threshold (usually $0.5$) and compares its output with it to decide the resulting class label.
-The logistic regression problem can be transformed into an optimization problem that seeks a parameter $\bm{\beta}$ to maximize $L(\bm{\beta})=\prod_{i=1}^{n}\Pr(y_{i}|\mathbf{x}_{i},\bm{\beta})$ or its log-likelihood function $l(\bm{\beta})$ for convenience in the calculation:
-$l(\bm{\beta})=\ln L(\bm{\beta})=-\sum_{i=1}^{n}\ln(1+e^{-y_{i}\bm{\beta}^{\intercal}\mathbf{x}_{i}}),$
+The logistic regression problem can be transformed into an optimization problem that seeks a parameter $\boldsymbol{\beta}$ to maximize $L(\boldsymbol{\beta})=\prod_{i=1}^{n}\Pr(y_{i}|\mathbf{x}_{i},\boldsymbol{\beta})$ or its log-likelihood function $l(\boldsymbol{\beta})$ for convenience in the calculation:
+$l(\boldsymbol{\beta})=\ln L(\boldsymbol{\beta})=-\sum_{i=1}^{n}\ln(1+e^{-y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}}),$
 where $n$ is the number of examples in the training dataset.
-LR does not have a closed form of maximizing $l(\bm{\beta})$ and two main methods are adopted to estimate the parameters of an LR model: (a) gradient descent method via the gradient; and (b) Newton’s method by the Hessian matrix.
-The gradient and Hessian of the log-likelihood function $l(\bm{\beta})$ are given by, respectively:
-$\nabla_{\bm{\beta}}l(\bm{\beta})$ $=\sum_{i}(1-\sigma(y_{i}\bm{\beta}^{\intercal}\mathbf{x}_{i}))y_{i}\mathbf{x}_{i},$
-$\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$ $=\sum_{i}(y_{i}\mathbf{x}_{i})(\sigma(y_{i}\bm{\beta}^{\intercal}\mathbf{x}_{i})-1)\sigma(y_{i}\bm{\beta}^{\intercal}\mathbf{x}_{i})(y_{i}\mathbf{x}_{i})$
-$=X^{\intercal}SX,$
-where $S$ is a diagonal matrix with entries $S_{ii}=(\sigma(y_{i}\bm{\beta}^{\intercal}\mathbf{x}_{i})-1)\sigma(y_{i}\bm{\beta}^{\intercal}\mathbf{x}_{i})$ and $X$ the dataset.
-The log-likelihood function $l(\bm{\beta})$ of LR has at most a unique global maximum [1], where its gradient is zero.
-Newton’s method is a second-order technique to numerically find the roots of a real-valued differentiable function, and thus can be used to solve the $\bm{\beta}$ in $\nabla_{\bm{\beta}}l(\bm{\beta})=0$ for LR.
+LR does not have a closed form of maximizing $l(\boldsymbol{\beta})$ and two main methods are adopted to estimate the parameters of an LR model: (a) gradient descent method via the gradient; and (b) Newton’s method by the Hessian matrix.
+The gradient and Hessian of the log-likelihood function $l(\boldsymbol{\beta})$ are given by, respectively:
+$\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$ $=\sum_{i}(1-\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}))y_{i}\mathbf{x}_{i},$
+$\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$ $=\sum_{i}(y_{i}\mathbf{x}_{i})(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})(y_{i}\mathbf{x}_{i})$
+$=X^{\top}SX,$
+where $S$ is a diagonal matrix with entries $S_{ii}=(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})$ and $X$ the dataset.
+The log-likelihood function $l(\boldsymbol{\beta})$ of LR has at most a unique global maximum (Allison, 2008), where its gradient is zero.
+Newton’s method is a second-order technique to numerically find the roots of a real-valued differentiable function, and thus can be used to solve the $\boldsymbol{\beta}$ in $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})=0$ for LR.
 3 Technical Details
@@
 This novel technique is called the fixed Hessian Newton’s method.
-Böhning and Lindsay [bohning1988monotonicity] have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\bm{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})-\bar{H}$ is non-negative definite.
+Böhning and Lindsay (1988) have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\boldsymbol{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})-\bar{H}$ is non-negative definite.
 With such a fixed Hessian matrix $\bar{H}$, the iteration for Newton’s method can be simplified to:
-$\bm{\beta}_{t+1}=\bm{\beta}_{t}-\bar{H}^{-1}\nabla_{\bm{\beta}}l(\bm{\beta}).$
-Böhning and Lindsay also suggest the fixed matrix $\bar{H}=-\frac{1}{4}X^{\intercal}X$ is a good lower bound for the Hessian of the log-likelihood function $l(\bm{\beta})$ in LR.
+$\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-\bar{H}^{-1}\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}).$
+Böhning and Lindsay also suggest the fixed matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ is a good lower bound for the Hessian of the log-likelihood function $l(\boldsymbol{\beta})$ in LR.
 3.1 Simplified Fixed Hessian
-Bonte and Vercauteren [4] simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain.
+Bonte and Vercauteren (2018) simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain.
 They replace the matrix $\bar{H}$ with a diagonal matrix $B$ whose diagonal elements are simply the sums of each row in $\bar{H}$.
@@
 Their simplified fixed Hessian method can be formulated as follows:
-$\bm{\beta}_{t+1}$ $=\bm{\beta}_{t}-B^{-1}\cdot\nabla_{\bm{\beta}}l(\bm{\beta}),$
-$=\bm{\beta}_{t}-\left[\begin{array}[]{cccc}b_{00}&0&\ldots&0\\ 0&b_{11}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&b_{dd}\\ \end{array}\right]\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]$
-$=\bm{\beta}_{t}-\left[\begin{array}[]{c}b_{00}\cdot\nabla_{0}\\ b_{11}\cdot\nabla_{1}\\ \vdots\\ b_{dd}\cdot\nabla_{d}\\ \end{array}\right],$
-where $b_{ii}$ is the reciprocal of $\sum_{i=0}^{d}\bar{h}_{0i}$ and $\nabla_{i}$ is the element of $\nabla_{\bm{\beta}}l(\bm{\beta})$.
+$\boldsymbol{\beta}_{t+1}$ $=\boldsymbol{\beta}_{t}-B^{-1}\cdot\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}),$
+$=\boldsymbol{\beta}_{t}-\left[\begin{array}[]{cccc}b_{00}&0&\ldots&0\\ 0&b_{11}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&b_{dd}\\ \end{array}\right]\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]$
+$=\boldsymbol{\beta}_{t}-\left[\begin{array}[]{c}b_{00}\cdot\nabla_{0}\\ b_{11}\cdot\nabla_{1}\\ \vdots\\ b_{dd}\cdot\nabla_{d}\\ \end{array}\right],$
+where $b_{ii}$ is the reciprocal of $\sum_{i=0}^{d}\bar{h}_{0i}$ and $\nabla_{i}$ is the element of $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$.
 Consider a special situation: if all the elements $b_{00},\ldots,b_{dd}$ had the same value $-\eta$ with $\eta>0$, the iterative formula of the SFH method could be given as:
-$\bm{\beta}_{t+1}$ $=\bm{\beta}_{t}-(-\eta)\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\bm{\beta}_{t}+\eta\cdot\nabla_{\bm{\beta}}l(\bm{\beta}),$
+$\boldsymbol{\beta}_{t+1}$ $=\boldsymbol{\beta}_{t}-(-\eta)\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\boldsymbol{\beta}_{t}+\eta\cdot\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}),$
 which is the same as the formula of the naive gradient $ascent$ method.
@@
 As long as we ensure that this new learning rate decreases from a positive floating-point number greater than 1 (such as 2) to 1 in a bounded number of iteration steps, the fixed Hessian Newton’s method guarantees the algorithm will converge eventually.
-The SFH method proposed by Bonte and Vercauteren [4] has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive.
+The SFH method proposed by Bonte and Vercauteren (2018) has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive.
 For machine learning applications, datasets are typically normalized in advance to the range [0,1], satisfying the convergence condition of the SFH method.
-However, in other cases, such as numerical optimization, this condition may not always hold; and (b) the simplified fixed Hessian matrix $B$, as well as the fixed Hessian matrix $\bar{H}=-\frac{1}{4}X^{\intercal}X$ , can still be singular, especially when the dataset is a high-dimensional sparse matrix, such as the MNIST datasets.
+However, in other cases, such as numerical optimization, this condition may not always hold; and (b) the simplified fixed Hessian matrix $B$, as well as the fixed Hessian matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ , can still be singular, especially when the dataset is a high-dimensional sparse matrix, such as the MNIST datasets.
 We extend their work by removing these limitations so as to generalize this simplified fixed Hessian to be invertible in any case and propose a faster gradient variant, which we term quadratic gradient.
 3.2 Quadratic Gradient Definition
-Suppose that a differentiable scalar-valued function $F(\mathbf{x})$ has its gradient $\bm{g}$ and Hessian matrix $H$, with any matrix $\bar{H}\leq H$ in the Loewner ordering for a maximization problem as follows:
-$\bm{g}=\left[\begin{array}[]{c}g_{0}\\ g_{1}\\ \vdots\\ g_{d}\\ \end{array}\right],$ $H=\left[\begin{array}[]{cccc}\nabla_{00}^{2}&\nabla_{01}^{2}&\ldots&\nabla_{0d}^{2}\\ \nabla_{10}^{2}&\nabla_{11}^{2}&\ldots&\nabla_{1d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \nabla_{d0}^{2}&\nabla_{d1}^{2}&\ldots&\nabla_{dd}^{2}\\ \end{array}\right],\quad$
+Suppose that a differentiable scalar-valued function $F(\mathbf{x})$ has its gradient $\boldsymbol{g}$ and Hessian matrix $H$, with any matrix $\bar{H}\leq H$ in the Loewner ordering for a maximization problem as follows:
+$\boldsymbol{g}=\left[\begin{array}[]{c}g_{0}\\ g_{1}\\ \vdots\\ g_{d}\\ \end{array}\right],$ $H=\left[\begin{array}[]{cccc}\nabla_{00}^{2}&\nabla_{01}^{2}&\ldots&\nabla_{0d}^{2}\\ \nabla_{10}^{2}&\nabla_{11}^{2}&\ldots&\nabla_{1d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \nabla_{d0}^{2}&\nabla_{d1}^{2}&\ldots&\nabla_{dd}^{2}\\ \end{array}\right],\quad$
 $\bar{H}$ $=\left[\begin{array}[]{cccc}\bar{h}_{00}&\bar{h}_{01}&\ldots&\bar{h}_{0d}\\ \bar{h}_{10}&\bar{h}_{11}&\ldots&\bar{h}_{1d}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{h}_{d0}&\bar{h}_{d1}&\ldots&\bar{h}_{dd}\\ \end{array}\right],$
@@
 Since we already assume that $\bar{H}\leq H$, it will suffice to show that $\tilde{B}\leq\bar{H}$.
-We prove $\tilde{B}\leq\bar{H}$ in a similar way that [4] did.
-Lemma 1.
+We prove $\tilde{B}\leq\bar{H}$ in a similar way that (Bonte & Vercauteren, 2018) did.
+Lemma 3.1.
 Let $A\in\mathbb{R}^{n\times n}$ be a symmetric matrix, and let $B$ be the diagonal matrix whose diagonal entries $B_{kk}=-\epsilon-\sum_{i}|A_{ki}|$ for $k=1,\ldots,n$, then $B\leq A$.
@@
 Definition 3.2 (Quadratic Gradient).
-Given such a $\tilde{B}$ above, we define the quadratic gradient as $G=\bar{B}\cdot\bm{g}$ with a new learning rate $\eta$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$, and $\eta$ should be always no less than 1 and decrease to 1 in a limited number of iteration steps.
-Note that $G$ is still a column vector of the same size as the gradient $\bm{g}$.
+Given such a $\tilde{B}$ above, we define the quadratic gradient as $G=\bar{B}\cdot\boldsymbol{g}$ with a new learning rate $\eta$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$, and $\eta$ should be always no less than 1 and decrease to 1 in a limited number of iteration steps.
+Note that $G$ is still a column vector of the same size as the gradient $\boldsymbol{g}$.
 To maximize the function $F(\mathbf{x})$, we can use the iterative formulas: $\mathbf{x}_{t+1}=\mathbf{x}_{t}+\eta\cdot G$, just like the naive gradient.
@@
 Algorithm 1 The Enhanced Nesterov’s Accelerated Gradient Algorithm
-[…]
-2:the parameter vector $V\in\mathbb{R}^{(1+d)}$
-3:Set $\bar{H}\leftarrow-\frac{1}{4}X^{\intercal}X$triangleright $\bar{H}\in\mathbb{R}^{(1+d)\times(1+d)}$
-4:Set $V\leftarrow\bm{0}$, $W\leftarrow\bm{0}$, $\bar{B}\leftarrow\bm{0}$triangleright $V\in\mathbb{R}^{(1+d)}$, $W\in\mathbb{R}^{(1+d)}$, $\bar{B}\in\mathbb{R}^{(1+d)\times(1+d)}$
+1:  Require: training dataset $X\in\mathbb{R}^{n\times(1+d)}$; training label $Y\in\mathbb{R}^{n\times 1}$; learning rate $lr\in\mathbb{R}$; and the number $\kappa$ of iterations;
+2:  Ensure: the parameter vector $V\in\mathbb{R}^{(1+d)}$
+3:  Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$ $\bar{H}\in\mathbb{R}^{(1+d)\times(1+d)}$
+4:  Set $V\leftarrow\boldsymbol{0}$, $W\leftarrow\boldsymbol{0}$, $\bar{B}\leftarrow\boldsymbol{0}$ $V\in\mathbb{R}^{(1+d)}$, $W\in\mathbb{R}^{(1+d)}$, $\bar{B}\in\mathbb{R}^{(1+d)\times(1+d)}$
 5:  for $i:=0$ to $d$ do
-6:     $\bar{B}[i][i]\leftarrow\epsilon$triangleright $\epsilon$ is a small positive constant such as $81e-$ 8
-7:     for $j:=0$ to $d$ do
-8:         $\bar{B}[i][i]\leftarrow\bar{B}[i][i]+|\bar{H}[i][j]|$
-9:     end for
-10:     $\bar{B}[i][i]\leftarrow 1/\bar{B}[i][i]$
+6:     # $\epsilon$ is a small positive constant such as $81e-$ 8
+7:     $\bar{B}[i][i]\leftarrow\epsilon$
+8:     for $j:=0$ to $d$ do
+9:        $\bar{B}[i][i]\leftarrow\bar{B}[i][i]+|\bar{H}[i][j]|$
+10:     end for
 11:  end for
@@
 14:  for $count:=1$ to $\kappa$ do
-15:     Set $Z\leftarrow\bm{0}$triangleright $Z\in\mathbb{R}^{n}$ will store the inputs for Sigmoid function
-16:     for $i:=1$ to $n$ do
-17:         for $j:=0$ to $d$ do
-18:              $Z[i]\leftarrow Z[i]+Y[i]\times V[j]\times X[i][j]$
-19:         end for
+15:     # $Z\in\mathbb{R}^{n}$ will store the inputs for Sigmoid function
+16:     Set $Z\leftarrow\boldsymbol{0}$
+17:     for $i:=1$ to $n$ do
+18:        for $j:=0$ to $d$ do
+19:           $Z[i]\leftarrow Z[i]+Y[i]\times V[j]\times X[i][j]$
 20:        end for
-21:     Set $\bm{\sigma}\leftarrow\bm{0}$triangleright $\bm{\sigma}\in\mathbb{R}^{n}$ will store the outputs of Sigmoid function
-22:     for $i:=1$ to $n$ do
-23:         $\bm{\sigma}[i]\leftarrow 1/(1+\exp(-Z[i]))$
-24:     end for
-25:     Set $\bm{g}\leftarrow\bm{0}$
-26:     for $j:=0$ to $d$ do
-27:         for $i:=1$ to $n$ do
-28:              $\bm{g}[j]\leftarrow\bm{g}[j]+(1-\bm{\sigma}[i])\times Y[i]\times X[i][j]$
-29:         end for
-30:     end for
-31:     Set $G\leftarrow\bm{0}$
-32:     for $j:=0$ to $d$ do
-33:         $G[j]\leftarrow\bar{B}[j][j]\times\bm{g}[j]$
-34:     end for
-35:     Set $\eta\leftarrow(1-alpha_{0})/alpha_{1}$
-36:     Set $\gamma\leftarrow lr/(n\times count)$triangleright $n$ is the size of training data; $lr$ is set to 10.0 in this work
-37:     for $j:=0$ to $d$ do
-38:         $w_{temp}\leftarrow V[j]+(1+\gamma)\times G[j]$
-39:         $V[j]\leftarrow(1-\eta)\times w_{temp}+\eta\times W[j]$
-40:         $W[j]\leftarrow w_{temp}$
-41:     end for
-42:     $alpha_{0}\leftarrow alpha_{1}$
-43:     $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
-44:end for
-45:return $V$
+21:     end for
+22:     # $\boldsymbol{\sigma}\in\mathbb{R}^{n}$ will store the outputs of Sigmoid function
+23:     Set $\boldsymbol{\sigma}\leftarrow\boldsymbol{0}$
+24:     for $i:=1$ to $n$ do
+25:        $\boldsymbol{\sigma}[i]\leftarrow 1/(1+\exp(-Z[i]))$
+26:     end for
+27:     Set $\boldsymbol{g}\leftarrow\boldsymbol{0}$
+28:     for $j:=0$ to $d$ do
+29:        for $i:=1$ to $n$ do
+30:           $\boldsymbol{g}[j]\leftarrow\boldsymbol{g}[j]+(1-\boldsymbol{\sigma}[i])\times Y[i]\times X[i][j]$
+31:        end for
+32:     end for
+33:     Set $G\leftarrow\boldsymbol{0}$
+34:     for $j:=0$ to $d$ do
+35:        $G[j]\leftarrow\bar{B}[j][j]\times\boldsymbol{g}[j]$
+36:     end for
+37:     Set $\eta\leftarrow(1-alpha_{0})/alpha_{1}$
+38:     Set $\gamma\leftarrow lr/(n\times count)$
+39:     for $j:=0$ to $d$ do
+40:        $w_{temp}\leftarrow V[j]+(1+\gamma)\times G[j]$
+41:        $V[j]\leftarrow(1-\eta)\times w_{temp}+\eta\times W[j]$
+42:        $W[j]\leftarrow w_{temp}$
+43:     end for
+44:     $alpha_{0}\leftarrow alpha_{1}$
+45:     $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
+46:  end for
+47:  Return $V$
 3.3 Quadratic Gradient Algorithms
-This gradient variant, Quadratic Gradient, can be used to enhance various first-order gradient algorithms, such as NAG and Adagrad.
+This gradient variant, Quadratic Gradient, can be used to enhance various first-order gradient algorithms:
+  1. 1.
     NAG is a different variant of the momentum method to give the momentum term much more prescience. The iterative formulas of the gradient $ascent$ method for NAG are as follows:
-$V_{t+1}$ $=\bm{\beta}_{t}+\alpha_{t}\cdot\nabla J(\bm{\beta}_{t}),$ (3)
-$\bm{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
-where $V_{t+1}$ is the intermediate variable used for updating the final weight $\bm{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position [11].
-The enhanced NAG is to replace (3) with $V_{t+1}=\bm{\beta}_{t}+(1+\alpha_{t})\cdot G$.
-Our enhanced NAG method is described in Algorithm 1 .
+    $V_{t+1}$ $=\boldsymbol{\beta}_{t}+\eta_{t}\cdot\nabla J(\boldsymbol{\beta}_{t}),$ (3)
+    $\boldsymbol{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
+    where $V_{t+1}$ is the intermediate variable used for updating the final weight $\boldsymbol{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position (Kim et al., 2018a).
+The enhanced NAG is to replace (3) with $V_{t+1}=\boldsymbol{\beta}_{t}+N_{t}\cdot G,$ where $N_{t}$ is the new learning rate for the enhanced algorithm usually setted to $1+\eta_{t}$.
+Our enhanced NAG is described in Algorithm 1 .
+  2. 2.
     Adagrad is a gradient-based algorithm suitable for dealing with sparse data.
-The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\bm{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
-$\bm{\beta}_{[i]}^{(t+1)}$ $=\bm{\beta}_{[i]}^{(t)}-\frac{\eta}{\epsilon+\sqrt{\sum_{k=1}^{t}\bm{g}_{[i]}^{(t)}\cdot\bm{g}_{[i]}^{(t)}}}\cdot\bm{g}_{[i]}^{(t)},$
-$\bm{\beta}_{[i]}^{(t+1)}$ $=\bm{\beta}_{[i]}^{(t)}-\frac{1+\eta}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
+The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\boldsymbol{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
+    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{\eta_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}\boldsymbol{g}_{[i]}^{(t)}\cdot\boldsymbol{g}_{[i]}^{(t)}}}\cdot\boldsymbol{g}_{[i]}^{(t)},$
+    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{N_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
+  3. 3.
+    Adam is an optimization algorithm that combines the benefits of Adagrad, which adapts learning rates for each parameter, with momentum, allowing it to converge faster and handle sparse gradients more effectively.
+Similar to how Adagrad transitions to its enhanced versions, the update formula for enhanced Adam, also employs a new learning rate and replaces the gradient with its quadratic form.
 (a) The iDASH dataset
@@
 (h) The private financial dataset
-Figure 1: The training results of Adagrad and Enhanced Adagrad in the clear setting
-Performance Evaluation We assess the performance of various algorithms in a clear setting using the Python programming language on a desktop computer with an Intel Core CPU G640 at 1.60 GHz and 7.3 GB RAM.
-Since our focus is on how fast the algorithms converge in the training phase, the loss function, maximum likelihood estimation (MLE), is selected as the only indicator.
-We evaluate four algorithms, NAG, Adagrad, and their quadratic-gradient versions (denoted as Enhanced NAG and Enhanced Adagrad, respectively) on the datasets that Kim et al.
-[11] adopted: the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (Edinburgh), Low Birth weight Study (lbw), Nhanes III (nhanes3), Prostate Cancer study (pcs), and Umaru Impact Study datasets (uis).
-The genomic dataset is provided by the third task in the iDASH competition of 2017, which consists of 1579 records.
-Each record has 103 binary genotypes and a binary phenotype indicating if the patient has cancer.
-The other five datasets all have a single binary dependent variable.
-We also evaluate these algorithms on two large datasets from Han et al.
-[9]: a real financial data consisting of 422,108 samples over 200 features and the restructured public MNIST dataset that consists of 11,982 samples of the training dataset with 196 features.
-For a fair comparison with the baseline work [11], we use the learning rate $1+10/(1+t)$ since Kim et al.
-[11] select $10/(1+t)$ as their learning rate.
-In our experiments, we always use $\bar{H}$ = $-\frac{1}{4}X^{\intercal}X$ to construct our $\tilde{B}$ for the binary LR model.
-Figures 1 and 2 show that except for the enhanced Adagrad method on the iDASH genomic dataset our enhanced methods all converge faster than their original ones in other cases.
-In all the Python experiments, the time to calculate the $\bar{B}$ in quadratic gradient $G$ before running the iterations and the time to run each iteration for various algorithms are not important and can be considered negligible.
-(few seconds).
-An important observation is that the enhanced algorithms have a better performance when using learning rates between $1$ and $2$ than using learning rate otherelse.
-When using a learning rate greater than $3$, the algorithm is almost certain not to converge.
-When employing quadratic gradient algorithms, we recommend adopting the learning rate setting $1+A\cdot\gamma^{t}$ as the optimal choice, where $t$ represents the iteration number, $A$ is a positive float number typically set to no less than $1$, and $\gamma$ is usually a positive number less than 1, controlling the rate of decay.
-We strongly recommend that $A$ is set to $1.0$ and $\gamma$ is tuned to search for optimal results, or that one can set $A$ to a small number less than $1.0$ and $\gamma$ to a larger number greater than $0.9$.
+Figure 1: The training results of NAG and Enhanced NAG in the clear
 (a) The iDASH dataset
@@
 (h) The private financial dataset
-Figure 2: The training results of NAG and Enhanced NAG in the clear setting
+Figure 2: The training results of Adagrad and Enhanced Adagrad in the clear
+(a) The iDASH dataset
+(b) The Edinburgh dataset
+(c) The lbw dataset
+(d) The nhanes3 dataset
+(e) The pcs dataset
+(f) The uis dataset
+(g) restructured MNIST dataset
+(h) The private financial dataset
+Figure 3: The training results of Adam and Enhanced Adam in the clear
+Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM.
+Given that our focus is on the convergence speed of the algorithms during training, we use the loss function, specifically maximum likelihood estimation (MLE), as the only indicator.
+We evaluate six algorithms—NAG, Adagrad, Adam, and their quadratic-gradient variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam, respectively)—using the datasets adopted by Kim et al.
+(2018a): the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (Edinburgh), the Low Birth Weight Study (lbw), Nhanes III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study datasets (uis).
+The genomic dataset from the third task of the 2017 iDASH competition comprises 1,579 records, each featuring 103 binary genotypes and a binary phenotype indicating whether the patient has cancer.
+The remaining five datasets each contain a single binary dependent variable.
+We also evaluate these algorithms on two large datasets from Han et al.
+(2019): a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features.
+For a fair comparison with the baseline work (Kim et al., 2018a), we use the learning rate $1+\frac{10}{1+t}$, in line with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate.
+In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the binary logistic regression model.
+Figures 21, and 3 show that, except for the Enhanced Adagrad and Adam methods on the iDASH genomic dataset, our enhanced methods converge faster than their original counterparts in all other cases.
+All figures demonstrate that the enhanced algorithms converge empirically.
+In all the Python experiments, the time required to calculate $\bar{B}$ in the quadratic gradient $G$ before running the iterations, as well as the time to execute each iteration for various algorithms, is minimal and can be considered negligible (a few seconds).
+An important observation is that the enhanced algorithms have a better performance when using learning rates between $1$ and $2$ than using learning rate otherelse.
+When using a learning rate greater than $3$, the algorithm is almost certain not to converge.
+When employing quadratic gradient algorithms, we recommend adopting the learning rate setting $1+A\cdot\gamma^{t}$ as the optimal choice, where $t$ represents the iteration number, $A$ is a positive float number typically set to no less than $1$, and $\gamma$ is usually a positive number less than 1, controlling the rate of decay.
+We strongly recommend that $A$ is set to $1.0$ and $\gamma$ is tuned to search for optimal results, or that one can set $A$ to a small number less than $1.0$ and $\gamma$ to a larger number greater than $0.9$.
 Analysis In Figure 2(a), the enhanced Adagrad algorithm failed to outperform the original Adagrad algorithm.
@@
 The reasons why the enhanced algorithms have a better performance might be that quadratic gradient can incorporate curves into first-order gradient algorithms.
-(a) The raw first-order gradient ascent algorithm with iterative formulas: $\bm{\beta}_{t+1}=\bm{\beta}_{t}+\mathrm{lr}\cdot\bm{g}$
-(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\bm{\beta}_{t+1}=\bm{\beta}_{t}+\mathrm{LR}\cdot G$
+(a) The raw first-order gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{lr}\cdot\boldsymbol{g}$
+(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{LR}\cdot G$
 Figure 4: The training outcomes for both the raw gradient algorithm and the raw quadratic gradient ascent algorithm in the clear setting, conducted on the lbw dataset
@@
 We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al.
-[11], utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
+(2018a), utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
 Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al.
-[11] used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$.
+(2018a) used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$.
 The weight vector $\beta^{(0)}$ consisting of zeros and the diagnoal elements of $\bar{B}$ are copied $n$ times to form two matrices.
@@
 $\text{ct}_{\bar{B}}$ $=Enc\left[\begin{array}[]{cccc}\bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \end{array}\right],$
-where $\bar{B}_{[i][i]}$ is the diagonal element of $\bar{B}$ that is built from $-\frac{1}{4}X^{\intercal}X$.
+where $\bar{B}_{[i][i]}$ is the diagonal element of $\bar{B}$ that is built from $-\frac{1}{4}X^{\top}X$.
 The pulbic cloud takes the three ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ and evaluates the enhanced NAG algorithm to find a decent weight vector by updating the vector $\text{ct}_{\beta}^{(0)}$.
-Refer to [11] for a detailed description about how to calculate the gradient by HE programming.
+Refer to (Kim et al., 2018a) for a detailed description about how to calculate the gradient by HE programming.
 Limitations
@@
 Num
-Method
+Method deg $g$
 Iter
@@
 AUC
-iDASH 1579 18 [11] 7 0.04 6.07 62.87 0.689
-Ours 4 0.08 4.43 61.46 0.696
+iDASH 1579 18 Ours 5 4 0.08 4.43 61.46 0.696
+(Kim et al., 2018a) 5 7 0.04 6.07 62.87 0.689
 Table 2: Implementation Results for other datasets with 5-fold CV
@@
 Num
-Method
+Method deg $g$
 Iter
@@
 AUC
-Edinburgh 1253 9 [11] 7 0.02 3.6 91.04 0.958
-Ours 4 0.04 0.6 89.52 0.943
-lbw 189 9 [11] 7 0.02 3.3 69.19 0.689
-Ours 4 0.04 0.6 71.35 0.667
-nhanes3 15649 15 [11] 7 0.16 7.3 79.22 0.717
-Ours 4 0.31 4.5 79.23 0.637
-pcs 379 9 [11] 7 0.02 3.5 68.27 0.740
-Ours 4 0.04 0.6 63.20 0.733
-uis 575 8 [11] 7 0.02 3.5 74.44 0.603
-Ours 4 0.04 0.6 74.43 0.597
+Edinburgh 1253 9 Ours 5 4 0.04 0.6 89.52 0.943
+(Kim et al., 2018a) 5 7 0.02 3.6 91.04 0.958
+lbw 189 9 Ours 5 4 0.04 0.6 71.35 0.667
+(Kim et al., 2018a) 5 7 0.02 3.3 69.19 0.689
+nhanes3 15649 15 Ours 5 4 0.31 4.5 79.23 0.637
+(Kim et al., 2018a) 5 7 0.16 7.3 79.22 0.717
+pcs 379 9 Ours 5 4 0.04 0.6 63.20 0.733
+(Kim et al., 2018a) 5 7 0.02 3.5 68.27 0.740
+uis 575 8 Ours 5 4 0.04 0.6 74.43 0.597
+(Kim et al., 2018a) 5 7 0.02 3.5 74.44 0.603
 5 Experiments
@@
 All the experiments on the ciphertexts were conducted on a public cloud with 32 vCPUs and 64 GB RAM.
-For a fair comparison with [11], we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets.
-Like [11], We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators.
+For a fair comparison with (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets.
+Like (Kim et al., 2018a), We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators.
 Tables 1 and 2 display the results of the two experiments, respectively.
@@
 We adopt the same packing method that Kim et al.
-[11] proposed and hence our solution has similar storage of ciphertexts to [11] with some extra ciphertexts to encrypt the $\bar{B}$.
+(2018a) proposed and hence our solution has similar storage of ciphertexts to (Kim et al., 2018a) with some extra ciphertexts to encrypt the $\bar{B}$.
 We chose $1+0.9^{t}$ as our learning rate configuration.
-The parameters of HEAAN we set are same to [11]: $logN=16$, $logQ=1200$, $logp=30$, $slots=32768$, which ensure the security level $\lambda=80$.
+The parameters of HEAAN we set are same to (Kim et al., 2018a): $logN=16$, $logQ=1200$, $logp=30$, $slots=32768$, which ensure the security level $\lambda=80$.
 We use a larger $logp=40$ to encrypt the matrix ${\bar{B}}$ for preserving the precision of ${\bar{B}}$.
-Refer to [11] for details on these parameters.
+Refer to (Kim et al., 2018a) for details on these parameters.
 Since our enhanced NAG method need one more ciphertext multiplication than the baseline work, consuming more modulus, our solution thus can only perform $4$ iterations of the enhanced NAG method.
@@
 Also, quadratic gradient might substitute and supersede the line-search method, for example when using enhanced Adagrad-like methods, and could use gradient descent methods to accelerate Newton’s method, resulting in super-quadratic algorithms.
+Impact Statements
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v8

@@
 Training logistic regression over encrypted data has been a compelling approach in addressing security concerns for several years.
-In this paper, we introduce a more efficient gradient variant, called quadratic gradient, for privacy-preserving logistic regression training.
-We enhance Nesterov’s Accelerated Gradient (NAG) and Adaptive Gradient Algorithm (Adagrad) by incorporating their quadratic gradients and evaluate these improved algorithms on various datasets.
-Experiments show that the enhanced algorithms have state-of-the-art performance in terms of convergence speed when compared to their raw first-order gradient methods.
-Also, we employ the enhanced NAG method to implement homomorphic logistic regression training, achieving comparable results in only 4 iterations.
-There is a promising possibility that quadratic gradient could combine first-order gradient descent/ascent algorithms with the second-order Newton-Raphson method, making it applicable to a wide range of numerical optimization problems.
-Homomorphic Encryption, Logistic Regression, Quadratic Gradient, Simplified Fixed Hessian, Nesterov’s Accelerated Gradient, Gradient Ascent Algorithm
+In this paper, we introduce an efficient gradient variant, called $quadratic$ $gradient$, for privacy-preserving logistic regression training.
+We enhance Nesterov’s Accelerated Gradient (NAG), Adaptive Gradient Algorithm (Adagrad) and Adam algorithms by incorporating their quadratic gradients and evaluate these improved algorithms on various datasets.
+Experimental results demonstrate that the enhanced algorithms achieve significantly improved convergence speed compared to traditional first-order gradient methods.
+Moreover, we applied the enhanced NAG method to implement homomorphic logistic regression training, achieving comparable results within just 4 iterations.
+There is a good chance that the quadratic gradient approach could integrate first-order gradient descent/ascent algorithms with the second-order Newton-Raphson methods, and that it could be applied to a wide range of numerical optimization problems.
 1 Introduction
@@
 Since 2014, iDASH has included the theme of genomics and biomedical privacy.
-The third track of the 2017 iDASH competition and the second track of the 2018 iDASH competition were both to develop homomorphic-encryption-based solutions for building an LR model over encrypted data.
+The third track of the 2017 iDASH competition and the second track of the 2018 iDASH competition were both to develop HE-based solutions for building an LR model over encrypted data.
 Several studies on logistic regression models are based on homomorphic encryption.
+Kim et al.
 Kim et al.
@@
 They used the full batch gradient descent in the training process and the least-squares method to get the approximation of the sigmoid function.
-In the iDASH 2017 competition, Bonte and Vercauteren (2018), Kim et al.
+In the iDASH 2017 competition, Bonte and Vercauteren Bonte & Vercauteren (2018), Kim et al.
+Kim et al.
 (2018a), Chen et al.
+Chen et al.
 (2018), and Crawford et al.
+Crawford et al.
 (2018) all investigated the same problem that Kim et al.
+Kim et al.
 (2018b) studied.
 In the iDASH competition of 2018, Kim et al.
+Kim et al.
 (2019b) and Blatt et al.
+Blatt et al.
 (2019) further worked on it for an efficient packing and semi-parallel algorithm.
-There are other related works (Kim et al., 2019a; Bergamaschi et al., 2019) focusing on various aspects but the papers most relevant to this work are (Bonte & Vercauteren, 2018) and (Kim et al., 2018a).
-Bonte and Vercauteren (2018) developed a practical algorithm called the simplified fixed Hessian (SFH) method.
+There are other related works (Kim et al., 2019a; Bergamaschi et al., 2019; Ogilvie et al., 2020) focusing on various aspects but the papers most relevant to this work are (Bonte & Vercauteren, 2018) and (Kim et al., 2018a).
+Bonte and Vercauteren Bonte & Vercauteren (2018) developed a practical algorithm called the simplified fixed Hessian (SFH) method.
 Our study extends their work and adopts the ciphertext packing technique proposed by Kim et al.
+Kim et al.
 (2018a) for efficient homomorphic computation.
@@
   1. 1.
-    We propose a new gradient variant, quadratic gradient, which can combine the first-order gradient algorithms and the second-order Newton-Raphson method (aka Newton’s method) as one.
+    We propose a new gradient variant, quadratic gradient, which can combine the first-order gradient algorithms and the second-order Newton-Raphson method (aka Newton’s method).
   2. 2.
-    We develop two enhanced gradient algorithms by equipping the original ones with quadratic gradient. The resulting algorithms exhibit state-of-the-art performance in terms of convergence speed.
+    We develop three enhanced gradient algorithms by equipping the original ones with quadratic gradient.
+The resulting algorithms in most cases demonstrate high performance in terms of convergence speed.
   3. 3.
-    We implement privacy-preserving logistic regression training using the enhanced NAG method, to our best knowledge, which seems to be a good choice without compromising much on computation and storage.
+    We implement privacy-preserving logistic regression training using the enhanced NAG method, to our best knowledge, which seems to be a great choice without compromising much on computation and storage.
 2 Preliminaries
 We use the square brackets “[][\ ][ ]” to denote the index of a vector or matrix element in what follows.
-For example, for a vector $\boldsymbol{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\boldsymbol{v}[i]$ or $\boldsymbol{v}_{[i]}$ means the $i$-th element of vector $\boldsymbol{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
+For example, for a vector $\bm{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\bm{v}[i]$ or $\bm{v}_{[i]}$ means the $i$-th element of vector $\bm{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
 2.1 Fully Homomorphic Encryption
 Fully Homomorphic Encryption (FHE) is a type of cryptographic scheme that can be used to compute an arbitrary number of additions and multiplications directly on the encrypted data.
-It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation (2009).
+It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation (Gentry, 2009).
 FHE schemes themselves are computationally time-consuming; the choice of dataset encoding matters likewise to the efficiency.
 In addition to these two limits, how to manage the magnitude of plaintext (Jäschke & Armknecht, 2016) also contributes to the slowdown.
+Cheon et al.
 Cheon et al.
@@
 HEAAN supports a parallel technique (aka SIMD) to pack multiple complex numbers in a single polynomial and provides rotation operation on plaintext slots.
-The underlying HE scheme in HEAAN is well described in (Kim et al., 2018a, b; Han et al., 2019).
+The underlying HE scheme in HEAAN is well described in (Kim et al., 2018a; b; Han et al., 2019).
 2.2 Database Encoding Method
+Kim et al.
 Kim et al.
@@
 Han et al.
+Han et al.
 (2019) mentioned several basic but important operations used by Kim et al.
+Kim et al.
 (2018a) in their implementation, such as a procedure named “SumColVec” to compute the summation of the columns of a matrix.
@@
 Logistic regression is widely used in binary classification tasks to infer whether a binary-valued variable belongs to a certain class or not.
-LR can be generalized from linear regression (Murphy, 2012) by mapping the whole real line $(\boldsymbol{\beta}^{\top}\mathbf{x})$ to $(0,1)$ via the sigmoid function $\sigma(z)=1/(1+\exp(-z))$, where the vector $\boldsymbol{\beta}\in\mathbb{R}^{(1+d)}$ is the main parameter of LR and the vector $\mathbf{x}=(1,x_{1},\ldots,x_{d})\in\mathbb{R}^{(1+d)}$ the input covariate.
+LR can be generalized from linear regression (Murphy, 2012) by mapping the whole real line $(\bm{\beta}^{\top}\mathbf{x})$ to $(0,1)$ via the sigmoid function $\sigma(z)=1/(1+\exp(-z))$, where the vector $\bm{\beta}\in\mathbb{R}^{(1+d)}$ is the main parameter of LR and the vector $\mathbf{x}=(1,x_{1},\ldots,x_{d})\in\mathbb{R}^{(1+d)}$ the input covariate.
 Thus logistic regression can be formulated with the class label $y\in\{\pm 1\}$ as follows:
-$\Pr(y=+1|\mathbf{x},\boldsymbol{\beta})$ $=\sigma(\boldsymbol{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{-\boldsymbol{\beta}^{\top}\mathbf{x}}},$
-$\Pr(y=-1|\mathbf{x},\boldsymbol{\beta})$ $=1-\sigma(\boldsymbol{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{+\boldsymbol{\beta}^{\top}\mathbf{x}}}.$
+$\Pr(y=+1|\mathbf{x},\bm{\beta})$ $=\sigma(\bm{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{-\bm{\beta}^{\top}\mathbf{x}}},$
+$\Pr(y=-1|\mathbf{x},\bm{\beta})$ $=1-\sigma(\bm{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{+\bm{\beta}^{\top}\mathbf{x}}}.$
 LR sets a threshold (usually $0.5$) and compares its output with it to decide the resulting class label.
-The logistic regression problem can be transformed into an optimization problem that seeks a parameter $\boldsymbol{\beta}$ to maximize $L(\boldsymbol{\beta})=\prod_{i=1}^{n}\Pr(y_{i}|\mathbf{x}_{i},\boldsymbol{\beta})$ or its log-likelihood function $l(\boldsymbol{\beta})$ for convenience in the calculation:
-$l(\boldsymbol{\beta})=\ln L(\boldsymbol{\beta})=-\sum_{i=1}^{n}\ln(1+e^{-y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}}),$
+The logistic regression problem can be transformed into an optimization problem that seeks a parameter $\bm{\beta}$ to maximize $L(\bm{\beta})=\prod_{i=1}^{n}\Pr(y_{i}|\mathbf{x}_{i},\bm{\beta})$ or its log-likelihood function $l(\bm{\beta})$ for convenience in the calculation:
+$l(\bm{\beta})=\ln L(\bm{\beta})=-\sum_{i=1}^{n}\ln(1+e^{-y_{i}\bm{\beta}^{\top}\mathbf{x}_{i}}),$
 where $n$ is the number of examples in the training dataset.
-LR does not have a closed form of maximizing $l(\boldsymbol{\beta})$ and two main methods are adopted to estimate the parameters of an LR model: (a) gradient descent method via the gradient; and (b) Newton’s method by the Hessian matrix.
-The gradient and Hessian of the log-likelihood function $l(\boldsymbol{\beta})$ are given by, respectively:
-$\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$ $=\sum_{i}(1-\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}))y_{i}\mathbf{x}_{i},$
-$\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$ $=\sum_{i}(y_{i}\mathbf{x}_{i})(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})(y_{i}\mathbf{x}_{i})$
+LR does not have a closed form of maximizing $l(\bm{\beta})$ and two main methods are adopted to estimate the parameters of an LR model: (a) gradient descent method via the gradient; and (b) Newton’s method by the Hessian matrix.
+The gradient and Hessian of the log-likelihood function $l(\bm{\beta})$ are given by, respectively:
+$\nabla_{\bm{\beta}}l(\bm{\beta})$ $=\sum_{i}(1-\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i}))y_{i}\mathbf{x}_{i},$
+$\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$ $=\sum_{i}(y_{i}\mathbf{x}_{i})(\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i})(y_{i}\mathbf{x}_{i})$
 $=X^{\top}SX,$
-where $S$ is a diagonal matrix with entries $S_{ii}=(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})$ and $X$ the dataset.
-The log-likelihood function $l(\boldsymbol{\beta})$ of LR has at most a unique global maximum (Allison, 2008), where its gradient is zero.
-Newton’s method is a second-order technique to numerically find the roots of a real-valued differentiable function, and thus can be used to solve the $\boldsymbol{\beta}$ in $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})=0$ for LR.
+where $S$ is a diagonal matrix with entries $S_{ii}=(\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i})$ and $X$ the dataset.
+The log-likelihood function $l(\bm{\beta})$ of LR has at most a unique global maximum (Allison, 2008), where its gradient is zero.
+Newton’s method is a second-order technique to numerically find the roots of a real-valued differentiable function, and thus can be used to solve the $\bm{\beta}$ in $\nabla_{\bm{\beta}}l(\bm{\beta})=0$ for LR.
 3 Technical Details
@@
 This novel technique is called the fixed Hessian Newton’s method.
-Böhning and Lindsay (1988) have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\boldsymbol{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})-\bar{H}$ is non-negative definite.
+Böhning and Lindsay Böhning & Lindsay (1988) have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\bm{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})-\bar{H}$ is non-negative definite.
 With such a fixed Hessian matrix $\bar{H}$, the iteration for Newton’s method can be simplified to:
-$\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-\bar{H}^{-1}\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}).$
-Böhning and Lindsay also suggest the fixed matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ is a good lower bound for the Hessian of the log-likelihood function $l(\boldsymbol{\beta})$ in LR.
+$\bm{\beta}_{t+1}=\bm{\beta}_{t}-\bar{H}^{-1}\nabla_{\bm{\beta}}l(\bm{\beta}).$
+Böhning and Lindsay also suggest the fixed matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ is a good lower bound for the Hessian of the log-likelihood function $l(\bm{\beta})$ in LR.
 3.1 Simplified Fixed Hessian
-Bonte and Vercauteren (2018) simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain.
+Bonte and Vercauteren Bonte & Vercauteren (2018) simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain.
 They replace the matrix $\bar{H}$ with a diagonal matrix $B$ whose diagonal elements are simply the sums of each row in $\bar{H}$.
@@
 Their simplified fixed Hessian method can be formulated as follows:
-$\boldsymbol{\beta}_{t+1}$ $=\boldsymbol{\beta}_{t}-B^{-1}\cdot\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}),$
-$=\boldsymbol{\beta}_{t}-\left[\begin{array}[]{cccc}b_{00}&0&\ldots&0\\ 0&b_{11}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&b_{dd}\\ \end{array}\right]\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]$
-$=\boldsymbol{\beta}_{t}-\left[\begin{array}[]{c}b_{00}\cdot\nabla_{0}\\ b_{11}\cdot\nabla_{1}\\ \vdots\\ b_{dd}\cdot\nabla_{d}\\ \end{array}\right],$
-where $b_{ii}$ is the reciprocal of $\sum_{i=0}^{d}\bar{h}_{0i}$ and $\nabla_{i}$ is the element of $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$.
+$\bm{\beta}_{t+1}$ $=\bm{\beta}_{t}-B^{-1}\cdot\nabla_{\bm{\beta}}l(\bm{\beta}),$
+$=\bm{\beta}_{t}-\left[\begin{array}[]{cccc}b_{00}&0&\ldots&0\\ 0&b_{11}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&b_{dd}\\ \end{array}\right]\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\bm{\beta}_{t}-\left[\begin{array}[]{c}b_{00}\cdot\nabla_{0}\\ b_{11}\cdot\nabla_{1}\\ \vdots\\ b_{dd}\cdot\nabla_{d}\\ \end{array}\right],$
+where $b_{ii}$ is the reciprocal of $\sum_{i=0}^{d}\bar{h}_{0i}$ and $\nabla_{i}$ is the element of $\nabla_{\bm{\beta}}l(\bm{\beta})$.
 Consider a special situation: if all the elements $b_{00},\ldots,b_{dd}$ had the same value $-\eta$ with $\eta>0$, the iterative formula of the SFH method could be given as:
-$\boldsymbol{\beta}_{t+1}$ $=\boldsymbol{\beta}_{t}-(-\eta)\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\boldsymbol{\beta}_{t}+\eta\cdot\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}),$
+$\bm{\beta}_{t+1}$ $=\bm{\beta}_{t}-(-\eta)\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\bm{\beta}_{t}+\eta\cdot\nabla_{\bm{\beta}}l(\bm{\beta}),$
 which is the same as the formula of the naive gradient $ascent$ method.
@@
 As long as we ensure that this new learning rate decreases from a positive floating-point number greater than 1 (such as 2) to 1 in a bounded number of iteration steps, the fixed Hessian Newton’s method guarantees the algorithm will converge eventually.
-The SFH method proposed by Bonte and Vercauteren (2018) has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive.
+The SFH method proposed by Bonte and Vercauteren Bonte & Vercauteren (2018) has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive.
 For machine learning applications, datasets are typically normalized in advance to the range [0,1], satisfying the convergence condition of the SFH method.
 However, in other cases, such as numerical optimization, this condition may not always hold; and (b) the simplified fixed Hessian matrix $B$, as well as the fixed Hessian matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ , can still be singular, especially when the dataset is a high-dimensional sparse matrix, such as the MNIST datasets.
-We extend their work by removing these limitations so as to generalize this simplified fixed Hessian to be invertible in any case and propose a faster gradient variant, which we term quadratic gradient.
+We extend SFH by removing these limitations to generalize this simplified fixed Hessian to be invertible in any case and propose a faster gradient variant, which we term $quadratic$ $gradient$.
 3.2 Quadratic Gradient Definition
-Suppose that a differentiable scalar-valued function $F(\mathbf{x})$ has its gradient $\boldsymbol{g}$ and Hessian matrix $H$, with any matrix $\bar{H}\leq H$ in the Loewner ordering for a maximization problem as follows:
-$\boldsymbol{g}=\left[\begin{array}[]{c}g_{0}\\ g_{1}\\ \vdots\\ g_{d}\\ \end{array}\right],$ $H=\left[\begin{array}[]{cccc}\nabla_{00}^{2}&\nabla_{01}^{2}&\ldots&\nabla_{0d}^{2}\\ \nabla_{10}^{2}&\nabla_{11}^{2}&\ldots&\nabla_{1d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \nabla_{d0}^{2}&\nabla_{d1}^{2}&\ldots&\nabla_{dd}^{2}\\ \end{array}\right],\quad$
-$\bar{H}$ $=\left[\begin{array}[]{cccc}\bar{h}_{00}&\bar{h}_{01}&\ldots&\bar{h}_{0d}\\ \bar{h}_{10}&\bar{h}_{11}&\ldots&\bar{h}_{1d}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{h}_{d0}&\bar{h}_{d1}&\ldots&\bar{h}_{dd}\\ \end{array}\right],$
+Suppose that a differentiable scalar-valued function $F(\mathbf{x})$ has its gradient $\bm{g}$ and Hessian matrix $H$, with any matrix $\bar{H}\leq H$ in the Loewner ordering for a maximization problem as follows:
+$\bm{g}=\left[\begin{array}[]{c}g_{0}\\ g_{1}\\ \vdots\\ g_{d}\\ \end{array}\right],$ $H=\left[\begin{array}[]{cccc}\nabla_{00}^{2}&\nabla_{01}^{2}&\ldots&\nabla_{0d}^{2}\\ \nabla_{10}^{2}&\nabla_{11}^{2}&\ldots&\nabla_{1d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \nabla_{d0}^{2}&\nabla_{d1}^{2}&\ldots&\nabla_{dd}^{2}\\ \end{array}\right],\quad\bar{H}$ $=\left[\begin{array}[]{cccc}\bar{h}_{00}&\bar{h}_{01}&\ldots&\bar{h}_{0d}\\ \bar{h}_{10}&\bar{h}_{11}&\ldots&\bar{h}_{1d}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{h}_{d0}&\bar{h}_{d1}&\ldots&\bar{h}_{dd}\\ \end{array}\right],$
 where $\nabla_{ij}^{2}=\nabla_{ji}^{2}=\frac{\partial^{2}F}{\partial x_{i}\partial x_{j}}$.
@@
 Definition 3.2 (Quadratic Gradient).
-Given such a $\tilde{B}$ above, we define the quadratic gradient as $G=\bar{B}\cdot\boldsymbol{g}$ with a new learning rate $\eta$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$, and $\eta$ should be always no less than 1 and decrease to 1 in a limited number of iteration steps.
-Note that $G$ is still a column vector of the same size as the gradient $\boldsymbol{g}$.
+Given such a $\tilde{B}$ above, we define the quadratic gradient as $G=\bar{B}\cdot\bm{g}$ with a new learning rate $\eta$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$, and $\eta$ should be always no less than 1 and decrease to 1 in a limited number of iteration steps.
+Note that $G$ is still a column vector of the same size as the gradient $\bm{g}$.
 To maximize the function $F(\mathbf{x})$, we can use the iterative formulas: $\mathbf{x}_{t+1}=\mathbf{x}_{t}+\eta\cdot G$, just like the naive gradient.
-To minimize the function $F(x)$ is the same as to just maximize the function $-F(x)$, in which case we need to construct the $\tilde{B}$ by any good lower bound $\bar{H}$ of the Hessian $-H$ of $-F(x)$ or any good upper bound $\bar{H}$ of the Hessian $H$ of $F(x)$.
-We point out here that $\bar{H}$ could be the Hessian matrix $H$ itself.
+Minimizing the function $F(x)$ is equivalent to maximizing the function $-F(x)$.
+In this context, we need to construct $\tilde{B}$ using either a good $lower$ bound $\bar{H}$ for the Hessian $-H$ of $-F(x)$ or a good $upper$ bound $\bar{H}$ for the Hessian $H$ of $F(x)$.
+It is worth noting that $\bar{H}$ can be taken as the Hessian $H$ itself, thereby eliminating the need to find either lower or upper bounds.
+Böhning and Lindsay Böhning & Lindsay (1988) did not proposed a systemic way to find or build the constant Hessian approximate for their fixed Hessian method.
+Perhaps there are no such matrices for most objective functions.
+We mention here that our quadratic gradient doesn’t need a constant (fixed) Hessian replacement to construct and that we can use the original Hessian itself to build our varying diaginal matrix $\tilde{B}$ for the quadratic gradient.
+Because of this loose condition, we propose a system way satisfying sufficient condition to find if there is a fixed Hessian matrix for efficient calculation.
+That is, we first build $\tilde{B}$ from the Hessian of the given objective function and then check if for every its diaginal element $\tilde{B}_{kk}$ there is a constant maximum for $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$ (or constant minimum $|\tilde{B}_{kk}|$).
+If each diagnal element $|\tilde{B}_{kk}|$ all had their minimum $|\tilde{M}_{kk}|$, then the constant diagnial matrix with corresponding diagnal elements $|\tilde{M}_{kk}|$ would satisfy the convengence condiion for SFH method.
+Finally, we should also evaluate the performance of such fixed Hessian replacement since it might not be a $good$ bound.
+This searching method can be employed to any optimization functions.
+For example, we can build $|\tilde{B}_{kk}|$ from the Hessian $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$ for the logistic regressian and attempt to maximise $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$.
+The resulting constant diagnoal matrix is the same as the one built from the $-\frac{1}{4}X^{\top}X$ using the quadratic-gradient construction.
+Böhning and Lindsay Böhning & Lindsay (1988) did not propose a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method.
+This may be due to the absence of such matrices for most objective functions.
+We note that our quadratic gradient approach does not require a constant (fixed) Hessian replacement and instead allows for the use of the original Hessian itself to construct our varying diagonal matrix $\tilde{B}$ for the quadratic gradient.
+Under this less restrictive condition, we present a methodical approach to evaluate whether a fixed Hessian matrix can be employed for efficient computation.
+Specifically, we first construct $\tilde{B}$ from the Hessian of the given objective function and then check if, for each diagonal element $\tilde{B}_{kk}$, there is a constant maximum for $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$ (or a constant minimum $|\tilde{B}_{kk}|$).
+If each diagonal element $|\tilde{B}_{kk}|$ achieves its minimum $|\tilde{M}_{kk}|$, then a constant diagonal matrix with the corresponding diagonal elements $|\tilde{M}_{kk}|$ would satisfy the convergence condition for the SFH method.
+Finally, we should evaluate the performance of such a fixed Hessian replacement, as it might not always provide a good bound.
+This search method can be applied to any optimization function to identify a fixed Hessian.
+However, the absence of a constant Hessian replacement via this method does not necessarily imply that such an approximation does not exist.
+Our method can be employed to identify a fixed Hessian for binary logistic regression.
+For instance, we can construct $|\tilde{B}_{kk}|$ from the Hessian $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$ and then try to maximize $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$.
+The resulting constant diagonal matrix is equivalent to the one obtained from $-\frac{1}{4}X^{\top}X$ using the quadratic-gradient construction.
+Therefore, $-\frac{1}{4}X^{\top}X$ could serve as the fixed Hessian for the fixed Hessian method, precisely as suggested by Böhning and Lindsay Böhning & Lindsay (1988).
 Algorithm 1 The Enhanced Nesterov’s Accelerated Gradient Algorithm
-1:  Require: training dataset $X\in\mathbb{R}^{n\times(1+d)}$; training label $Y\in\mathbb{R}^{n\times 1}$; learning rate $lr\in\mathbb{R}$; and the number $\kappa$ of iterations;
-2:  Ensure: the parameter vector $V\in\mathbb{R}^{(1+d)}$
-3:  Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$ $\bar{H}\in\mathbb{R}^{(1+d)\times(1+d)}$
-4:  Set $V\leftarrow\boldsymbol{0}$, $W\leftarrow\boldsymbol{0}$, $\bar{B}\leftarrow\boldsymbol{0}$ $V\in\mathbb{R}^{(1+d)}$, $W\in\mathbb{R}^{(1+d)}$, $\bar{B}\in\mathbb{R}^{(1+d)\times(1+d)}$
-5:  for $i:=0$ to $d$ do
-6:     # $\epsilon$ is a small positive constant such as $81e-$ 8
-7:     $\bar{B}[i][i]\leftarrow\epsilon$
-8:     for $j:=0$ to $d$ do
-9:        $\bar{B}[i][i]\leftarrow\bar{B}[i][i]+|\bar{H}[i][j]|$
-10:     end for
-11:  end for
-12:  Set $alpha_{0}\leftarrow 0.01$
-13:  Set $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
-14:  for $count:=1$ to $\kappa$ do
-15:     # $Z\in\mathbb{R}^{n}$ will store the inputs for Sigmoid function
-16:     Set $Z\leftarrow\boldsymbol{0}$
-17:     for $i:=1$ to $n$ do
-18:        for $j:=0$ to $d$ do
-19:           $Z[i]\leftarrow Z[i]+Y[i]\times V[j]\times X[i][j]$
-20:        end for
+1:training dataset $X\in\mathbb{R}^{n\times(1+d)}$; training label $Y\in\mathbb{R}^{n\times 1}$; learning rate $lr\in\mathbb{R}$; and the number $\kappa$ of iterations;
+2:the parameter vector $V\in\mathbb{R}^{(1+d)}$
+3:Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$, $V\leftarrow\bm{0}$, $W\leftarrow\bm{0}$, $\bar{B}\leftarrow\bm{0}$
+4:for $i:=0$ to $d$ do $\triangleright$ $\epsilon$ is a small positive constant such as $81e-$ 8
+5:     $\bar{B}[i][i]\leftarrow\epsilon$
+6:     for $j:=0$ to $d$ do
+7:         $\bar{B}[i][i]\leftarrow\bar{B}[i][i]+|\bar{H}[i][j]|$
+8:     end for
+9:end for
+10:Set $alpha_{0}\leftarrow 0.01$, $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
+11:for $count:=1$ to $\kappa$ do
+12:     Set $Z\leftarrow\bm{0}$triangleright $Z\in\mathbb{R}^{n}$ will store the inputs for Sigmoid function
+13:     for $i:=1$ to $n$ do
+14:         for $j:=0$ to $d$ do
+15:              $Z[i]\leftarrow Z[i]+Y[i]\times V[j]\times X[i][j]$
+16:         end for
+17:     end for
+18:     Set $\bm{\sigma}\leftarrow\bm{0}$triangleright $\bm{\sigma}\in\mathbb{R}^{n}$ will store the outputs of Sigmoid function
+19:     for $i:=1$ to $n$ do
+20:         $\bm{\sigma}[i]\leftarrow 1/(1+\exp(-Z[i]))$
 21:     end for
-22:     # $\boldsymbol{\sigma}\in\mathbb{R}^{n}$ will store the outputs of Sigmoid function
-23:     Set $\boldsymbol{\sigma}\leftarrow\boldsymbol{0}$
+22:     Set $\bm{g}\leftarrow\bm{0}$
+23:     for $j:=0$ to $d$ do
 24:         for $i:=1$ to $n$ do
-25:        $\boldsymbol{\sigma}[i]\leftarrow 1/(1+\exp(-Z[i]))$
+25:              $\bm{g}[j]\leftarrow\bm{g}[j]+(1-\bm{\sigma}[i])\times Y[i]\times X[i][j]$
 26:         end for
-27:     Set $\boldsymbol{g}\leftarrow\boldsymbol{0}$
-28:     for $j:=0$ to $d$ do
-29:        for $i:=1$ to $n$ do
-30:           $\boldsymbol{g}[j]\leftarrow\boldsymbol{g}[j]+(1-\boldsymbol{\sigma}[i])\times Y[i]\times X[i][j]$
+27:     end for
+28:     Set $G\leftarrow\bm{0}$
+29:     for $j:=0$ to $d$ do
+30:         $G[j]\leftarrow\bar{B}[j][j]\times\bm{g}[j]$
 31:     end for
-32:     end for
-33:     Set $G\leftarrow\boldsymbol{0}$
-34:     for $j:=0$ to $d$ do
-35:        $G[j]\leftarrow\bar{B}[j][j]\times\boldsymbol{g}[j]$
-36:     end for
-37:     Set $\eta\leftarrow(1-alpha_{0})/alpha_{1}$
-38:     Set $\gamma\leftarrow lr/(n\times count)$
-39:     for $j:=0$ to $d$ do
-40:        $w_{temp}\leftarrow V[j]+(1+\gamma)\times G[j]$
-41:        $V[j]\leftarrow(1-\eta)\times w_{temp}+\eta\times W[j]$
-42:        $W[j]\leftarrow w_{temp}$
-43:     end for
-44:     $alpha_{0}\leftarrow alpha_{1}$
-45:     $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
-46:  end for
-47:  Return $V$
+32:     Set $\eta\leftarrow(1-alpha_{0})/alpha_{1}$, $\gamma\leftarrow(\frac{10}{1+t})/n$
+33:     for $j:=0$ to $d$ do
+34:         $w_{temp}\leftarrow V[j]+(1+\gamma)\times G[j]$, $V[j]\leftarrow(1-\eta)\times w_{temp}+\eta\times W[j]$, $W[j]\leftarrow w_{temp}$
+35:     end for
+36:     $alpha_{0}\leftarrow alpha_{1}$, $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
+37:end for
+38:Return $V$
 3.3 Quadratic Gradient Algorithms
-This gradient variant, Quadratic Gradient, can be used to enhance various first-order gradient algorithms:
+Quadratic gradient can be used to enhance various first-order gradient algorithms:
   1. 1.
     NAG is a different variant of the momentum method to give the momentum term much more prescience. The iterative formulas of the gradient $ascent$ method for NAG are as follows:
-    $V_{t+1}$ $=\boldsymbol{\beta}_{t}+\eta_{t}\cdot\nabla J(\boldsymbol{\beta}_{t}),$ (3)
-    $\boldsymbol{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
-    where $V_{t+1}$ is the intermediate variable used for updating the final weight $\boldsymbol{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position (Kim et al., 2018a).
-The enhanced NAG is to replace (3) with $V_{t+1}=\boldsymbol{\beta}_{t}+N_{t}\cdot G,$ where $N_{t}$ is the new learning rate for the enhanced algorithm usually setted to $1+\eta_{t}$.
+    $V_{t+1}$ $=\bm{\beta}_{t}+\eta_{t}\cdot\nabla J(\bm{\beta}_{t}),$ (3)
+    $\bm{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
+    where $V_{t+1}$ is the intermediate variable used for updating the final weight $\bm{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position (Kim et al., 2018a).
+The enhanced NAG is to replace equation 3 with $V_{t+1}=\bm{\beta}_{t}+N_{t}\cdot G,$ where $N_{t}$ is the new learning rate for the enhanced algorithm usually setted to $1+\eta_{t}$.
 Our enhanced NAG is described in Algorithm 1 .
@@
     Adagrad is a gradient-based algorithm suitable for dealing with sparse data.
-The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\boldsymbol{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
-    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{\eta_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}\boldsymbol{g}_{[i]}^{(t)}\cdot\boldsymbol{g}_{[i]}^{(t)}}}\cdot\boldsymbol{g}_{[i]}^{(t)},$
-    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{N_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
+The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\bm{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
+    $\bm{\beta}_{[i]}^{(t+1)}$ $=\bm{\beta}_{[i]}^{(t)}-\frac{\eta_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}\bm{g}_{[i]}^{(t)}\cdot\bm{g}_{[i]}^{(t)}}}\cdot\bm{g}_{[i]}^{(t)},$
+    $\bm{\beta}_{[i]}^{(t+1)}$ $=\bm{\beta}_{[i]}^{(t)}-\frac{N_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
   3. 3.
     Adam is an optimization algorithm that combines the benefits of Adagrad, which adapts learning rates for each parameter, with momentum, allowing it to converge faster and handle sparse gradients more effectively.
-Similar to how Adagrad transitions to its enhanced versions, the update formula for enhanced Adam, also employs a new learning rate and replaces the gradient with its quadratic form.
+Similar to how Adagrad transitions to its enhanced version, the update formula for enhanced Adam, also employs a new learning rate $N_{t}$ and replaces the gradient with its quadratic form.
 (a) The iDASH dataset
-(b) The Edinburgh dataset
+(b) The edin dataset
 (c) The lbw dataset
@@
 (f) The uis dataset
-(g) restructured MNIST dataset
-(h) The private financial dataset
-Figure 1: The training results of NAG and Enhanced NAG in the clear
+(g) The MNIST dataset
+(h) The financial dataset
+Figure 1: The training results of NAG and Enhanced NAG in the clear.
 (a) The iDASH dataset
-(b) The Edinburgh dataset
+(b) The edin dataset
 (c) The lbw dataset
@@
 (f) The uis dataset
-(g) restructured MNIST dataset
-(h) The private financial dataset
-Figure 2: The training results of Adagrad and Enhanced Adagrad in the clear
+(g) The MNIST dataset
+(h) The financial dataset
+Figure 2: The training results of Adagrad and Enhanced Adagrad in the clear.
 (a) The iDASH dataset
-(b) The Edinburgh dataset
+(b) The edin dataset
 (c) The lbw dataset
@@
 (f) The uis dataset
-(g) restructured MNIST dataset
-(h) The private financial dataset
-Figure 3: The training results of Adam and Enhanced Adam in the clear
+(g) The MNIST dataset
+(h) The financial dataset
+Figure 3: The training results of Adam and Enhanced Adam in the clear.
 Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM.
-Given that our focus is on the convergence speed of the algorithms during training, we use the loss function, specifically maximum likelihood estimation (MLE), as the only indicator.
+Given that our focus is on the convergence speed of the algorithms during training, we use the loss function $l(\bm{\beta})$, specifically maximum likelihood estimation (MLE), as the only indicator.
 We evaluate six algorithms—NAG, Adagrad, Adam, and their quadratic-gradient variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam, respectively)—using the datasets adopted by Kim et al.
-(2018a): the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (Edinburgh), the Low Birth Weight Study (lbw), Nhanes III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study datasets (uis).
+Kim et al.
+(2018a): the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (edin), the Low Birth Weight Study (lbw), Nhanes III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study datasets (uis).
 The genomic dataset from the third task of the 2017 iDASH competition comprises 1,579 records, each featuring 103 binary genotypes and a binary phenotype indicating whether the patient has cancer.
@@
 We also evaluate these algorithms on two large datasets from Han et al.
+Han et al.
 (2019): a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features.
 For a fair comparison with the baseline work (Kim et al., 2018a), we use the learning rate $1+\frac{10}{1+t}$, in line with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate.
-In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the binary logistic regression model.
-Figures 21, and 3 show that, except for the Enhanced Adagrad and Adam methods on the iDASH genomic dataset, our enhanced methods converge faster than their original counterparts in all other cases.
-All figures demonstrate that the enhanced algorithms converge empirically.
-In all the Python experiments, the time required to calculate $\bar{B}$ in the quadratic gradient $G$ before running the iterations, as well as the time to execute each iteration for various algorithms, is minimal and can be considered negligible (a few seconds).
-An important observation is that the enhanced algorithms have a better performance when using learning rates between $1$ and $2$ than using learning rate otherelse.
-When using a learning rate greater than $3$, the algorithm is almost certain not to converge.
-When employing quadratic gradient algorithms, we recommend adopting the learning rate setting $1+A\cdot\gamma^{t}$ as the optimal choice, where $t$ represents the iteration number, $A$ is a positive float number typically set to no less than $1$, and $\gamma$ is usually a positive number less than 1, controlling the rate of decay.
-We strongly recommend that $A$ is set to $1.0$ and $\gamma$ is tuned to search for optimal results, or that one can set $A$ to a small number less than $1.0$ and $\gamma$ to a larger number greater than $0.9$.
-Analysis In Figure 2(a), the enhanced Adagrad algorithm failed to outperform the original Adagrad algorithm.
-The possible reason for that might be related to the limitations of the raw Adagrad method.
-Without a doubt, Adagrad is a novel algorithm initiated to accelerate each element of the gradient with different learning rates.
-However, Adagrad tends to converge to a suboptimal solution due to its aggressive, monotonically decreasing learning rates.
-This would lead to its main limitation that in the later training phase every learning rate for different components of the gradient is too close to zero due to keeping adding positive additional terms to the denominator, stopping the algorithm from learning anything.
-On the other hand, the original Adagrad method has another little-noticed limitation: the learning rate in the first few iterations tends to be large.
-While this limitation does not affect the performance of the original Adagrad method to some extent, the enhanced Adagrad method exacerbates this phenomenon significantly, leading to the Learning-Rate Explosion .
-Therefore, the enhanced Adagrad cannot be applied to general optimization problems such as Rosenbrock’s function.
-The exploding learning rate would be too large for the algorithm to survive the first several iterations, finally leading the optimization function to some point where its output cannot be represented by the computer system.
-This might explain why the performance of this algorithm in all cases, not just on the iDASH genome dataset, seems to be meaningless and unstable in the first few iterations.
-Several improved algorithms upon the Adagrad method, such as RMSProp, have been proposed in order to address these issues existed, via using an exponential moving average of historical gradients rather than just the sum of all squared gradients from the beginning of training.
-We might be able to overcome the problems existing in the enhanced Adagrad method by adopting the enhanced Adagrad-like variants, like the enhanced Adadelta method and the enhanced RMSProp method.
-The reasons why the enhanced algorithms have a better performance might be that quadratic gradient can incorporate curves into first-order gradient algorithms.
-(a) The raw first-order gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{lr}\cdot\boldsymbol{g}$
-(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{LR}\cdot G$
-Figure 4: The training outcomes for both the raw gradient algorithm and the raw quadratic gradient ascent algorithm in the clear setting, conducted on the lbw dataset
-Gradient And Quadratic Gradient We executed the raw gradient ascent algorithm and raw quadratic gradient algorithm on the lbw dataset using various learning rates.
-Figure 4 shows the detailed results of this experiment.
-It is precisely these results that inspired the present authors to come up with the idea of quadratic gradient.
-Similar to the gradient, the quadratic gradient carrying the curve also exhibits the characteristic of smooth progression.
-That is, when the learning rate is gradually increased, both the gradient and the quadratic gradient will show corresponding gradual changes in performance rather than abrupt shifts.
-4 Homomorphic Training
-Adagrad method is not a practical solution for homomorphic LR due to its frequent inversion operations.
-It seems plausible that the enhanced NAG is likely the reasonable choice for homomorphic LR training.
+The enhanced Adagrad algorithm use $N_{t}=1+0.01$ as the learning rate while the enhanced Adam algorithms adopt the following learning rate setting: $\alpha=1+0.001,$ $\beta_{1}=0.9,$ and $\beta_{2}=0.999.$ In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\bm{\beta})$.
+From the empirical results presented in Figures 12, and 3, we conclude that all the enhanced algorithms converge, and that the enhanced NAG algorithms outperform the original NAG method in terms of convergence rate.
+Although the enhanced Adagrad and Adam methods do not achieve faster convergence than their original counterparts on the iDASH genomic dataset, they demonstrate clear advantages in all other cases.
+The reason for this discrepancy is beyond the scope of this paper and will be addressed as part of future work.
+One possible explanation for the superior performance of the enhanced algorithms is that the quadratic gradient integrates curvature information into first-order gradient methods.
+An important observation is that the enhanced algorithms demonstrate better performance with learning rates between $1$ and $2$ compared to other values.
+When the learning rate exceeds $3$, the algorithm is nearly guaranteed not to converge.
+For quadratic gradient algorithms, we recommend employing an exponentially decaying learning rate, such as $1+A\cdot\gamma^{t}$, where $t$ denotes the iteration number, $A$ is a positive constant typically set to no less than $1$, and $\gamma$ is a positive number less than $1$ that controls the rate of decay.
+Gradient And Quadratic Gradient We executed the raw gradient ascent algorithm and the raw quadratic gradient algorithm on the LBW dataset using various learning rates.
+Figure 4 presents the detailed results of this experiment.
+It was precisely these results that directly inspired the authors to develop the concept of the quadratic gradient.
+Similar to the gradient, the quadratic gradient, which incorporates curvature, also demonstrates smooth progression.
+In particular, as the learning rate is gradually increased, both the gradient and the quadratic gradient reveal corresponding gradual changes in performance, rather than sudden jumps.
+(a) The raw first-order gradient ascent algorithm with iterative formulas: $\bm{\beta}_{t+1}=\bm{\beta}_{t}+\mathrm{lr}\cdot\bm{g}$
+(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\bm{\beta}_{t+1}=\bm{\beta}_{t}+\mathrm{LR}\cdot G$
+Figure 4: The training outcomes for both the raw gradient algorithm and the raw quadratic gradient ascent algorithm in the clear setting, conducted on the lbw dataset.
+4 Secure Training
 We adopt the enhanced NAG method to implement secure logistic regression training based on HE.
@@
 We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al.
+Kim et al.
 (2018a), utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
 Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al.
+Kim et al.
 (2018a) used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$.
@@
 The ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ are as follows:
-X $=\left[\begin{array}[]{cccc}1&x_{11}&\ldots&x_{1d}\\ 1&x_{21}&\ldots&x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n1}&\ldots&x_{nd}\\ \end{array}\right],\text{Y}=\left[\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\\ \end{array}\right],$
-$\text{ct}_{Z}$ $=Enc\left[\begin{array}[]{cccc}y_{1}&y_{1}x_{11}&\ldots&y_{1}x_{1d}\\ y_{2}&y_{2}x_{21}&\ldots&y_{2}x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ y_{n}&y_{n}x_{n1}&\ldots&y_{n}x_{nd}\\ \end{array}\right],$
-$\text{ct}_{\beta}^{(0)}$ $=Enc\left[\begin{array}[]{cccc}\beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \vdots&\vdots&\ddots&\vdots\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \end{array}\right],$
-$\text{ct}_{\bar{B}}$ $=Enc\left[\begin{array}[]{cccc}\bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \end{array}\right],$
+$\text{X}=\left[\begin{array}[]{cccc}1&x_{11}&\ldots&x_{1d}\\ 1&x_{21}&\ldots&x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n1}&\ldots&x_{nd}\\ \end{array}\right],\text{Y}=\left[\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\\ \end{array}\right],$ $\text{ct}_{Z}=Enc\left[\begin{array}[]{cccc}y_{1}&y_{1}x_{11}&\ldots&y_{1}x_{1d}\\ y_{2}&y_{2}x_{21}&\ldots&y_{2}x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ y_{n}&y_{n}x_{n1}&\ldots&y_{n}x_{nd}\\ \end{array}\right],$
+$\text{ct}_{\beta}^{(0)}=Enc\left[\begin{array}[]{cccc}\beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \vdots&\vdots&\ddots&\vdots\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \end{array}\right],$ $\text{ct}_{\bar{B}}=Enc\left[\begin{array}[]{cccc}\bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \end{array}\right],$
 where $\bar{B}_{[i][i]}$ is the diagonal element of $\bar{B}$ that is built from $-\frac{1}{4}X^{\top}X$.
@@
 Num
-Method deg $g$
+Method
 Iter
@@
 AUC
-iDASH 1579 18 Ours 5 4 0.08 4.43 61.46 0.696
-(Kim et al., 2018a) 5 7 0.04 6.07 62.87 0.689
+iDASH 1579 18 Ours 4 0.08 4.43 61.46 0.696
+Baseline 7 0.04 6.07 62.87 0.689
 Table 2: Implementation Results for other datasets with 5-fold CV
@@
 Num
-Method deg $g$
+Method
 Iter
@@
 AUC
-Edinburgh 1253 9 Ours 5 4 0.04 0.6 89.52 0.943
-(Kim et al., 2018a) 5 7 0.02 3.6 91.04 0.958
-lbw 189 9 Ours 5 4 0.04 0.6 71.35 0.667
-(Kim et al., 2018a) 5 7 0.02 3.3 69.19 0.689
-nhanes3 15649 15 Ours 5 4 0.31 4.5 79.23 0.637
-(Kim et al., 2018a) 5 7 0.16 7.3 79.22 0.717
-pcs 379 9 Ours 5 4 0.04 0.6 63.20 0.733
-(Kim et al., 2018a) 5 7 0.02 3.5 68.27 0.740
-uis 575 8 Ours 5 4 0.04 0.6 74.43 0.597
-(Kim et al., 2018a) 5 7 0.02 3.5 74.44 0.603
+Edin 1253 8 Ours 4 0.04 0.6 89.52 0.943
+Baseline 7 0.02 3.6 91.04 0.958
+lbw 189 8 Ours 4 0.04 0.6 71.35 0.667
+Baseline 7 0.02 3.3 69.19 0.689
+nhanes3 15649 15 Ours 4 0.31 4.5 79.23 0.637
+Baseline 7 0.16 7.3 79.22 0.717
+pcs 379 9 Ours 4 0.04 0.6 63.20 0.733
+Baseline 7 0.02 3.5 68.27 0.740
+uis 575 8 Ours 4 0.04 0.6 74.43 0.597
+Baseline 7 0.02 3.5 74.44 0.603
 5 Experiments
@@
 All the experiments on the ciphertexts were conducted on a public cloud with 32 vCPUs and 64 GB RAM.
-For a fair comparison with (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets.
+For a fair comparison with the baseline (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets.
 Like (Kim et al., 2018a), We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators.
@@
 We adopt the same packing method that Kim et al.
+Kim et al.
 (2018a) proposed and hence our solution has similar storage of ciphertexts to (Kim et al., 2018a) with some extra ciphertexts to encrypt the $\bar{B}$.
@@
 The quadratic gradient presented in this work can be constructed from the Hessian matrix directly, and thus somehow combines the first-order Newton’s method and the second-order gradient (descent) method together.
-There is a promising chance that quadratic gradient could accelerate other gradient methods, which is an open future work.
+There is a promising chance that quadratic gradient could accelerate other gradient methods.
 Also, quadratic gradient might substitute and supersede the line-search method, for example when using enhanced Adagrad-like methods, and could use gradient descent methods to accelerate Newton’s method, resulting in super-quadratic algorithms.
-Impact Statements
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+It is highly likely that the quadratic gradient approach could combine first-order gradient descent/ascent algorithms with the second-order Newtonian method, and that it could be applicable to a wide range of numerical optimization problems.

### v9

@@
 Kim et al.
-Kim et al.
 (2018b) discussed the problem of performing LR training in an encrypted environment.
-They used the full batch gradient descent in the training process and the least-squares method to get the approximation of the sigmoid function.
-In the iDASH 2017 competition, Bonte and Vercauteren Bonte & Vercauteren (2018), Kim et al.
-Kim et al.
+They employed full-batch gradient descent during the training process and utilized the least-squares method to approximate the sigmoid function.
+In the 2017 iDASH competition, Bonte & Vercauteren (2018), Kim et al.
 (2018a), Chen et al.
-Chen et al.
 (2018), and Crawford et al.
-Crawford et al.
-(2018) all investigated the same problem that Kim et al.
-Kim et al.
-(2018b) studied.
-In the iDASH competition of 2018, Kim et al.
-Kim et al.
+(2018) all addressed the same problem explored by Kim et al.
+(2018b).
+In the 2018 iDASH competition, Kim et al.
 (2019b) and Blatt et al.
-Blatt et al.
-(2019) further worked on it for an efficient packing and semi-parallel algorithm.
+(2019) further developed the problem, focusing on efficient packing and a semi-parallel algorithm.
 There are other related works (Kim et al., 2019a; Bergamaschi et al., 2019; Ogilvie et al., 2020) focusing on various aspects but the papers most relevant to this work are (Bonte & Vercauteren, 2018) and (Kim et al., 2018a).
-Bonte and Vercauteren Bonte & Vercauteren (2018) developed a practical algorithm called the simplified fixed Hessian (SFH) method.
+Bonte & Vercauteren (2018) developed a practical algorithm called the simplified fixed Hessian (SFH) method.
 Our study extends their work and adopts the ciphertext packing technique proposed by Kim et al.
-Kim et al.
 (2018a) for efficient homomorphic computation.
@@
     We develop three enhanced gradient algorithms by equipping the original ones with quadratic gradient.
-The resulting algorithms in most cases demonstrate high performance in terms of convergence speed.
+The resulting algorithms all converge and, in most cases, demonstrate strong performance in terms of convergence speed.
   3. 3.
@@
 Cheon et al.
-Cheon et al.
 (2017) proposed a method to construct an HE scheme with a rescaling procedure which could eliminate this technical bottleneck effectively.
@@
 Kim et al.
-Kim et al.
 (2018a) proposed an efficient and promising database-encoding method by using SIMD technique, which could make full use of the computation and storage resources.
-Suppose that a database has a training dataset consisting of $n$ samples with $(1+d)$ covariates, they packed the training dataset $Z$ into a single ciphertext in a row-by-row manner.
+Suppose a database contains a training dataset $Z$ comprising $n$ samples and $(1+d)$ covariates.
+The dataset matrix $Z$ is packed into a single ciphertext in a row-by-row manner.
 When employing this encoding scheme, we can manipulate the data matrix $Z$ through HE operations on the ciphertext $Enc[Z]$, utilizing only three HE operations - rotation, addition, and multiplication.
@@
 Han et al.
-Han et al.
 (2019) mentioned several basic but important operations used by Kim et al.
-Kim et al.
 (2018a) in their implementation, such as a procedure named “SumColVec” to compute the summation of the columns of a matrix.
@@
 This novel technique is called the fixed Hessian Newton’s method.
-Böhning and Lindsay Böhning & Lindsay (1988) have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\bm{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})-\bar{H}$ is non-negative definite.
+Böhning & Lindsay (1988) have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\bm{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})-\bar{H}$ is non-negative definite.
 With such a fixed Hessian matrix $\bar{H}$, the iteration for Newton’s method can be simplified to:
@@
 3.1 Simplified Fixed Hessian
-Bonte and Vercauteren Bonte & Vercauteren (2018) simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain.
+Bonte & Vercauteren (2018) simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain.
 They replace the matrix $\bar{H}$ with a diagonal matrix $B$ whose diagonal elements are simply the sums of each row in $\bar{H}$.
@@
 As long as we ensure that this new learning rate decreases from a positive floating-point number greater than 1 (such as 2) to 1 in a bounded number of iteration steps, the fixed Hessian Newton’s method guarantees the algorithm will converge eventually.
-The SFH method proposed by Bonte and Vercauteren Bonte & Vercauteren (2018) has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive.
+The SFH method proposed by Bonte & Vercauteren (2018) has two limitations: (a) in the construction of the simplified fixed Hessian, all entries in the symmetric matrix $\bar{H}$ need to be non-positive.
 For machine learning applications, datasets are typically normalized in advance to the range [0,1], satisfying the convergence condition of the SFH method.
-However, in other cases, such as numerical optimization, this condition may not always hold; and (b) the simplified fixed Hessian matrix $B$, as well as the fixed Hessian matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ , can still be singular, especially when the dataset is a high-dimensional sparse matrix, such as the MNIST datasets.
+However, in other cases, such as numerical optimization, this condition may not always hold; and (b) the simplified fixed Hessian matrix $B$, as well as the fixed Hessian $\bar{H}=-\frac{1}{4}X^{\top}X$ , can still be singular, especially when the dataset is a high-dimensional sparse matrix, such as the MNIST datasets.
 We extend SFH by removing these limitations to generalize this simplified fixed Hessian to be invertible in any case and propose a faster gradient variant, which we term $quadratic$ $gradient$.
@@
 where $\epsilon$ is a small positive constant to avoid division by zero (usually set to $81e-$ 8).
-As long as $\tilde{B}$ satisfies the convergence condition of the above fixed Hessian method, $\tilde{B}\leq H$, we can use this approximation $\tilde{B}$ of the Hessian matrix as a lower bound.
-Since we already assume that $\bar{H}\leq H$, it will suffice to show that $\tilde{B}\leq\bar{H}$.
-We prove $\tilde{B}\leq\bar{H}$ in a similar way that (Bonte & Vercauteren, 2018) did.
+As long as $\tilde{B}$ satisfies the convergence condition of the aforementioned fixed Hessian method, namely $\tilde{B}\leq H$, we can use this approximation $\tilde{B}$ of the Hessian matrix as a lower bound.
+Given that we have already assumed $\bar{H}\leq H$, it suffices to demonstrate that $\tilde{B}\leq\bar{H}$.
+We establish $\tilde{B}\leq\bar{H}$ in a manner similar to that employed by Bonte & Vercauteren (2018).
 Lemma 3.1.
@@
 Definition 3.2 (Quadratic Gradient).
-Given such a $\tilde{B}$ above, we define the quadratic gradient as $G=\bar{B}\cdot\bm{g}$ with a new learning rate $\eta$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$, and $\eta$ should be always no less than 1 and decrease to 1 in a limited number of iteration steps.
-Note that $G$ is still a column vector of the same size as the gradient $\bm{g}$.
-To maximize the function $F(\mathbf{x})$, we can use the iterative formulas: $\mathbf{x}_{t+1}=\mathbf{x}_{t}+\eta\cdot G$, just like the naive gradient.
+Given the previously defined $\tilde{B}$, we define the quadratic gradient as $G=\bar{B}\cdot\bm{g}$ with a new learning rate $N_{t}$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$.
+The learning rate $N_{t}$ should always be no less than 1 and is designed to decrease to 1 over a limited number of iterations.
+It is important to note that $G$ remains a column vector of the same dimension as the gradient $\bm{g}$.
+To maximize the function $F(\mathbf{x})$, we can employ the iterative formula: $\mathbf{x}_{t+1}=\mathbf{x}_{t}+N_{t}\cdot G$, similar to the naive gradient approach.
 Minimizing the function $F(x)$ is equivalent to maximizing the function $-F(x)$.
@@
 It is worth noting that $\bar{H}$ can be taken as the Hessian $H$ itself, thereby eliminating the need to find either lower or upper bounds.
-Böhning and Lindsay Böhning & Lindsay (1988) did not proposed a systemic way to find or build the constant Hessian approximate for their fixed Hessian method.
-Perhaps there are no such matrices for most objective functions.
-We mention here that our quadratic gradient doesn’t need a constant (fixed) Hessian replacement to construct and that we can use the original Hessian itself to build our varying diaginal matrix $\tilde{B}$ for the quadratic gradient.
-Because of this loose condition, we propose a system way satisfying sufficient condition to find if there is a fixed Hessian matrix for efficient calculation.
-That is, we first build $\tilde{B}$ from the Hessian of the given objective function and then check if for every its diaginal element $\tilde{B}_{kk}$ there is a constant maximum for $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$ (or constant minimum $|\tilde{B}_{kk}|$).
-If each diagnal element $|\tilde{B}_{kk}|$ all had their minimum $|\tilde{M}_{kk}|$, then the constant diagnial matrix with corresponding diagnal elements $|\tilde{M}_{kk}|$ would satisfy the convengence condiion for SFH method.
-Finally, we should also evaluate the performance of such fixed Hessian replacement since it might not be a $good$ bound.
-This searching method can be employed to any optimization functions.
-For example, we can build $|\tilde{B}_{kk}|$ from the Hessian $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$ for the logistic regressian and attempt to maximise $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$.
-The resulting constant diagnoal matrix is the same as the one built from the $-\frac{1}{4}X^{\top}X$ using the quadratic-gradient construction.
-Böhning and Lindsay Böhning & Lindsay (1988) did not propose a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method.
+Böhning & Lindsay (1988) did not propose a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method.
 This may be due to the absence of such matrices for most objective functions.
@@
 However, the absence of a constant Hessian replacement via this method does not necessarily imply that such an approximation does not exist.
-Our method can be employed to identify a fixed Hessian for binary logistic regression.
-For instance, we can construct $|\tilde{B}_{kk}|$ from the Hessian $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$ and then try to maximize $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$.
-The resulting constant diagonal matrix is equivalent to the one obtained from $-\frac{1}{4}X^{\top}X$ using the quadratic-gradient construction.
-Therefore, $-\frac{1}{4}X^{\top}X$ could serve as the fixed Hessian for the fixed Hessian method, precisely as suggested by Böhning and Lindsay Böhning & Lindsay (1988).
 Algorithm 1 The Enhanced Nesterov’s Accelerated Gradient Algorithm
@@
 We evaluate six algorithms—NAG, Adagrad, Adam, and their quadratic-gradient variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam, respectively)—using the datasets adopted by Kim et al.
-Kim et al.
 (2018a): the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (edin), the Low Birth Weight Study (lbw), Nhanes III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study datasets (uis).
@@
 We also evaluate these algorithms on two large datasets from Han et al.
-Han et al.
 (2019): a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features.
-For a fair comparison with the baseline work (Kim et al., 2018a), we use the learning rate $1+\frac{10}{1+t}$, in line with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate.
-The enhanced Adagrad algorithm use $N_{t}=1+0.01$ as the learning rate while the enhanced Adam algorithms adopt the following learning rate setting: $\alpha=1+0.001,$ $\beta_{1}=0.9,$ and $\beta_{2}=0.999.$ In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\bm{\beta})$.
-From the empirical results presented in Figures 12, and 3, we conclude that all the enhanced algorithms converge, and that the enhanced NAG algorithms outperform the original NAG method in terms of convergence rate.
+For a fair comparison with the baseline work by Kim et al.
+(2018a), the enhanced NAG algorithm employs the learning rate $1+\frac{10}{1+t}$, consistent with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate.
+The enhanced Adagrad algorithm utilizes $N_{t}=1+0.01$ as the learning rate, while the enhanced Adam algorithm adopts the following learning rate settings: $\alpha=1+0.001$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$.
+Refer to (Kingma & Ba, 2014) for further details regarding these parameters.
+In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\bm{\beta})$.
+From the empirical results presented in Figures 12, and 3, we conclude that all the enhanced algorithms converge, and that the enhanced NAG algorithm consistently outperforms the original NAG method in terms of convergence rate.
 Although the enhanced Adagrad and Adam methods do not achieve faster convergence than their original counterparts on the iDASH genomic dataset, they demonstrate clear advantages in all other cases.
@@
 We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al.
-Kim et al.
 (2018a), utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
 Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al.
-Kim et al.
 (2018a) used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$.
@@
 In addition, the data owner needs to upload one more ciphertext encrypting the matrix $\bar{B}$.
-However, the enhanced algorithm converges faster, and we believe it can compensate for the mentioned drawbacks.
+However, the enhanced algorithm converges faster, and we believe it can compensate for the mentioned shortcomings.
 Table 1: Implementation Results for iDASH datasets with 10-fold CV
@@
 We adopt the same packing method that Kim et al.
-Kim et al.
 (2018a) proposed and hence our solution has similar storage of ciphertexts to (Kim et al., 2018a) with some extra ciphertexts to encrypt the $\bar{B}$.
@@
 6 Conclusion
-In this paper, we proposed a faster gradient variant called quadratic gradient, and implemented the quadratic-gradient version of NAG in the encrypted domain to train the logistic regression model.
-The quadratic gradient presented in this work can be constructed from the Hessian matrix directly, and thus somehow combines the first-order Newton’s method and the second-order gradient (descent) method together.
-There is a promising chance that quadratic gradient could accelerate other gradient methods.
-Also, quadratic gradient might substitute and supersede the line-search method, for example when using enhanced Adagrad-like methods, and could use gradient descent methods to accelerate Newton’s method, resulting in super-quadratic algorithms.
-It is highly likely that the quadratic gradient approach could combine first-order gradient descent/ascent algorithms with the second-order Newtonian method, and that it could be applicable to a wide range of numerical optimization problems.
+In this paper, we proposed a faster gradient variant, termed quadratic gradient, and implemented the quadratic-gradient version of NAG in the encrypted domain to train a logistic regression model.
+The quadratic gradient introduced in this work can be directly constructed from the Hessian, effectively merging first-order gradient (descent) methods with the second-order Newton’s method.
+There is a promising potential for the quadratic gradient to accelerate other gradient methods, and it may serve as an alternative or even a replacement for the traditional line-search method.

### v10

@@
     We implement privacy-preserving logistic regression training using the enhanced NAG method, to our best knowledge, which seems to be a great choice without compromising much on computation and storage.
+  4. 4.
+    We present a systematic approach to determine whether constant Hessian approximations exist for fixed Hessian methods, a systematic approach to evaluate whether fixed Hessian matrices can be used for efficient computation.
 2 Preliminaries
@@
 Such a coincidence not only helps generate the idea behind this work but also leads us to believe that there is a connection between the Hessian matrix and the learning rate of the gradient (descent) method.
-We regard $B^{-1}\cdot\nabla_{i}$ as a novel enhanced gradient variant and allocate a distinct learning rate to it.
+We regard $B^{-1}\cdot\nabla_{\bm{\beta}}$ as a novel enhanced gradient variant and allocate a distinct learning rate to it.
 As long as we ensure that this new learning rate decreases from a positive floating-point number greater than 1 (such as 2) to 1 in a bounded number of iteration steps, the fixed Hessian Newton’s method guarantees the algorithm will converge eventually.
@@
     where $V_{t+1}$ is the intermediate variable used for updating the final weight $\bm{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position (Kim et al., 2018a).
-The enhanced NAG is to replace equation 3 with $V_{t+1}=\bm{\beta}_{t}+N_{t}\cdot G,$ where $N_{t}$ is the new learning rate for the enhanced algorithm usually setted to $1+\eta_{t}$.
+The enhanced NAG is to replace equation 3 with $V_{t+1}=\bm{\beta}_{t}+N_{t}\cdot G,$ where $N_{t}$ is the new learning rate for the enhanced algorithm usually set to $1+\eta_{t}$.
 Our enhanced NAG is described in Algorithm 1 .
@@
 We leave the computation of matrix $\bar{B}$ to data owner and let the data owner upload the ciphertext encrypting the $\bar{B}$ to the cloud.
-Since data owner has to prepare the dataset and normalize it, it would also be practicable for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
+Since data owner has to prepare the dataset and normalize it, it would also be practical for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
 Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model.
@@
 where $\bar{B}_{[i][i]}$ is the diagonal element of $\bar{B}$ that is built from $-\frac{1}{4}X^{\top}X$.
-The pulbic cloud takes the three ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ and evaluates the enhanced NAG algorithm to find a decent weight vector by updating the vector $\text{ct}_{\beta}^{(0)}$.
+The public cloud takes the three ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ and evaluates the enhanced NAG algorithm to find a decent weight vector by updating the vector $\text{ct}_{\beta}^{(0)}$.
 Refer to (Kim et al., 2018a) for a detailed description about how to calculate the gradient by HE programming.

### v12

@@
 Training logistic regression over encrypted data has been a compelling approach in addressing security concerns for several years.
-In this paper, we introduce an efficient gradient variant, called $quadratic$ $gradient$, for privacy-preserving logistic regression training.
+In this paper, we introduce an efficient gradient variant, called $quadratic$ $gradient$, which can be used for privacy-preserving logistic regression training.
 We enhance Nesterov’s Accelerated Gradient (NAG), Adaptive Gradient Algorithm (Adagrad) and Adam algorithms by incorporating their quadratic gradients and evaluate these improved algorithms on various datasets.
@@
 Moreover, we applied the enhanced NAG method to implement homomorphic logistic regression training, achieving comparable results within just 4 iterations.
-There is a good chance that the quadratic gradient approach could integrate first-order gradient descent/ascent algorithms with the second-order Newton-Raphson methods, and that it could be applied to a wide range of numerical optimization problems.
+There is a great chance that the quadratic gradient approach could integrate first-order gradient descent/ascent algorithms with the second-order Newton-Raphson methods, and that quadratic gradient could be applied to a wide range of numerical optimization problems.
+Homomorphic Encryption, Logistic Regression, Quadratic Gradient, Simplified Fixed Hessian, Nesterov’s Accelerated Gradient, Gradient Ascent Algorithm
 1 Introduction
@@
 (2018b) discussed the problem of performing LR training in an encrypted environment.
-They employed full-batch gradient descent during the training process and utilized the least-squares method to approximate the sigmoid function.
-In the 2017 iDASH competition, Bonte & Vercauteren (2018), Kim et al.
+They employed full batch gradient descent in the training process and utilized the least-squares method to approximate the sigmoid function.
+In the 2017 iDASH competition, Bonte and Vercauteren (2018), Kim et al.
 (2018a), Chen et al.
 (2018), and Crawford et al.
-(2018) all addressed the same problem explored by Kim et al.
-(2018b).
-In the 2018 iDASH competition, Kim et al.
+(2018) all investigated the same problem that Kim et al.
+(2018b) studied.
+In the iDASH competition of 2018, Kim et al.
 (2019b) and Blatt et al.
-(2019) further developed the problem, focusing on efficient packing and a semi-parallel algorithm.
-There are other related works (Kim et al., 2019a; Bergamaschi et al., 2019; Ogilvie et al., 2020) focusing on various aspects but the papers most relevant to this work are (Bonte & Vercauteren, 2018) and (Kim et al., 2018a).
-Bonte & Vercauteren (2018) developed a practical algorithm called the simplified fixed Hessian (SFH) method.
+(2019) further worked on it for an efficient packing and semi-parallel algorithm.
+There are other related works (Kim et al., 2019a; Ogilvie et al., 2020) focusing on various aspects but the papers most relevant to this work are (Bonte & Vercauteren, 2018) and (Kim et al., 2018a).
+Bonte and Vercauteren (2018) developed a practical algorithm called the simplified fixed Hessian (SFH) method.
 Our study extends their work and adopts the ciphertext packing technique proposed by Kim et al.
@@
   2. 2.
-    We develop three enhanced gradient algorithms by equipping the original ones with quadratic gradient.
-The resulting algorithms all converge and, in most cases, demonstrate strong performance in terms of convergence speed.
+    We develop three enhanced gradient algorithms by equipping the original ones with quadratic gradient. The resulting algorithms exhibit state-of-the-art performance in terms of convergence speed.
   3. 3.
@@
   4. 4.
-    We present a systematic approach to determine whether constant Hessian approximations exist for fixed Hessian methods, a systematic approach to evaluate whether fixed Hessian matrices can be used for efficient computation.
+    We propose a systematic framework to determine the existence of constant Hessian approximations in fixed-Hessian methods (Böhning & Lindsay, 1988; Bonte & Vercauteren, 2018), and to assess the feasibility of employing fixed Hessian matrices for efficient computation.
 2 Preliminaries
 We use the square brackets “[][\ ][ ]” to denote the index of a vector or matrix element in what follows.
-For example, for a vector $\bm{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\bm{v}[i]$ or $\bm{v}_{[i]}$ means the $i$-th element of vector $\bm{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
+For example, for a vector $\boldsymbol{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\boldsymbol{v}[i]$ or $\boldsymbol{v}_{[i]}$ means the $i$-th element of vector $\boldsymbol{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
 2.1 Fully Homomorphic Encryption
 Fully Homomorphic Encryption (FHE) is a type of cryptographic scheme that can be used to compute an arbitrary number of additions and multiplications directly on the encrypted data.
-It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation (Gentry, 2009).
+It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation (2009).
 FHE schemes themselves are computationally time-consuming; the choice of dataset encoding matters likewise to the efficiency.
@@
 HEAAN supports a parallel technique (aka SIMD) to pack multiple complex numbers in a single polynomial and provides rotation operation on plaintext slots.
-The underlying HE scheme in HEAAN is well described in (Kim et al., 2018a; b; Han et al., 2019).
+The underlying HE scheme in HEAAN is well described in (Kim et al., 2018a, b; Han et al., 2019).
 2.2 Database Encoding Method
@@
 (2018a) proposed an efficient and promising database-encoding method by using SIMD technique, which could make full use of the computation and storage resources.
-Suppose a database contains a training dataset $Z$ comprising $n$ samples and $(1+d)$ covariates.
-The dataset matrix $Z$ is packed into a single ciphertext in a row-by-row manner.
+Suppose that a database has a training dataset consisting of $n$ samples with $(1+d)$ covariates, they packed the training dataset $Z$ into a single ciphertext in a row-by-row manner.
 When employing this encoding scheme, we can manipulate the data matrix $Z$ through HE operations on the ciphertext $Enc[Z]$, utilizing only three HE operations - rotation, addition, and multiplication.
@@
 Logistic regression is widely used in binary classification tasks to infer whether a binary-valued variable belongs to a certain class or not.
-LR can be generalized from linear regression (Murphy, 2012) by mapping the whole real line $(\bm{\beta}^{\top}\mathbf{x})$ to $(0,1)$ via the sigmoid function $\sigma(z)=1/(1+\exp(-z))$, where the vector $\bm{\beta}\in\mathbb{R}^{(1+d)}$ is the main parameter of LR and the vector $\mathbf{x}=(1,x_{1},\ldots,x_{d})\in\mathbb{R}^{(1+d)}$ the input covariate.
+LR can be generalized from linear regression (Murphy, 2012) by mapping the whole real line $(\boldsymbol{\beta}^{\top}\mathbf{x})$ to $(0,1)$ via the sigmoid function $\sigma(z)=1/(1+\exp(-z))$, where the vector $\boldsymbol{\beta}\in\mathbb{R}^{(1+d)}$ is the main parameter of LR and the vector $\mathbf{x}=(1,x_{1},\ldots,x_{d})\in\mathbb{R}^{(1+d)}$ the input covariate.
 Thus logistic regression can be formulated with the class label $y\in\{\pm 1\}$ as follows:
-$\Pr(y=+1|\mathbf{x},\bm{\beta})$ $=\sigma(\bm{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{-\bm{\beta}^{\top}\mathbf{x}}},$
-$\Pr(y=-1|\mathbf{x},\bm{\beta})$ $=1-\sigma(\bm{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{+\bm{\beta}^{\top}\mathbf{x}}}.$
+$\Pr(y=+1|\mathbf{x},\boldsymbol{\beta})$ $=\sigma(\boldsymbol{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{-\boldsymbol{\beta}^{\top}\mathbf{x}}},$
+$\Pr(y=-1|\mathbf{x},\boldsymbol{\beta})$ $=1-\sigma(\boldsymbol{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{+\boldsymbol{\beta}^{\top}\mathbf{x}}}.$
 LR sets a threshold (usually $0.5$) and compares its output with it to decide the resulting class label.
-The logistic regression problem can be transformed into an optimization problem that seeks a parameter $\bm{\beta}$ to maximize $L(\bm{\beta})=\prod_{i=1}^{n}\Pr(y_{i}|\mathbf{x}_{i},\bm{\beta})$ or its log-likelihood function $l(\bm{\beta})$ for convenience in the calculation:
-$l(\bm{\beta})=\ln L(\bm{\beta})=-\sum_{i=1}^{n}\ln(1+e^{-y_{i}\bm{\beta}^{\top}\mathbf{x}_{i}}),$
+The logistic regression problem can be transformed into an optimization problem that seeks a parameter $\boldsymbol{\beta}$ to maximize $L(\boldsymbol{\beta})=\prod_{i=1}^{n}\Pr(y_{i}|\mathbf{x}_{i},\boldsymbol{\beta})$ or its log-likelihood function $l(\boldsymbol{\beta})$ for convenience in the calculation:
+$l(\boldsymbol{\beta})=\ln L(\boldsymbol{\beta})=-\sum_{i=1}^{n}\ln(1+e^{-y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}}),$
 where $n$ is the number of examples in the training dataset.
-LR does not have a closed form of maximizing $l(\bm{\beta})$ and two main methods are adopted to estimate the parameters of an LR model: (a) gradient descent method via the gradient; and (b) Newton’s method by the Hessian matrix.
-The gradient and Hessian of the log-likelihood function $l(\bm{\beta})$ are given by, respectively:
-$\nabla_{\bm{\beta}}l(\bm{\beta})$ $=\sum_{i}(1-\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i}))y_{i}\mathbf{x}_{i},$
-$\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$ $=\sum_{i}(y_{i}\mathbf{x}_{i})(\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i})(y_{i}\mathbf{x}_{i})$
+LR does not have a closed form of maximizing $l(\boldsymbol{\beta})$ and two main methods are adopted to estimate the parameters of an LR model: (a) gradient descent method via the gradient; and (b) Newton’s method by the Hessian matrix.
+The gradient and Hessian of the log-likelihood function $l(\boldsymbol{\beta})$ are given by, respectively:
+$\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$ $=\sum_{i}(1-\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}))y_{i}\mathbf{x}_{i},$
+$\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$ $=\sum_{i}(y_{i}\mathbf{x}_{i})(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})(y_{i}\mathbf{x}_{i})$
 $=X^{\top}SX,$
-where $S$ is a diagonal matrix with entries $S_{ii}=(\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i})$ and $X$ the dataset.
-The log-likelihood function $l(\bm{\beta})$ of LR has at most a unique global maximum (Allison, 2008), where its gradient is zero.
-Newton’s method is a second-order technique to numerically find the roots of a real-valued differentiable function, and thus can be used to solve the $\bm{\beta}$ in $\nabla_{\bm{\beta}}l(\bm{\beta})=0$ for LR.
+where $S$ is a diagonal matrix with entries $S_{ii}=(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})$ and $X$ the dataset.
+The log-likelihood function $l(\boldsymbol{\beta})$ of LR has at most a unique global maximum (Allison, 2008), where its gradient is zero.
+Newton’s method is a second-order technique to numerically find the roots of a real-valued differentiable function, and thus can be used to solve the $\boldsymbol{\beta}$ in $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})=0$ for LR.
 3 Technical Details
@@
 This novel technique is called the fixed Hessian Newton’s method.
-Böhning & Lindsay (1988) have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\bm{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})-\bar{H}$ is non-negative definite.
+Böhning and Lindsay (1988) have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\boldsymbol{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})-\bar{H}$ is non-negative definite.
 With such a fixed Hessian matrix $\bar{H}$, the iteration for Newton’s method can be simplified to:
-$\bm{\beta}_{t+1}=\bm{\beta}_{t}-\bar{H}^{-1}\nabla_{\bm{\beta}}l(\bm{\beta}).$
-Böhning and Lindsay also suggest the fixed matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ is a good lower bound for the Hessian of the log-likelihood function $l(\bm{\beta})$ in LR.
+$\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-\bar{H}^{-1}\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}).$
+Böhning and Lindsay also suggest the fixed matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ is a good lower bound for the Hessian of the log-likelihood function $l(\boldsymbol{\beta})$ in LR.
 3.1 Simplified Fixed Hessian
-Bonte & Vercauteren (2018) simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain.
+Bonte and Vercauteren (2018) simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain.
 They replace the matrix $\bar{H}$ with a diagonal matrix $B$ whose diagonal elements are simply the sums of each row in $\bar{H}$.
@@
 Their simplified fixed Hessian method can be formulated as follows:
-$\bm{\beta}_{t+1}$ $=\bm{\beta}_{t}-B^{-1}\cdot\nabla_{\bm{\beta}}l(\bm{\beta}),$
-$=\bm{\beta}_{t}-\left[\begin{array}[]{cccc}b_{00}&0&\ldots&0\\ 0&b_{11}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&b_{dd}\\ \end{array}\right]\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\bm{\beta}_{t}-\left[\begin{array}[]{c}b_{00}\cdot\nabla_{0}\\ b_{11}\cdot\nabla_{1}\\ \vdots\\ b_{dd}\cdot\nabla_{d}\\ \end{array}\right],$
-where $b_{ii}$ is the reciprocal of $\sum_{i=0}^{d}\bar{h}_{0i}$ and $\nabla_{i}$ is the element of $\nabla_{\bm{\beta}}l(\bm{\beta})$.
+$\boldsymbol{\beta}_{t+1}$ $=\boldsymbol{\beta}_{t}-B^{-1}\cdot\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}),$
+$=\boldsymbol{\beta}_{t}-\left[\begin{array}[]{cccc}b_{00}&0&\ldots&0\\ 0&b_{11}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&b_{dd}\\ \end{array}\right]\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]$
+$=\boldsymbol{\beta}_{t}-\left[\begin{array}[]{c}b_{00}\cdot\nabla_{0}\\ b_{11}\cdot\nabla_{1}\\ \vdots\\ b_{dd}\cdot\nabla_{d}\\ \end{array}\right],$
+where $b_{ii}$ is the reciprocal of $\sum_{i=0}^{d}\bar{h}_{0i}$ and $\nabla_{i}$ is the element of $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$.
 Consider a special situation: if all the elements $b_{00},\ldots,b_{dd}$ had the same value $-\eta$ with $\eta>0$, the iterative formula of the SFH method could be given as:
-$\bm{\beta}_{t+1}$ $=\bm{\beta}_{t}-(-\eta)\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\bm{\beta}_{t}+\eta\cdot\nabla_{\bm{\beta}}l(\bm{\beta}),$
+$\boldsymbol{\beta}_{t+1}$ $=\boldsymbol{\beta}_{t}-(-\eta)\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\boldsymbol{\beta}_{t}+\eta\cdot\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}),$
 which is the same as the formula of the naive gradient $ascent$ method.
 Such a coincidence not only helps generate the idea behind this work but also leads us to believe that there is a connection between the Hessian matrix and the learning rate of the gradient (descent) method.
-We regard $B^{-1}\cdot\nabla_{\bm{\beta}}$ as a novel enhanced gradient variant and allocate a distinct learning rate to it.
+We regard $B^{-1}\cdot\nabla_{i}$ as a novel enhanced gradient variant and allocate a distinct learning rate to it.
 As long as we ensure that this new learning rate decreases from a positive floating-point number greater than 1 (such as 2) to 1 in a bounded number of iteration steps, the fixed Hessian Newton’s method guarantees the algorithm will converge eventually.
-The SFH method proposed by Bonte & Vercauteren (2018) has two limitations: (a) in the construction of the simplified fixed Hessian, all entries in the symmetric matrix $\bar{H}$ need to be non-positive.
+The SFH method proposed by Bonte and Vercauteren (2018) has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive.
 For machine learning applications, datasets are typically normalized in advance to the range [0,1], satisfying the convergence condition of the SFH method.
-However, in other cases, such as numerical optimization, this condition may not always hold; and (b) the simplified fixed Hessian matrix $B$, as well as the fixed Hessian $\bar{H}=-\frac{1}{4}X^{\top}X$ , can still be singular, especially when the dataset is a high-dimensional sparse matrix, such as the MNIST datasets.
-We extend SFH by removing these limitations to generalize this simplified fixed Hessian to be invertible in any case and propose a faster gradient variant, which we term $quadratic$ $gradient$.
+However, in other cases, such as numerical optimization, this condition may not always hold; and (b) the simplified fixed Hessian matrix $B$, as well as the fixed Hessian matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ , can still be singular, especially when the dataset is a high-dimensional sparse matrix, such as the MNIST datasets.
+We extend their work by removing these limitations so as to generalize this simplified fixed Hessian to be invertible in any case and propose a faster gradient variant, which we term quadratic gradient.
 3.2 Quadratic Gradient Definition
-Suppose that a differentiable scalar-valued function $F(\mathbf{x})$ has its gradient $\bm{g}$ and Hessian matrix $H$, with any matrix $\bar{H}\leq H$ in the Loewner ordering for a $maximization$ problem as follows:
-$\bm{g}=\left[\begin{array}[]{c}g_{0}\\ g_{1}\\ \vdots\\ g_{d}\\ \end{array}\right],$ $H=\left[\begin{array}[]{cccc}\nabla_{00}^{2}&\nabla_{01}^{2}&\ldots&\nabla_{0d}^{2}\\ \nabla_{10}^{2}&\nabla_{11}^{2}&\ldots&\nabla_{1d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \nabla_{d0}^{2}&\nabla_{d1}^{2}&\ldots&\nabla_{dd}^{2}\\ \end{array}\right],\quad\bar{H}$ $=\left[\begin{array}[]{cccc}\bar{h}_{00}&\bar{h}_{01}&\ldots&\bar{h}_{0d}\\ \bar{h}_{10}&\bar{h}_{11}&\ldots&\bar{h}_{1d}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{h}_{d0}&\bar{h}_{d1}&\ldots&\bar{h}_{dd}\\ \end{array}\right],$
+Suppose that a differentiable scalar-valued function $F(\mathbf{x})$ has its gradient $\boldsymbol{g}$ and Hessian matrix $H$, with any matrix $\bar{H}\leq H$ in the Loewner ordering for a maximization problem as follows:
+$\boldsymbol{g}=\left[\begin{array}[]{c}g_{0}\\ g_{1}\\ \vdots\\ g_{d}\\ \end{array}\right],$ $H=\left[\begin{array}[]{cccc}\nabla_{00}^{2}&\nabla_{01}^{2}&\ldots&\nabla_{0d}^{2}\\ \nabla_{10}^{2}&\nabla_{11}^{2}&\ldots&\nabla_{1d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \nabla_{d0}^{2}&\nabla_{d1}^{2}&\ldots&\nabla_{dd}^{2}\\ \end{array}\right],\quad$
+$\bar{H}$ $=\left[\begin{array}[]{cccc}\bar{h}_{00}&\bar{h}_{01}&\ldots&\bar{h}_{0d}\\ \bar{h}_{10}&\bar{h}_{11}&\ldots&\bar{h}_{1d}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{h}_{d0}&\bar{h}_{d1}&\ldots&\bar{h}_{dd}\\ \end{array}\right],$
 where $\nabla_{ij}^{2}=\nabla_{ji}^{2}=\frac{\partial^{2}F}{\partial x_{i}\partial x_{j}}$.
@@
 where $\epsilon$ is a small positive constant to avoid division by zero (usually set to $81e-$ 8).
-As long as $\tilde{B}$ satisfies the convergence condition of the aforementioned fixed Hessian method, namely $\tilde{B}\leq H$, we can use this approximation $\tilde{B}$ of the Hessian matrix as a lower bound.
-Given that we have already assumed $\bar{H}\leq H$, it suffices to demonstrate that $\tilde{B}\leq\bar{H}$.
-We establish $\tilde{B}\leq\bar{H}$ in a manner similar to that employed by Bonte & Vercauteren (2018).
+As long as $\tilde{B}$ satisfies the convergence condition of the above fixed Hessian method, $\tilde{B}\leq H$, we can use this approximation $\tilde{B}$ of the Hessian matrix as a lower bound.
+Since we already assume that $\bar{H}\leq H$, it will suffice to show that $\tilde{B}\leq\bar{H}$.
+We prove $\tilde{B}\leq\bar{H}$ in a similar way that (Bonte & Vercauteren, 2018) did.
 Lemma 3.1.
@@
 Definition 3.2 (Quadratic Gradient).
-Given the previously defined $\tilde{B}$, we define the quadratic gradient as $G=\bar{B}\cdot\bm{g}$ with a new learning rate $N_{t}$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$.
+Given the previously defined $\tilde{B}$, we define the quadratic gradient as $G=\bar{B}\cdot\boldsymbol{g}$ with a new learning rate $N_{t}$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$.
 The learning rate $N_{t}$ should always be no less than 1 and is designed to decrease to 1 over a limited number of iterations.
-It is important to note that $G$ remains a column vector of the same dimension as the gradient $\bm{g}$.
+It is important to note that $G$ remains a column vector of the same dimension as the gradient $\boldsymbol{g}$.
 To maximize the function $F(\mathbf{x})$, we can employ the iterative formula: $\mathbf{x}_{t+1}=\mathbf{x}_{t}+N_{t}\cdot G$, similar to the naive gradient approach.
@@
 It is worth noting that $\bar{H}$ can be taken as the Hessian $H$ itself, thereby eliminating the need to find either lower or upper bounds.
-Böhning & Lindsay (1988) did not propose a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method.
+Böhning and Lindsay (1988) did not propose a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method.
 This may be due to the absence of such matrices for most objective functions.
+In this work, we present a systematic approach for identifying fixed Hessian matrices when they exist.
+Our detection method is based on a necessary condition—if it fails to identify a valid fixed Hessian matrix, it does not rule out the existence of one; instead, it indicates that no such matrix can be constructed through our particular framework.
 We note that our quadratic gradient approach does not require a constant (fixed) Hessian replacement and instead allows for the use of the original Hessian itself to construct our varying diagonal matrix $\tilde{B}$ for the quadratic gradient.
@@
 Algorithm 1 The Enhanced Nesterov’s Accelerated Gradient Algorithm
-1:training dataset $X\in\mathbb{R}^{n\times(1+d)}$; training label $Y\in\mathbb{R}^{n\times 1}$; learning rate $lr\in\mathbb{R}$; and the number $\kappa$ of iterations;
-2:the parameter vector $V\in\mathbb{R}^{(1+d)}$
-3:Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$, $V\leftarrow\bm{0}$, $W\leftarrow\bm{0}$, $\bar{B}\leftarrow\bm{0}$
-4:for $i:=0$ to $d$ do $\triangleright$ $\epsilon$ is a small positive constant such as $81e-$ 8
-5:     $\bar{B}[i][i]\leftarrow\epsilon$
-6:     for $j:=0$ to $d$ do
-7:         $\bar{B}[i][i]\leftarrow\bar{B}[i][i]+|\bar{H}[i][j]|$
-8:     end for
-9:end for
-10:Set $alpha_{0}\leftarrow 0.01$, $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
-11:for $count:=1$ to $\kappa$ do
-12:     Set $Z\leftarrow\bm{0}$triangleright $Z\in\mathbb{R}^{n}$ will store the inputs for Sigmoid function
-13:     for $i:=1$ to $n$ do
-14:         for $j:=0$ to $d$ do
-15:              $Z[i]\leftarrow Z[i]+Y[i]\times V[j]\times X[i][j]$
-16:         end for
-17:     end for
-18:     Set $\bm{\sigma}\leftarrow\bm{0}$triangleright $\bm{\sigma}\in\mathbb{R}^{n}$ will store the outputs of Sigmoid function
-19:     for $i:=1$ to $n$ do
-20:         $\bm{\sigma}[i]\leftarrow 1/(1+\exp(-Z[i]))$
+1:  Require: training dataset $X\in\mathbb{R}^{n\times(1+d)}$; training label $Y\in\mathbb{R}^{n\times 1}$; learning rate $lr\in\mathbb{R}$; and the number $\kappa$ of iterations;
+2:  Ensure: the parameter vector $V\in\mathbb{R}^{(1+d)}$
+3:  Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$ $\bar{H}\in\mathbb{R}^{(1+d)\times(1+d)}$
+4:  Set $V\leftarrow\boldsymbol{0}$, $W\leftarrow\boldsymbol{0}$, $\bar{B}\leftarrow\boldsymbol{0}$ $V\in\mathbb{R}^{(1+d)}$, $W\in\mathbb{R}^{(1+d)}$, $\bar{B}\in\mathbb{R}^{(1+d)\times(1+d)}$
+5:  for $i:=0$ to $d$ do
+6:     # $\epsilon$ is a small positive constant such as $81e-$ 8
+7:     $\bar{B}[i][i]\leftarrow\epsilon$
+8:     for $j:=0$ to $d$ do
+9:        $\bar{B}[i][i]\leftarrow\bar{B}[i][i]+|\bar{H}[i][j]|$
+10:     end for
+11:  end for
+12:  Set $alpha_{0}\leftarrow 0.01$
+13:  Set $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
+14:  for $count:=1$ to $\kappa$ do
+15:     # $Z\in\mathbb{R}^{n}$ will store the inputs for Sigmoid function
+16:     Set $Z\leftarrow\boldsymbol{0}$
+17:     for $i:=1$ to $n$ do
+18:        for $j:=0$ to $d$ do
+19:           $Z[i]\leftarrow Z[i]+Y[i]\times V[j]\times X[i][j]$
+20:        end for
 21:     end for
-22:     Set $\bm{g}\leftarrow\bm{0}$
-23:     for $j:=0$ to $d$ do
+22:     # $\boldsymbol{\sigma}\in\mathbb{R}^{n}$ will store the outputs of Sigmoid function
+23:     Set $\boldsymbol{\sigma}\leftarrow\boldsymbol{0}$
 24:     for $i:=1$ to $n$ do
-25:              $\bm{g}[j]\leftarrow\bm{g}[j]+(1-\bm{\sigma}[i])\times Y[i]\times X[i][j]$
+25:        $\boldsymbol{\sigma}[i]\leftarrow 1/(1+\exp(-Z[i]))$
 26:     end for
-27:     end for
-28:     Set $G\leftarrow\bm{0}$
-29:     for $j:=0$ to $d$ do
-30:         $G[j]\leftarrow\bar{B}[j][j]\times\bm{g}[j]$
+27:     Set $\boldsymbol{g}\leftarrow\boldsymbol{0}$
+28:     for $j:=0$ to $d$ do
+29:        for $i:=1$ to $n$ do
+30:           $\boldsymbol{g}[j]\leftarrow\boldsymbol{g}[j]+(1-\boldsymbol{\sigma}[i])\times Y[i]\times X[i][j]$
 31:        end for
-32:     Set $\eta\leftarrow(1-alpha_{0})/alpha_{1}$, $\gamma\leftarrow(\frac{10}{1+t})/n$
-33:     for $j:=0$ to $d$ do
-34:         $w_{temp}\leftarrow V[j]+(1+\gamma)\times G[j]$, $V[j]\leftarrow(1-\eta)\times w_{temp}+\eta\times W[j]$, $W[j]\leftarrow w_{temp}$
-35:     end for
-36:     $alpha_{0}\leftarrow alpha_{1}$, $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
-37:end for
-38:Return $V$
+32:     end for
+33:     Set $G\leftarrow\boldsymbol{0}$
+34:     for $j:=0$ to $d$ do
+35:        $G[j]\leftarrow\bar{B}[j][j]\times\boldsymbol{g}[j]$
+36:     end for
+37:     Set $\eta\leftarrow(1-alpha_{0})/alpha_{1}$
+38:     Set $\gamma\leftarrow lr/(n\times count)$
+39:     for $j:=0$ to $d$ do
+40:        $w_{temp}\leftarrow V[j]+(1+\gamma)\times G[j]$
+41:        $V[j]\leftarrow(1-\eta)\times w_{temp}+\eta\times W[j]$
+42:        $W[j]\leftarrow w_{temp}$
+43:     end for
+44:     $alpha_{0}\leftarrow alpha_{1}$
+45:     $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
+46:  end for
+47:  Return $V$
 3.3 Quadratic Gradient Algorithms
@@
     NAG is a different variant of the momentum method to give the momentum term much more prescience. The iterative formulas of the gradient $ascent$ method for NAG are as follows:
-    $V_{t+1}$ $=\bm{\beta}_{t}+\eta_{t}\cdot\nabla J(\bm{\beta}_{t}),$ (3)
-    $\bm{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
-    where $V_{t+1}$ is the intermediate variable used for updating the final weight $\bm{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position (Kim et al., 2018a).
-The enhanced NAG is to replace equation 3 with $V_{t+1}=\bm{\beta}_{t}+N_{t}\cdot G,$ where $N_{t}$ is the new learning rate for the enhanced algorithm usually set to $1+\eta_{t}$.
+    $V_{t+1}$ $=\boldsymbol{\beta}_{t}+\eta_{t}\cdot\nabla J(\boldsymbol{\beta}_{t}),$ (3)
+    $\boldsymbol{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
+    where $V_{t+1}$ is the intermediate variable used for updating the final weight $\boldsymbol{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position (Kim et al., 2018a).
+The enhanced NAG is to replace (3) with $V_{t+1}=\boldsymbol{\beta}_{t}+N_{t}\cdot G,$ where $N_{t}$ is the new learning rate for the enhanced NAG algorithm usually set to $1+\eta_{t}$.
 Our enhanced NAG is described in Algorithm 1 .
@@
     Adagrad is a gradient-based algorithm suitable for dealing with sparse data.
-The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\bm{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
-    $\bm{\beta}_{[i]}^{(t+1)}$ $=\bm{\beta}_{[i]}^{(t)}-\frac{\eta_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}\bm{g}_{[i]}^{(t)}\cdot\bm{g}_{[i]}^{(t)}}}\cdot\bm{g}_{[i]}^{(t)},$
-    $\bm{\beta}_{[i]}^{(t+1)}$ $=\bm{\beta}_{[i]}^{(t)}-\frac{N_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
+The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\boldsymbol{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
+    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{\eta_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}\boldsymbol{g}_{[i]}^{(t)}\cdot\boldsymbol{g}_{[i]}^{(t)}}}\cdot\boldsymbol{g}_{[i]}^{(t)},$
+    $and$
+    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{N_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
   3. 3.
     Adam is an optimization algorithm that combines the benefits of Adagrad, which adapts learning rates for each parameter, with momentum, allowing it to converge faster and handle sparse gradients more effectively.
-Similar to how Adagrad transitions to its enhanced version, the update formula for enhanced Adam, also employs a new learning rate $N_{t}$ and replaces the gradient with its quadratic form.
+Similar to how Adagrad transitions to its enhanced version, the update formula for enhanced Adam, also employs a new learning rate and replaces the gradient with its quadratic form.
 (a) The iDASH dataset
-(b) The edin dataset
+(b) The Edinburgh dataset
 (c) The lbw dataset
@@
 (f) The uis dataset
-(g) The MNIST dataset
-(h) The financial dataset
-Figure 1: The training results of NAG and Enhanced NAG in the clear.
+(g) restructured MNIST dataset
+(h) The private financial dataset
+Figure 1: The training results of NAG and Enhanced NAG in the clear
 (a) The iDASH dataset
-(b) The edin dataset
+(b) The Edinburgh dataset
 (c) The lbw dataset
@@
 (f) The uis dataset
-(g) The MNIST dataset
-(h) The financial dataset
-Figure 2: The training results of Adagrad and Enhanced Adagrad in the clear.
+(g) restructured MNIST dataset
+(h) The private financial dataset
+Figure 2: The training results of Adagrad and Enhanced Adagrad in the clear
 (a) The iDASH dataset
-(b) The edin dataset
+(b) The Edinburgh dataset
 (c) The lbw dataset
@@
 (f) The uis dataset
-(g) The MNIST dataset
-(h) The financial dataset
-Figure 3: The training results of Adam and Enhanced Adam in the clear.
+(g) restructured MNIST dataset
+(h) The private financial dataset
+Figure 3: The training results of Adam and Enhanced Adam in the clear
 Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM.
-Given that our focus is on the convergence speed of the algorithms during training, we use the loss function $l(\bm{\beta})$, specifically maximum likelihood estimation (MLE), as the only indicator.
+Given that our focus is on the convergence speed of the algorithms during training, we use the loss function $l(\boldsymbol{\beta})$, specifically maximum likelihood estimation (MLE), as the only indicator.
 We evaluate six algorithms—NAG, Adagrad, Adam, and their quadratic-gradient variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam, respectively)—using the datasets adopted by Kim et al.
@@
 (2018a), the enhanced NAG algorithm employs the learning rate $1+\frac{10}{1+t}$, consistent with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate.
-The enhanced Adagrad algorithm utilizes $N_{t}=1+0.01$ as the learning rate, while the enhanced Adam algorithm adopts the following learning rate settings: $\alpha=1+0.001$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$.
+The enhanced Adagrad algorithm utilizes $N_{t}=0.06+0.01$ as the learning rate, while the enhanced Adam algorithm adopts the following learning rate settings: $\alpha=0.01+0.001$ (the new learning rate), $\beta_{1}=0.9$, and $\beta_{2}=0.999$.
 Refer to (Kingma & Ba, 2014) for further details regarding these parameters.
-In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\bm{\beta})$.
-From the empirical results presented in Figures 12, and 3, we conclude that all the enhanced algorithms converge, and that the enhanced NAG algorithm consistently outperforms the original NAG method in terms of convergence rate.
-Although the enhanced Adagrad and Adam methods do not achieve faster convergence than their original counterparts on the iDASH genomic dataset, they demonstrate clear advantages in all other cases.
-The reason for this discrepancy is beyond the scope of this paper and will be addressed as part of future work.
-One possible explanation for the superior performance of the enhanced algorithms is that the quadratic gradient integrates curvature information into first-order gradient methods.
+In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\boldsymbol{\beta})$.
+From the empirical results presented in Figures 21, and 3, we conclude that all the enhanced algorithms converge and consistently outperforms the original NAG method in terms of convergence rate.
 An important observation is that the enhanced algorithms demonstrate better performance with learning rates between $1$ and $2$ compared to other values.
@@
 For quadratic gradient algorithms, we recommend employing an exponentially decaying learning rate, such as $1+A\cdot\gamma^{t}$, where $t$ denotes the iteration number, $A$ is a positive constant typically set to no less than $1$, and $\gamma$ is a positive number less than $1$ that controls the rate of decay.
-Gradient And Quadratic Gradient We executed the raw gradient ascent algorithm and the raw quadratic gradient algorithm on the lbw dataset using various learning rates.
-Figure 4 presents the detailed results of this experiment.
-It was precisely these results that directly inspired the authors to develop the concept of the quadratic gradient.
-Similar to the gradient, the quadratic gradient, which incorporates curvature, also demonstrates smooth progression.
-In particular, as the learning rate is gradually increased, both the gradient and the quadratic gradient reveal corresponding gradual changes in performance, rather than sudden jumps.
-(a) The raw first-order gradient ascent algorithm with iterative formulas: $\bm{\beta}_{t+1}=\bm{\beta}_{t}+\mathrm{lr}\cdot\bm{g}$
-(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\bm{\beta}_{t+1}=\bm{\beta}_{t}+\mathrm{LR}\cdot G$
-Figure 4: The training outcomes for both the raw gradient algorithm and the raw quadratic gradient ascent algorithm in the clear setting, conducted on the lbw dataset.
+The reasons why the enhanced algorithms have a better performance might be that quadratic gradient can incorporate curves into first-order gradient algorithms.
+(a) The raw first-order gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{lr}\cdot\boldsymbol{g}$
+(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{LR}\cdot G$
+Figure 4: The training outcomes for both the raw gradient algorithm and the raw quadratic gradient ascent algorithm in the clear setting, conducted on the lbw dataset
+Gradient And Quadratic Gradient We executed the raw gradient ascent algorithm and raw quadratic gradient algorithm on the lbw dataset using various learning rates.
+Figure 4 shows the detailed results of this experiment.
+It is precisely these results that inspired the present authors to come up with the idea of quadratic gradient.
+Similar to the gradient, the quadratic gradient carrying the curve also exhibits the characteristic of smooth progression.
+That is, when the learning rate is gradually increased, both the gradient and the quadratic gradient will show corresponding gradual changes in performance rather than abrupt shifts.
 4 Secure Training
+Adagrad method is not a practical solution for homomorphic LR due to its frequent inversion operations.
+It seems plausible that the enhanced NAG is likely the reasonable choice for homomorphic LR training.
 We adopt the enhanced NAG method to implement secure logistic regression training based on HE.
@@
 We leave the computation of matrix $\bar{B}$ to data owner and let the data owner upload the ciphertext encrypting the $\bar{B}$ to the cloud.
-Since data owner has to prepare the dataset and normalize it, it would also be practical for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
+Since data owner has to prepare the dataset and normalize it, it would also be practicable for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
 Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model.
@@
 The ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ are as follows:
-$\text{X}=\left[\begin{array}[]{cccc}1&x_{11}&\ldots&x_{1d}\\ 1&x_{21}&\ldots&x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n1}&\ldots&x_{nd}\\ \end{array}\right],\text{Y}=\left[\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\\ \end{array}\right],$ $\text{ct}_{Z}=Enc\left[\begin{array}[]{cccc}y_{1}&y_{1}x_{11}&\ldots&y_{1}x_{1d}\\ y_{2}&y_{2}x_{21}&\ldots&y_{2}x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ y_{n}&y_{n}x_{n1}&\ldots&y_{n}x_{nd}\\ \end{array}\right],$
-$\text{ct}_{\beta}^{(0)}=Enc\left[\begin{array}[]{cccc}\beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \vdots&\vdots&\ddots&\vdots\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \end{array}\right],$ $\text{ct}_{\bar{B}}=Enc\left[\begin{array}[]{cccc}\bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \end{array}\right],$
+X $=\left[\begin{array}[]{cccc}1&x_{11}&\ldots&x_{1d}\\ 1&x_{21}&\ldots&x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n1}&\ldots&x_{nd}\\ \end{array}\right],\text{Y}=\left[\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\\ \end{array}\right],$
+$\text{ct}_{Z}$ $=Enc\left[\begin{array}[]{cccc}y_{1}&y_{1}x_{11}&\ldots&y_{1}x_{1d}\\ y_{2}&y_{2}x_{21}&\ldots&y_{2}x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ y_{n}&y_{n}x_{n1}&\ldots&y_{n}x_{nd}\\ \end{array}\right],$
+$\text{ct}_{\beta}^{(0)}$ $=Enc\left[\begin{array}[]{cccc}\beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \vdots&\vdots&\ddots&\vdots\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \end{array}\right],$
+$\text{ct}_{\bar{B}}$ $=Enc\left[\begin{array}[]{cccc}\bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \end{array}\right],$
 where $\bar{B}_{[i][i]}$ is the diagonal element of $\bar{B}$ that is built from $-\frac{1}{4}X^{\top}X$.
-The public cloud takes the three ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ and evaluates the enhanced NAG algorithm to find a decent weight vector by updating the vector $\text{ct}_{\beta}^{(0)}$.
+The pulbic cloud takes the three ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ and evaluates the enhanced NAG algorithm to find a decent weight vector by updating the vector $\text{ct}_{\beta}^{(0)}$.
 Refer to (Kim et al., 2018a) for a detailed description about how to calculate the gradient by HE programming.
@@
 In addition, the data owner needs to upload one more ciphertext encrypting the matrix $\bar{B}$.
-However, the enhanced algorithm converges faster, and we believe it can compensate for the mentioned shortcomings.
+However, the enhanced algorithm converges faster, and we believe it can compensate for the mentioned drawbacks.
 Table 1: Implementation Results for iDASH datasets with 10-fold CV
@@
 Num
-Method
+Method deg $g$
 Iter
@@
 AUC
-iDASH 1579 18 Ours 4 0.08 4.43 61.46 0.696
-Baseline 7 0.04 6.07 62.87 0.689
+iDASH 1579 18 Ours 5 4 0.08 4.43 61.46 0.696
+(Kim et al., 2018a) 5 7 0.04 6.07 62.87 0.689
 Table 2: Implementation Results for other datasets with 5-fold CV
@@
 Num
-Method
+Method deg $g$
 Iter
@@
 AUC
-Edin 1253 8 Ours 4 0.04 0.6 89.52 0.943
-Baseline 7 0.02 3.6 91.04 0.958
-lbw 189 8 Ours 4 0.04 0.6 71.35 0.667
-Baseline 7 0.02 3.3 69.19 0.689
-nhanes3 15649 15 Ours 4 0.31 4.5 79.23 0.637
-Baseline 7 0.16 7.3 79.22 0.717
-pcs 379 9 Ours 4 0.04 0.6 63.20 0.733
-Baseline 7 0.02 3.5 68.27 0.740
-uis 575 8 Ours 4 0.04 0.6 74.43 0.597
-Baseline 7 0.02 3.5 74.44 0.603
+Edinburgh 1253 9 Ours 5 4 0.04 0.6 89.52 0.943
+(Kim et al., 2018a) 5 7 0.02 3.6 91.04 0.958
+lbw 189 9 Ours 5 4 0.04 0.6 71.35 0.667
+(Kim et al., 2018a) 5 7 0.02 3.3 69.19 0.689
+nhanes3 15649 15 Ours 5 4 0.31 4.5 79.23 0.637
+(Kim et al., 2018a) 5 7 0.16 7.3 79.22 0.717
+pcs 379 9 Ours 5 4 0.04 0.6 63.20 0.733
+(Kim et al., 2018a) 5 7 0.02 3.5 68.27 0.740
+uis 575 8 Ours 5 4 0.04 0.6 74.43 0.597
+(Kim et al., 2018a) 5 7 0.02 3.5 74.44 0.603
 5 Experiments
@@
 All the experiments on the ciphertexts were conducted on a public cloud with 32 vCPUs and 64 GB RAM.
-For a fair comparison with the baseline (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets.
+For a fair comparison with (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets.
 Like (Kim et al., 2018a), We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators.
 Tables 1 and 2 display the results of the two experiments, respectively.
-The two tables also provide the average evaluation running time for each iteration and the storage (encrypted dataset for the baseline work and encrypted dataset and $\bar{B}$ for our method).
+The two tables also provide the average evaluation running time and the storage (encrypted dataset for the baseline work and encrypted dataset and $\bar{B}$ for our method).
 We adopt the same packing method that Kim et al.
@@
 Yet despite only $4$ iterations, our enhanced NAG method still produces a comparable result.
+Our learning time measurement reflects the time required for 4 iterations of our enhanced algorithms while the baseline’s reported average running time includes time for evaluation and encryption.
+It is worth noting that we relied on their open-source code to implement our homomorphic logistic regression training.
+Therefore, theoretically, the running time for one iteration of our algorithm should be very similar to theirs.
 6 Conclusion
-In this paper, we proposed a faster gradient variant, termed quadratic gradient, and implemented the quadratic-gradient version of NAG in the encrypted domain to train a logistic regression model.
-The quadratic gradient introduced in this work can be directly constructed from the Hessian, effectively merging first-order gradient (descent) methods with the second-order Newton’s method.
-There is a promising potential for the quadratic gradient to accelerate other gradient methods, and it may serve as an alternative or even a replacement for the traditional line-search method.
+In this paper, we introduce a novel gradient variant, referred to as the quadratic gradient, and implement a quadratic-gradient-enhanced version of Nesterov’s Accelerated Gradient (NAG) to train a logistic regression model in the encrypted domain.
+The proposed quadratic gradient is constructed directly from the Hessian matrix, effectively bridging first-order gradient-based methods with second-order Newton-type techniques.
+This approach exhibits strong potential for accelerating a broad range of optimization algorithms.
+In particular, it offers a viable alternative to traditional line-search procedures, can be incorporated into adaptive methods such as enhanced AdaGrad, and may serve to accelerate Newton’s method via embedded gradient descent steps.
+Impact Statements
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v13

@@
 There is a great chance that the quadratic gradient approach could integrate first-order gradient descent/ascent algorithms with the second-order Newton-Raphson methods, and that quadratic gradient could be applied to a wide range of numerical optimization problems.
+Index Terms:
 Homomorphic Encryption, Logistic Regression, Quadratic Gradient, Simplified Fixed Hessian, Nesterov’s Accelerated Gradient, Gradient Ascent Algorithm
-1 Introduction
+I Introduction
 Given a person’s healthcare data related to a certain disease, we can train a logistic regression (LR) model capable of telling whether or not this person is likely to develop this disease.
@@
 Kim et al.
-(2018b) discussed the problem of performing LR training in an encrypted environment.
+[1] discussed the problem of performing LR training in an encrypted environment.
 They employed full batch gradient descent in the training process and utilized the least-squares method to approximate the sigmoid function.
-In the 2017 iDASH competition, Bonte and Vercauteren (2018), Kim et al.
-(2018a), Chen et al.
-(2018), and Crawford et al.
-(2018) all investigated the same problem that Kim et al.
-(2018b) studied.
+In the 2017 iDASH competition, Bonte and Vercauteren [2], Kim et al.
+[3], Chen et al.
+[4], and Crawford et al.
+[5] all investigated the same problem that Kim et al.
+[1] studied.
 In the iDASH competition of 2018, Kim et al.
-(2019b) and Blatt et al.
-(2019) further worked on it for an efficient packing and semi-parallel algorithm.
-There are other related works (Kim et al., 2019a; Ogilvie et al., 2020) focusing on various aspects but the papers most relevant to this work are (Bonte & Vercauteren, 2018) and (Kim et al., 2018a).
-Bonte and Vercauteren (2018) developed a practical algorithm called the simplified fixed Hessian (SFH) method.
+[6] and Blatt et al.
+[7] further worked on it for an efficient packing and semi-parallel algorithm.
+There are other related works [8, 9] focusing on various aspects but the papers most relevant to this work are [2] and [3].
+Bonte and Vercauteren [2] developed a practical algorithm called the simplified fixed Hessian (SFH) method.
 Our study extends their work and adopts the ciphertext packing technique proposed by Kim et al.
-(2018a) for efficient homomorphic computation.
+[3] for efficient homomorphic computation.
 Our specific contributions in this paper are as follows:
   1. 1.
-    We propose a new gradient variant, quadratic gradient, which can combine the first-order gradient algorithms and the second-order Newton-Raphson method (aka Newton’s method).
+    We introduce a novel gradient variant, termed quadratic gradient, that bridges first-order gradient-based algorithms and the second-order Newton–Raphson (i.e., Newton’s) method, thereby enabling a unified framework that leverages the strengths of both approaches.
   2. 2.
@@
   3. 3.
-    We implement privacy-preserving logistic regression training using the enhanced NAG method, to our best knowledge, which seems to be a great choice without compromising much on computation and storage.
+    We implement privacy-preserving logistic regression training using the enhanced Nesterov’s Accelerated Gradient (NAG) method, which, to the best of our knowledge, provides an effective balance between computation and storage without significant compromise.
   4. 4.
-    We propose a systematic framework to determine the existence of constant Hessian approximations in fixed-Hessian methods (Böhning & Lindsay, 1988; Bonte & Vercauteren, 2018), and to assess the feasibility of employing fixed Hessian matrices for efficient computation.
-2 Preliminaries
+    We propose a systematic framework to determine the existence of constant Hessian approximations in fixed-Hessian methods [10, 2], and to assess the feasibility of employing fixed Hessian matrices for efficient computation.
+II Preliminaries
 We use the square brackets “[][\ ][ ]” to denote the index of a vector or matrix element in what follows.
 For example, for a vector $\boldsymbol{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\boldsymbol{v}[i]$ or $\boldsymbol{v}_{[i]}$ means the $i$-th element of vector $\boldsymbol{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
-2.1 Fully Homomorphic Encryption
+II-A Fully Homomorphic Encryption
 Fully Homomorphic Encryption (FHE) is a type of cryptographic scheme that can be used to compute an arbitrary number of additions and multiplications directly on the encrypted data.
-It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation (2009).
+It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation [11].
 FHE schemes themselves are computationally time-consuming; the choice of dataset encoding matters likewise to the efficiency.
-In addition to these two limits, how to manage the magnitude of plaintext (Jäschke & Armknecht, 2016) also contributes to the slowdown.
+In addition to these two limits, how to manage the magnitude of plaintext [12] also contributes to the slowdown.
 Cheon et al.
-(2017) proposed a method to construct an HE scheme with a rescaling procedure which could eliminate this technical bottleneck effectively.
+[13] proposed a method to construct an HE scheme with a rescaling procedure which could eliminate this technical bottleneck effectively.
 We adopt their open-source implementation HEAAN while implementing our homomorphic LR algorithms.
@@
 HEAAN supports a parallel technique (aka SIMD) to pack multiple complex numbers in a single polynomial and provides rotation operation on plaintext slots.
-The underlying HE scheme in HEAAN is well described in (Kim et al., 2018a, b; Han et al., 2019).
-2.2 Database Encoding Method
+The underlying HE scheme in HEAAN is well described in [3, 1, 14].
+II-B Database Encoding Method
 Kim et al.
-(2018a) proposed an efficient and promising database-encoding method by using SIMD technique, which could make full use of the computation and storage resources.
-Suppose that a database has a training dataset consisting of $n$ samples with $(1+d)$ covariates, they packed the training dataset $Z$ into a single ciphertext in a row-by-row manner.
+[3] proposed an efficient and promising database encoding method leveraging the SIMD technique, which enables effective utilization of both computational and storage resources.
+Given a training dataset $Z$ consisting of $n$ samples with $(1+d)$ covariates each, the authors packed $Z$ into a single ciphertext in a row-wise manner, thereby maximizing the use of available ciphertext slots.
 When employing this encoding scheme, we can manipulate the data matrix $Z$ through HE operations on the ciphertext $Enc[Z]$, utilizing only three HE operations - rotation, addition, and multiplication.
@@
 Han et al.
-(2019) mentioned several basic but important operations used by Kim et al.
-(2018a) in their implementation, such as a procedure named “SumColVec” to compute the summation of the columns of a matrix.
+[14] mentioned several basic but important operations used by Kim et al.
+[3] in their implementation, such as a procedure named “SumColVec” to compute the summation of the columns of a matrix.
 With these fundamental operations, more intricate computations, like computing gradients in logistic regression models, become achievable.
-2.3 Logistic Regression Model
+II-C Logistic Regression Model
 Logistic regression is widely used in binary classification tasks to infer whether a binary-valued variable belongs to a certain class or not.
-LR can be generalized from linear regression (Murphy, 2012) by mapping the whole real line $(\boldsymbol{\beta}^{\top}\mathbf{x})$ to $(0,1)$ via the sigmoid function $\sigma(z)=1/(1+\exp(-z))$, where the vector $\boldsymbol{\beta}\in\mathbb{R}^{(1+d)}$ is the main parameter of LR and the vector $\mathbf{x}=(1,x_{1},\ldots,x_{d})\in\mathbb{R}^{(1+d)}$ the input covariate.
+LR can be generalized from linear regression [15] by mapping the whole real line $(\boldsymbol{\beta}^{\top}\mathbf{x})$ to $(0,1)$ via the sigmoid function $\sigma(z)=1/(1+\exp(-z))$, where the vector $\boldsymbol{\beta}\in\mathbb{R}^{(1+d)}$ is the main parameter of LR and the vector $\mathbf{x}=(1,x_{1},\ldots,x_{d})\in\mathbb{R}^{(1+d)}$ the input covariate.
 Thus logistic regression can be formulated with the class label $y\in\{\pm 1\}$ as follows:
@@
 where $S$ is a diagonal matrix with entries $S_{ii}=(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})$ and $X$ the dataset.
-The log-likelihood function $l(\boldsymbol{\beta})$ of LR has at most a unique global maximum (Allison, 2008), where its gradient is zero.
+The log-likelihood function $l(\boldsymbol{\beta})$ of LR has at most a unique global maximum [16], where its gradient is zero.
 Newton’s method is a second-order technique to numerically find the roots of a real-valued differentiable function, and thus can be used to solve the $\boldsymbol{\beta}$ in $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})=0$ for LR.
-3 Technical Details
-It is quite time-consuming to compute the Hessian matrix and its inverse in Newton’s method for each iteration.
-One way to limit this downside is to replace the varying Hessian with a fixed matrix $\bar{H}$.
-This novel technique is called the fixed Hessian Newton’s method.
-Böhning and Lindsay (1988) have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\boldsymbol{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})-\bar{H}$ is non-negative definite.
+III Technical Details
+Newton’s method requires computing the Hessian matrix and its inverse at every iteration, which is computationally intensive.
+To address this, the fixed Hessian Newton’s method (also known as the lower bound principle) replaces the varying Hessian with a constant matrix $\bar{H}$, offering improved efficiency.
+Böhning and Lindsay [10] have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\boldsymbol{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})-\bar{H}$ is non-negative definite.
 With such a fixed Hessian matrix $\bar{H}$, the iteration for Newton’s method can be simplified to:
@@
 Böhning and Lindsay also suggest the fixed matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ is a good lower bound for the Hessian of the log-likelihood function $l(\boldsymbol{\beta})$ in LR.
-3.1 Simplified Fixed Hessian
-Bonte and Vercauteren (2018) simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain.
+III-A Simplified Fixed Hessian
+Bonte and Vercauteren [2] further simplified the fixed Hessian $\bar{H}$ to facilitate its inversion in the encrypted domain.
 They replace the matrix $\bar{H}$ with a diagonal matrix $B$ whose diagonal elements are simply the sums of each row in $\bar{H}$.
@@
 Such a coincidence not only helps generate the idea behind this work but also leads us to believe that there is a connection between the Hessian matrix and the learning rate of the gradient (descent) method.
-We regard $B^{-1}\cdot\nabla_{i}$ as a novel enhanced gradient variant and allocate a distinct learning rate to it.
+We introduce $B^{-1}\cdot\nabla_{\boldsymbol{\beta}}\ell(\boldsymbol{\beta})$ as a novel enhanced gradient variant, and allocate a dedicated learning rate to better capture its dynamics.
 As long as we ensure that this new learning rate decreases from a positive floating-point number greater than 1 (such as 2) to 1 in a bounded number of iteration steps, the fixed Hessian Newton’s method guarantees the algorithm will converge eventually.
-The SFH method proposed by Bonte and Vercauteren (2018) has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive.
+The SFH method proposed by Bonte and Vercauteren [2] has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive.
 For machine learning applications, datasets are typically normalized in advance to the range [0,1], satisfying the convergence condition of the SFH method.
@@
 We extend their work by removing these limitations so as to generalize this simplified fixed Hessian to be invertible in any case and propose a faster gradient variant, which we term quadratic gradient.
-3.2 Quadratic Gradient Definition
-Suppose that a differentiable scalar-valued function $F(\mathbf{x})$ has its gradient $\boldsymbol{g}$ and Hessian matrix $H$, with any matrix $\bar{H}\leq H$ in the Loewner ordering for a maximization problem as follows:
+III-B Quadratic Gradient: Definition
+Suppose a differentiable scalar-valued function $F(\mathbf{x})$ has gradient $\boldsymbol{g}$ and Hessian matrix $H$.
+For the $maximization$ problem, let $\bar{H}$ be any matrix satisfying $\bar{H}\preceq H$ in the Loewner partial ordering.
+We represent these quantities as:
 $\boldsymbol{g}=\left[\begin{array}[]{c}g_{0}\\ g_{1}\\ \vdots\\ g_{d}\\ \end{array}\right],$ $H=\left[\begin{array}[]{cccc}\nabla_{00}^{2}&\nabla_{01}^{2}&\ldots&\nabla_{0d}^{2}\\ \nabla_{10}^{2}&\nabla_{11}^{2}&\ldots&\nabla_{1d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \nabla_{d0}^{2}&\nabla_{d1}^{2}&\ldots&\nabla_{dd}^{2}\\ \end{array}\right],\quad$
@@
 where $\nabla_{ij}^{2}=\nabla_{ji}^{2}=\frac{\partial^{2}F}{\partial x_{i}\partial x_{j}}$.
-We construct a new diagnoal Hessian matrix $\tilde{B}$ with each diagnoal element $\tilde{B}_{kk}$ being $-\epsilon-\sum_{i=0}^{d}|\bar{h}_{ki}|$,
+We construct a new diagonal matrix $\tilde{B}$, referred to as the diagonal Hessian approximation, where each diagonal entry $\tilde{B}_{kk}$ is defined as $-\epsilon-\sum_{i=0}^{d}|\bar{h}_{ki}|$.
+Formally, $\tilde{B}$ is given by:
 $\tilde{B}=\left[\begin{array}[]{ccc}-\epsilon-\sum_{i=0}^{d}|\bar{h}_{0i}|&0&\ldots\\ 0&-\epsilon-\sum_{i=0}^{d}|\bar{h}_{1i}|&\ldots\\ \vdots&\vdots&\ddots\\ 0&0&\ldots\\ \end{array}\right],$
-where $\epsilon$ is a small positive constant to avoid division by zero (usually set to $81e-$ 8).
+where $\epsilon>0$ is a small constant (typically set to $1\text{e}{-8}$) added for numerical stability, such as preventing division by zero.
 As long as $\tilde{B}$ satisfies the convergence condition of the above fixed Hessian method, $\tilde{B}\leq H$, we can use this approximation $\tilde{B}$ of the Hessian matrix as a lower bound.
 Since we already assume that $\bar{H}\leq H$, it will suffice to show that $\tilde{B}\leq\bar{H}$.
-We prove $\tilde{B}\leq\bar{H}$ in a similar way that (Bonte & Vercauteren, 2018) did.
-Lemma 3.1.
+We prove $\tilde{B}\leq\bar{H}$ in a similar way that [2] did.
+Lemma III.1.
 Let $A\in\mathbb{R}^{n\times n}$ be a symmetric matrix, and let $B$ be the diagonal matrix whose diagonal entries $B_{kk}=-\epsilon-\sum_{i}|A_{ki}|$ for $k=1,\ldots,n$, then $B\leq A$.
@@
 ∎
-Definition 3.2 (Quadratic Gradient).
+Definition III.2 (Quadratic Gradient).
 Given the previously defined $\tilde{B}$, we define the quadratic gradient as $G=\bar{B}\cdot\boldsymbol{g}$ with a new learning rate $N_{t}$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$.
@@
 It is worth noting that $\bar{H}$ can be taken as the Hessian $H$ itself, thereby eliminating the need to find either lower or upper bounds.
-Böhning and Lindsay (1988) did not propose a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method.
+Böhning and Lindsay [10] did not propose a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method.
 This may be due to the absence of such matrices for most objective functions.
-In this work, we present a systematic approach for identifying fixed Hessian matrices when they exist.
-Our detection method is based on a necessary condition—if it fails to identify a valid fixed Hessian matrix, it does not rule out the existence of one; instead, it indicates that no such matrix can be constructed through our particular framework.
+In this work, we present a systematic approach for identifying if there were fixed Hessian matrices existing.
+Our proposed method serves as one of many possible approaches to detect the existence of a fixed Hessian matrix.
+Specifically, if our approach identifies a fixed Hessian, this provides sufficient evidence for the existence of such a matrix.
+However, failure to detect a fixed Hessian via our solution does not imply its non-existence.
+In other words, detection by our technique constitutes a sufficient but not necessary condition for the existence of a fixed Hessian matrix.
 We note that our quadratic gradient approach does not require a constant (fixed) Hessian replacement and instead allows for the use of the original Hessian itself to construct our varying diagonal matrix $\tilde{B}$ for the quadratic gradient.
 Under this less restrictive condition, we present a methodical approach to evaluate whether a fixed Hessian matrix can be employed for efficient computation.
-Specifically, we first construct $\tilde{B}$ from the Hessian of the given objective function and then check if, for each diagonal element $\tilde{B}_{kk}$, there is a constant maximum for $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$ (or a constant minimum $|\tilde{B}_{kk}|$).
-If each diagonal element $|\tilde{B}_{kk}|$ achieves its minimum $|\tilde{M}_{kk}|$, then a constant diagonal matrix with the corresponding diagonal elements $|\tilde{M}_{kk}|$ would satisfy the convergence condition for the SFH method.
-Finally, we should evaluate the performance of such a fixed Hessian replacement, as it might not always provide a good bound.
-This search method can be applied to any optimization function to identify a fixed Hessian.
-However, the absence of a constant Hessian replacement via this method does not necessarily imply that such an approximation does not exist.
+Specifically, our method proceeds in the following steps:
+  • Step 1: Construct $\tilde{B}$, an approximation of the Hessian, directly from the Hessian of the given objective function.
+  • Step 2: For each diagonal element $\tilde{B}_{kk}$, examine whether there exists a constant maximum for $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$, or equivalently, a constant minimum of $|\tilde{B}_{kk}|$.
+If each diagonal element $|\tilde{B}_{kk}|$ achieves its minimum $|\tilde{M}_{kk}|$, then a constant diagonal matrix with the corresponding diagonal elements $1/|\tilde{M}_{kk}|$ would satisfy the convergence condition for the SFH method.
+  • Step 3: Evaluate the performance of the resulting fixed Hessian approximation.
+While this replacement satisfies theoretical convergence conditions, it may not always yield a tight bound to the true Hessian, potentially degrading the convergence rate in practice.
+This search method can be applied to obtain the same fixed Hessian matrix $-\frac{1}{4}X^{\top}X$ as originally proposed by the authors in [10], in the context of binary logistic regression.
 Algorithm 1 The Enhanced Nesterov’s Accelerated Gradient Algorithm
@@
 47:  Return $V$
-3.3 Quadratic Gradient Algorithms
+III-C Quadratic Gradient: Algorithms
 Quadratic gradient can be used to enhance various first-order gradient algorithms:
@@
     $\boldsymbol{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
-    where $V_{t+1}$ is the intermediate variable used for updating the final weight $\boldsymbol{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position (Kim et al., 2018a).
-The enhanced NAG is to replace (3) with $V_{t+1}=\boldsymbol{\beta}_{t}+N_{t}\cdot G,$ where $N_{t}$ is the new learning rate for the enhanced NAG algorithm usually set to $1+\eta_{t}$.
-Our enhanced NAG is described in Algorithm 1 .
+    where $V_{t+1}$ is the intermediate variable used for updating the final weight $\boldsymbol{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position [3].
+The enhanced NAG algorithm modifies the original update rule (3) by replacing it with
+    $V_{t+1}=\boldsymbol{\beta}_{t}+N_{t}\cdot B^{-1}\cdot\nabla J(\boldsymbol{\beta}_{t})=\boldsymbol{\beta}_{t}+N_{t}\cdot G,$
+    where $N_{t}$ denotes a learning rate adapted specifically for the enhanced variant.
+As detailed in Algorithm 1, we typically set $N_{t}=1+\eta_{t}$, where $\eta_{t}$ follows a decay schedule or fixed setting depending on the application.
   2. 2.
@@
     $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{\eta_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}\boldsymbol{g}_{[i]}^{(t)}\cdot\boldsymbol{g}_{[i]}^{(t)}}}\cdot\boldsymbol{g}_{[i]}^{(t)},$
-    $and$
     $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{N_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
@@
 (h) The private financial dataset
-Figure 1: The training results of NAG and Enhanced NAG in the clear
+Figure 1: The training results of NAG vs. Enhanced NAG in the clear domain.
 (a) The iDASH dataset
@@
 (h) The private financial dataset
-Figure 2: The training results of Adagrad and Enhanced Adagrad in the clear
+Figure 2: The training results of Adagrad vs. Enhanced Adagrad in the clear domain.
 (a) The iDASH dataset
@@
 (h) The private financial dataset
-Figure 3: The training results of Adam and Enhanced Adam in the clear
+Figure 3: The training results of Adam vs. Enhanced Adam in the clear domain.
 Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM.
 Given that our focus is on the convergence speed of the algorithms during training, we use the loss function $l(\boldsymbol{\beta})$, specifically maximum likelihood estimation (MLE), as the only indicator.
-We evaluate six algorithms—NAG, Adagrad, Adam, and their quadratic-gradient variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam, respectively)—using the datasets adopted by Kim et al.
-(2018a): the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (edin), the Low Birth Weight Study (lbw), Nhanes III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study datasets (uis).
+We conduct a comprehensive evaluation of six optimization algorithms—namely, NAG, Adagrad, Adam, and their respective quadratic-gradient enhanced variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam)—on the benchmark datasets utilized by [3], including the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (edin), the Low Birth Weight Study (lbw), NHANES III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study (uis).
 The genomic dataset from the third task of the 2017 iDASH competition comprises 1,579 records, each featuring 103 binary genotypes and a binary phenotype indicating whether the patient has cancer.
 The remaining five datasets each contain a single binary dependent variable.
-We also evaluate these algorithms on two large datasets from Han et al.
-(2019): a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features.
-For a fair comparison with the baseline work by Kim et al.
-(2018a), the enhanced NAG algorithm employs the learning rate $1+\frac{10}{1+t}$, consistent with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate.
-The enhanced Adagrad algorithm utilizes $N_{t}=0.06+0.01$ as the learning rate, while the enhanced Adam algorithm adopts the following learning rate settings: $\alpha=0.01+0.001$ (the new learning rate), $\beta_{1}=0.9$, and $\beta_{2}=0.999$.
-Refer to (Kingma & Ba, 2014) for further details regarding these parameters.
+We also evaluate these algorithms on two large datasets from [14]: a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features.
+For a fair comparison with the baseline work by [3], the enhanced NAG algorithm employs the learning rate $1+\frac{10}{1+t}$, consistent with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate.
+The Enhanced Adagrad algorithm employs a modified learning rate of $N_{t}=0.06+0.01$, in contrast to the default setting $N_{t}=0.01$ used by the original Adagrad algorithm.
+Similarly, the Enhanced Adam algorithm adopts an adjusted learning rate configuration with $\alpha=0.01+0.001$ (the new learning rate), $\beta_{1}=0.9$, and $\beta_{2}=0.999$, whereas the standard Adam algorithm follows its default settings of $\alpha=0.001$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$.
+For further details regarding these parameters, refer to [17].
 In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\boldsymbol{\beta})$.
@@
 When the learning rate exceeds $3$, the algorithm is nearly guaranteed not to converge.
-For quadratic gradient algorithms, we recommend employing an exponentially decaying learning rate, such as $1+A\cdot\gamma^{t}$, where $t$ denotes the iteration number, $A$ is a positive constant typically set to no less than $1$, and $\gamma$ is a positive number less than $1$ that controls the rate of decay.
-The reasons why the enhanced algorithms have a better performance might be that quadratic gradient can incorporate curves into first-order gradient algorithms.
-(a) The raw first-order gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{lr}\cdot\boldsymbol{g}$
-(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{LR}\cdot G$
-Figure 4: The training outcomes for both the raw gradient algorithm and the raw quadratic gradient ascent algorithm in the clear setting, conducted on the lbw dataset
+For quadratic-gradient NAG algorithms, we recommend employing an exponentially decaying learning rate, such as one of the form $1+A\cdot\gamma^{t}$, where $t$ denotes the iteration index, $A$ is a positive constant typically set to no less than $1$, and $\gamma$ is a positive scalar less than $1$ that governs the rate of decay.
+(a) First-order gradient ascent: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{lr}\cdot\boldsymbol{g}$
+(b) Quadratic gradient ascent: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{LR}\cdot G$
+Figure 4: Training results of standard first-order and quadratic gradient ascent algorithms on the lbw dataset in the plaintext domain.
 Gradient And Quadratic Gradient We executed the raw gradient ascent algorithm and raw quadratic gradient algorithm on the lbw dataset using various learning rates.
@@
 It is precisely these results that inspired the present authors to come up with the idea of quadratic gradient.
-Similar to the gradient, the quadratic gradient carrying the curve also exhibits the characteristic of smooth progression.
+Similar to the gradient, the quadratic gradient also exhibits the characteristic of smooth progression.
 That is, when the learning rate is gradually increased, both the gradient and the quadratic gradient will show corresponding gradual changes in performance rather than abrupt shifts.
-4 Secure Training
-Adagrad method is not a practical solution for homomorphic LR due to its frequent inversion operations.
+IV Secure Training
+Both Adagrad and Adam algorithms are not practical solutions for homomorphic logistic regression due to their frequent reliance on inversion operations.
 It seems plausible that the enhanced NAG is likely the reasonable choice for homomorphic LR training.
@@
 We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al.
-(2018a), utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
+[3], utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
 Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al.
-(2018a) used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$.
-The weight vector $\beta^{(0)}$ consisting of zeros and the diagnoal elements of $\bar{B}$ are copied $n$ times to form two matrices.
+[3] used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$.
+The initial weight vector $\beta^{(0)}$, initialized as a zero vector, and the diagonal elements of the matrix $\bar{B}$ are each replicated $n$ times to construct two corresponding matrices.
 The data owner then encrypt the two matrices into two ciphertexts $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$, respectively.
@@
 where $\bar{B}_{[i][i]}$ is the diagonal element of $\bar{B}$ that is built from $-\frac{1}{4}X^{\top}X$.
-The pulbic cloud takes the three ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ and evaluates the enhanced NAG algorithm to find a decent weight vector by updating the vector $\text{ct}_{\beta}^{(0)}$.
-Refer to (Kim et al., 2018a) for a detailed description about how to calculate the gradient by HE programming.
+The public cloud receives the ciphertexts $\text{ct}Z$, $\text{ct}{\beta}^{(0)}$, and $\text{ct}{\bar{B}}$, and executes the enhanced NAG algorithm to iteratively refine the encrypted weight vector by updating $\text{ct}{\beta}^{(0)}$ toward a more optimal solution.
+Refer to [3] for a detailed description about how to calculate the gradient by HE programming.
 Limitations
-In a privacy-preserving setting, when compared to the NAG method, the primary limitation of the Enhanced NAG method is that it requires one additional ciphertext multiplication to construct the quadratic gradient.
+In the privacy-preserving setting, when compared to the NAG method, the primary limitation of the Enhanced NAG method is that it requires one additional ciphertext multiplication to construct the quadratic gradient.
 In addition, the data owner needs to upload one more ciphertext encrypting the matrix $\bar{B}$.
 However, the enhanced algorithm converges faster, and we believe it can compensate for the mentioned drawbacks.
-Table 1: Implementation Results for iDASH datasets with 10-fold CV
+TABLE I: Implementation Results for iDASH datasets with 10-fold CV
 Dataset
@@
 iDASH 1579 18 Ours 5 4 0.08 4.43 61.46 0.696
-(Kim et al., 2018a) 5 7 0.04 6.07 62.87 0.689
-Table 2: Implementation Results for other datasets with 5-fold CV
+[3] 5 7 0.04 6.07 62.87 0.689
+TABLE II: Implementation Results for other datasets with 5-fold CV
 Dataset
@@
 Edinburgh 1253 9 Ours 5 4 0.04 0.6 89.52 0.943
-(Kim et al., 2018a) 5 7 0.02 3.6 91.04 0.958
+[3] 5 7 0.02 3.6 91.04 0.958
 lbw 189 9 Ours 5 4 0.04 0.6 71.35 0.667
-(Kim et al., 2018a) 5 7 0.02 3.3 69.19 0.689
+[3] 5 7 0.02 3.3 69.19 0.689
 nhanes3 15649 15 Ours 5 4 0.31 4.5 79.23 0.637
-(Kim et al., 2018a) 5 7 0.16 7.3 79.22 0.717
+[3] 5 7 0.16 7.3 79.22 0.717
 pcs 379 9 Ours 5 4 0.04 0.6 63.20 0.733
-(Kim et al., 2018a) 5 7 0.02 3.5 68.27 0.740
+[3] 5 7 0.02 3.5 68.27 0.740
 uis 575 8 Ours 5 4 0.04 0.6 74.43 0.597
-(Kim et al., 2018a) 5 7 0.02 3.5 74.44 0.603
-5 Experiments
-Implementation
-We implement the enhanced NAG based on HE with the library HEAAN.
+[3] 5 7 0.02 3.5 74.44 0.603
+V Experiments
+We implement the enhanced NAG algorithm over homomorphically encrypted data using the HEAAN library.
 The C++ source code is publicly available at https://anonymous.4open.science/r/IDASH2017-245B.
-All the experiments on the ciphertexts were conducted on a public cloud with 32 vCPUs and 64 GB RAM.
-For a fair comparison with (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets.
-Like (Kim et al., 2018a), We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators.
-Tables 1 and 2 display the results of the two experiments, respectively.
-The two tables also provide the average evaluation running time and the storage (encrypted dataset for the baseline work and encrypted dataset and $\bar{B}$ for our method).
-We adopt the same packing method that Kim et al.
-(2018a) proposed and hence our solution has similar storage of ciphertexts to (Kim et al., 2018a) with some extra ciphertexts to encrypt the $\bar{B}$.
-We chose $1+0.9^{t}$ as our learning rate configuration.
-The parameters of HEAAN we set are same to (Kim et al., 2018a): $logN=16$, $logQ=1200$, $logp=30$, $slots=32768$, which ensure the security level $\lambda=80$.
+All experiments on encrypted data were conducted on a public cloud instance equipped with 32 vCPUs and 64 GB of RAM.
+For a fair comparison with the baseline in [3], we adopt the same 10-fold cross-validation (CV) technique on the iDASH dataset, which contains 1,579 samples with 18 features, and apply the same 5-fold CV setup to the remaining five datasets.
+Following [3], we report average accuracy and Area Under the Curve (AUC) as the primary evaluation metrics.
+The results of these experiments are summarized in Tables I and II, which also include average evaluation runtime and storage consumption (i.e., encrypted dataset for the baseline, and encrypted dataset plus $\bar{B}$ for our method).
+To ensure consistency in ciphertext storage, we adopt the same packing strategy proposed by Kim et al.
+[3], resulting in comparable ciphertext sizes, with the only addition being the ciphertexts required to encode $\bar{B}$.
+We use a learning rate of $1+0.9^{t}$ throughout all training procedures.
+The parameters for HEAAN are configured identically to those in [3], with $\log N=16$, $\log Q=1200$, $\log p=30$, and the number of slots set to $32{,}768$, yielding a security level of $\lambda=80$.
 We use a larger $logp=40$ to encrypt the matrix ${\bar{B}}$ for preserving the precision of ${\bar{B}}$.
-Refer to (Kim et al., 2018a) for details on these parameters.
-Since our enhanced NAG method need one more ciphertext multiplication than the baseline work, consuming more modulus, our solution thus can only perform $4$ iterations of the enhanced NAG method.
-Yet despite only $4$ iterations, our enhanced NAG method still produces a comparable result.
-Our learning time measurement reflects the time required for 4 iterations of our enhanced algorithms while the baseline’s reported average running time includes time for evaluation and encryption.
+Refer to [3] for details on these parameters.
+Since our enhanced NAG method requires one additional ciphertext multiplication compared to the baseline approach—resulting in greater modulus consumption—our scheme is limited to performing only four iterations of the enhanced NAG algorithm.
+Nevertheless, even with just four iterations, the proposed method achieves results that remain comparable to those of the baseline.
+Our learning time measurement reflects the time required for 4 iterations of our enhanced algorithms while the baseline’s reported average running time includes time for encryption and evaluation.
 It is worth noting that we relied on their open-source code to implement our homomorphic logistic regression training.
 Therefore, theoretically, the running time for one iteration of our algorithm should be very similar to theirs.
-6 Conclusion
+VI Conclusion
 In this paper, we introduce a novel gradient variant, referred to as the quadratic gradient, and implement a quadratic-gradient-enhanced version of Nesterov’s Accelerated Gradient (NAG) to train a logistic regression model in the encrypted domain.
-The proposed quadratic gradient is constructed directly from the Hessian matrix, effectively bridging first-order gradient-based methods with second-order Newton-type techniques.
+The proposed quadratic gradient is preferably constructed directly from the Hessian matrix, effectively bridging first-order gradient-based methods with second-order Newton-type techniques.
 This approach exhibits strong potential for accelerating a broad range of optimization algorithms.
 In particular, it offers a viable alternative to traditional line-search procedures, can be incorporated into adaptive methods such as enhanced AdaGrad, and may serve to accelerate Newton’s method via embedded gradient descent steps.
-Impact Statements
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+As future work, we aim to establish a more rigorous theoretical analysis of the convergence properties of the enhanced algorithm.
+A more practical and application-driven direction involves systematically evaluating the performance of the enhanced method under different learning rate configurations.

### v14

@@
 Training logistic regression over encrypted data has been a compelling approach in addressing security concerns for several years.
-In this paper, we introduce an efficient gradient variant, called $quadratic$ $gradient$, which can be used for privacy-preserving logistic regression training.
+In this paper, we introduce an efficient gradient variant, called quadraticquadratic gradientgradient, which can be used for privacy-preserving logistic regression training.
 We enhance Nesterov’s Accelerated Gradient (NAG), Adaptive Gradient Algorithm (Adagrad) and Adam algorithms by incorporating their quadratic gradients and evaluate these improved algorithms on various datasets.
@@
 Moreover, we applied the enhanced NAG method to implement homomorphic logistic regression training, achieving comparable results within just 4 iterations.
-There is a great chance that the quadratic gradient approach could integrate first-order gradient descent/ascent algorithms with the second-order Newton-Raphson methods, and that quadratic gradient could be applied to a wide range of numerical optimization problems.
-Index Terms:
-Homomorphic Encryption, Logistic Regression, Quadratic Gradient, Simplified Fixed Hessian, Nesterov’s Accelerated Gradient, Gradient Ascent Algorithm
+The proposed quadratic gradient approach may offer a unified framework that combines the advantages of first-order gradient descent/ascent algorithms and second-order Newton–Raphson methods, potentially extending its applicability to diverse numerical optimization tasks.
 I Introduction
@@
   4. 4.
-    We propose a systematic framework to determine the existence of constant Hessian approximations in fixed-Hessian methods [10, 2], and to assess the feasibility of employing fixed Hessian matrices for efficient computation.
+    We introduce a systematic framework that not only identifies the existence of constant Hessian approximations in fixed-Hessian methods [10, 2], but also evaluates the practicality of employing fixed Hessian matrices for computational efficiency.
 II Preliminaries
-We use the square brackets “[][\ ][ ]” to denote the index of a vector or matrix element in what follows.
+We use the square brackets “[][\ ]” to denote the index of a vector or matrix element in what follows.
 For example, for a vector $\boldsymbol{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\boldsymbol{v}[i]$ or $\boldsymbol{v}_{[i]}$ means the $i$-th element of vector $\boldsymbol{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
@@
 Cheon et al.
-[13] proposed a method to construct an HE scheme with a rescaling procedure which could eliminate this technical bottleneck effectively.
+[13] proposed a method to construct an HE scheme with a rescalingrescaling procedure which could eliminate this technical bottleneck effectively.
 We adopt their open-source implementation HEAAN while implementing our homomorphic LR algorithms.
@@
 $=X^{\top}SX,$
-where $S$ is a diagonal matrix with entries $S_{ii}=(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})$ and $X$ the dataset.
+where $S$ is a diagonal matrix with entries $S_{[i][i]}=(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})$ and $X$ the dataset.
 The log-likelihood function $l(\boldsymbol{\beta})$ of LR has at most a unique global maximum [16], where its gradient is zero.
@@
 where $b_{ii}$ is the reciprocal of $\sum_{i=0}^{d}\bar{h}_{0i}$ and $\nabla_{i}$ is the element of $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$.
-Consider a special situation: if all the elements $b_{00},\ldots,b_{dd}$ had the same value $-\eta$ with $\eta>0$, the iterative formula of the SFH method could be given as:
+Consider a special situation: if all the elements $b_{00},\ldots,b_{dd}$ had the same value $\eta$ with $\eta>0$, the iterative formula of the SFH method could be given as:
 $\boldsymbol{\beta}_{t+1}$ $=\boldsymbol{\beta}_{t}-(-\eta)\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\boldsymbol{\beta}_{t}+\eta\cdot\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}),$
-which is the same as the formula of the naive gradient $ascent$ method.
-Such a coincidence not only helps generate the idea behind this work but also leads us to believe that there is a connection between the Hessian matrix and the learning rate of the gradient (descent) method.
+which is the same as the formula of the naive gradient ascentascent method.
+This coincidence sparked the idea of unifying fixed-Hessian approaches with gradient-based optimization, forming the foundation of this study.
 We introduce $B^{-1}\cdot\nabla_{\boldsymbol{\beta}}\ell(\boldsymbol{\beta})$ as a novel enhanced gradient variant, and allocate a dedicated learning rate to better capture its dynamics.
-As long as we ensure that this new learning rate decreases from a positive floating-point number greater than 1 (such as 2) to 1 in a bounded number of iteration steps, the fixed Hessian Newton’s method guarantees the algorithm will converge eventually.
+If the new learning rate decreases from a positive value greater than one (e.g., 2) to one within a finite number of iterations, the fixed-Hessian Newton’s method theoretically guarantees that our proposed approach with this new learning-rate scheme will eventually converge.
 The SFH method proposed by Bonte and Vercauteren [2] has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive.
@@
 Suppose a differentiable scalar-valued function $F(\mathbf{x})$ has gradient $\boldsymbol{g}$ and Hessian matrix $H$.
-For the $maximization$ problem, let $\bar{H}$ be any matrix satisfying $\bar{H}\preceq H$ in the Loewner partial ordering.
+For the maximizationmaximization problem, let $\bar{H}$ be any matrix satisfying $\bar{H}\leq H$ in the Loewner partial ordering.
 We represent these quantities as:
@@
 where $\nabla_{ij}^{2}=\nabla_{ji}^{2}=\frac{\partial^{2}F}{\partial x_{i}\partial x_{j}}$.
-We construct a new diagonal matrix $\tilde{B}$, referred to as the diagonal Hessian approximation, where each diagonal entry $\tilde{B}_{kk}$ is defined as $-\epsilon-\sum_{i=0}^{d}|\bar{h}_{ki}|$.
+We construct a new diagonal matrix $\tilde{B}$, referred to as the diagonal Hessian approximation, where each diagonal entry $\tilde{B}_{kk}$ is defined as $\epsilon-\sum_{i=0}^{d}|\bar{h}_{ki}|$.
 Formally, $\tilde{B}$ is given by:
@@
 Proof.
-By definition of the Loewner ordering, we have to prove the difference matrix $C=A-B$ is non-negative definite, which means that all the eigenvalues of $C$ need to be non-negative.
+By definition of the Loewner ordering, we have to prove the difference matrix C=ABC=A-B is non-negative definite, which means that all the eigenvalues of $C$ need to be non-negative.
 By construction of $C$ we have that $C_{ij}=A_{ij}+\epsilon+\sum_{k=1}^{n}|A_{ik}|$ for $i=j$ and $C_{ij}=A_{ij}$ for $i\neq j$.
@@
 To maximize the function $F(\mathbf{x})$, we can employ the iterative formula: $\mathbf{x}_{t+1}=\mathbf{x}_{t}+N_{t}\cdot G$, similar to the naive gradient approach.
-Minimizing the function $F(x)$ is equivalent to maximizing the function $-F(x)$.
-In this context, we need to construct $\tilde{B}$ using either a good $lower$ bound $\bar{H}$ for the Hessian $-H$ of $-F(x)$ or a good $upper$ bound $\bar{H}$ for the Hessian $H$ of $F(x)$.
+Minimizing the function $F(x)$ is equivalent to maximizing the function F(x)-F(x).
+In this context, we need to construct $\tilde{B}$ using either a good lowerlower bound $\bar{H}$ for the Hessian H-H of F(x)-F(x) or a good upperupper bound $\bar{H}$ for the Hessian $H$ of $F(x)$.
 It is worth noting that $\bar{H}$ can be taken as the Hessian $H$ itself, thereby eliminating the need to find either lower or upper bounds.
-Böhning and Lindsay [10] did not propose a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method.
+Böhning and Lindsay [10] did not provide a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method.
 This may be due to the absence of such matrices for most objective functions.
-In this work, we present a systematic approach for identifying if there were fixed Hessian matrices existing.
+In this work, we introduce a structured procedure for identifying if there were fixed Hessian matrices existing.
 Our proposed method serves as one of many possible approaches to detect the existence of a fixed Hessian matrix.
@@
 Under this less restrictive condition, we present a methodical approach to evaluate whether a fixed Hessian matrix can be employed for efficient computation.
-Specifically, our method proceeds in the following steps:
-  • Step 1: Construct $\tilde{B}$, an approximation of the Hessian, directly from the Hessian of the given objective function.
+Our method proceeds in the following steps:
+  • Step 1: Construct $\tilde{B}$, an approximation of the Hessian, directlydirectly from the Hessian of the given objective function.
   • Step 2: For each diagonal element $\tilde{B}_{kk}$, examine whether there exists a constant maximum for $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$, or equivalently, a constant minimum of $|\tilde{B}_{kk}|$.
 If each diagonal element $|\tilde{B}_{kk}|$ achieves its minimum $|\tilde{M}_{kk}|$, then a constant diagonal matrix with the corresponding diagonal elements $1/|\tilde{M}_{kk}|$ would satisfy the convergence condition for the SFH method.
-  • Step 3: Evaluate the performance of the resulting fixed Hessian approximation.
+  • Step 3: Evaluate the performance of the resulting fixed Hessian approximation derived in the previous step.
 While this replacement satisfies theoretical convergence conditions, it may not always yield a tight bound to the true Hessian, potentially degrading the convergence rate in practice.
-This search method can be applied to obtain the same fixed Hessian matrix $-\frac{1}{4}X^{\top}X$ as originally proposed by the authors in [10], in the context of binary logistic regression.
+This search method can be applied to obtain the same fixed Hessian matrix $\frac{1}{4}X^{\top}X$ as originally proposed by the authors in [10], in the context of binary logistic regression.
 Algorithm 1 The Enhanced Nesterov’s Accelerated Gradient Algorithm
@@
 2:Ensure: the parameter vector $V\in\mathbb{R}^{(1+d)}$
-3:  Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$ $\bar{H}\in\mathbb{R}^{(1+d)\times(1+d)}$
-4:  Set $V\leftarrow\boldsymbol{0}$, $W\leftarrow\boldsymbol{0}$, $\bar{B}\leftarrow\boldsymbol{0}$ $V\in\mathbb{R}^{(1+d)}$, $W\in\mathbb{R}^{(1+d)}$, $\bar{B}\in\mathbb{R}^{(1+d)\times(1+d)}$
+3:Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$ # $\bar{H}\in\mathbb{R}^{(1+d)\times(1+d)}$
+4:Set $V\leftarrow\boldsymbol{0}$, $W\leftarrow\boldsymbol{0}$, $\bar{B}\leftarrow\boldsymbol{0}$ # $V\in\mathbb{R}^{(1+d)}$, $W\in\mathbb{R}^{(1+d)}$, $\bar{B}\in\mathbb{R}^{(1+d)\times(1+d)}$
 5:for $i:=0$ to $d$ do
-6:     # $\epsilon$ is a small positive constant such as $81e-$ 8
+6:  # $\epsilon$ is a small positive constant such as 1e81e-8
 7:  $\bar{B}[i][i]\leftarrow\epsilon$
@@
   1. 1.
-    NAG is a different variant of the momentum method to give the momentum term much more prescience. The iterative formulas of the gradient $ascent$ method for NAG are as follows:
+    NAG is a different variant of the momentum method to give the momentum term much more prescience. The iterative formulas of the gradient ascentascent method for NAG are as follows:
     $V_{t+1}$ $=\boldsymbol{\beta}_{t}+\eta_{t}\cdot\nabla J(\boldsymbol{\beta}_{t}),$ (3)
@@
 Similarly, the Enhanced Adam algorithm adopts an adjusted learning rate configuration with $\alpha=0.01+0.001$ (the new learning rate), $\beta_{1}=0.9$, and $\beta_{2}=0.999$, whereas the standard Adam algorithm follows its default settings of $\alpha=0.001$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$.
-For further details regarding these parameters, refer to [17].
+For further details regarding these parameters, see [17].
 In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\boldsymbol{\beta})$.
 From the empirical results presented in Figures 21, and 3, we conclude that all the enhanced algorithms converge and consistently outperforms the original NAG method in terms of convergence rate.
-An important observation is that the enhanced algorithms demonstrate better performance with learning rates between $1$ and $2$ compared to other values.
-When the learning rate exceeds $3$, the algorithm is nearly guaranteed not to converge.
-For quadratic-gradient NAG algorithms, we recommend employing an exponentially decaying learning rate, such as one of the form $1+A\cdot\gamma^{t}$, where $t$ denotes the iteration index, $A$ is a positive constant typically set to no less than $1$, and $\gamma$ is a positive scalar less than $1$ that governs the rate of decay.
+An important observation is that, under a fixed learning rate, the enhanced algorithms achieve their better performance when the learning rate is between $1$ and $2$.
+By contrast, when the learning rate exceeds $3$, the algorithm almost never converges.
+For quadratic-gradient NAG algorithms, we recommend employing an exponentially decaying learning rate.
 (a) First-order gradient ascent: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{lr}\cdot\boldsymbol{g}$
@@
 We leave the computation of matrix $\bar{B}$ to data owner and let the data owner upload the ciphertext encrypting the $\bar{B}$ to the cloud.
-Since data owner has to prepare the dataset and normalize it, it would also be practicable for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
+Given that the data owner prepares and normalizes the dataset, computing $\bar{B}$ on their side is feasible and secure, as it does not reveal any sensitive data.
 Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model.
@@
 We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al.
-[3], utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
+[3], utilizing the least squares approach to approximate the sigmoid function over the interval [8,8][-8,8]: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$ .
 Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al.
@@
 $\text{ct}_{\bar{B}}$ $=Enc\left[\begin{array}[]{cccc}\bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \end{array}\right],$
-where $\bar{B}_{[i][i]}$ is the diagonal element of $\bar{B}$ that is built from $-\frac{1}{4}X^{\top}X$.
+where $\bar{B}_{[i][i]}$ represents the diagonal entry of $\bar{B}$, formed from $\frac{1}{4}X^{\top}X$.
 The public cloud receives the ciphertexts $\text{ct}Z$, $\text{ct}{\beta}^{(0)}$, and $\text{ct}{\bar{B}}$, and executes the enhanced NAG algorithm to iteratively refine the encrypted weight vector by updating $\text{ct}{\beta}^{(0)}$ toward a more optimal solution.
@@
 We use a learning rate of $1+0.9^{t}$ throughout all training procedures.
-The parameters for HEAAN are configured identically to those in [3], with $\log N=16$, $\log Q=1200$, $\log p=30$, and the number of slots set to $32{,}768$, yielding a security level of $\lambda=80$.
+The parameters for HEAAN are configured identically to those in [3], with $\log N=16$, $\log Q=1200$, $\log p=30$, and the number of slots set to 32,76832{,}768, yielding a security level of $\lambda=80$.
 We use a larger $logp=40$ to encrypt the matrix $\bar{B}$} for preserving the precision of $\bar{B}$}.
-Refer to [3] for details on these parameters.
+These parameters are described in greater detail in [3].
 Since our enhanced NAG method requires one additional ciphertext multiplication compared to the baseline approach—resulting in greater modulus consumption—our scheme is limited to performing only four iterations of the enhanced NAG algorithm.
@@
 VI Conclusion
-In this paper, we introduce a novel gradient variant, referred to as the quadratic gradient, and implement a quadratic-gradient-enhanced version of Nesterov’s Accelerated Gradient (NAG) to train a logistic regression model in the encrypted domain.
-The proposed quadratic gradient is preferably constructed directly from the Hessian matrix, effectively bridging first-order gradient-based methods with second-order Newton-type techniques.
+In this paper, we proposed a novel gradient variant, termed the quadratic gradient, and developed a quadratic-gradient-enhanced NAG for training logistic regression models in the encrypted domain.
+The proposed quadratic gradient could be directlydirectly constructed from the Hessian matrix, effectively bridging first-order gradient-based methods with second-order Newton-type techniques.
 This approach exhibits strong potential for accelerating a broad range of optimization algorithms.
-In particular, it offers a viable alternative to traditional line-search procedures, can be incorporated into adaptive methods such as enhanced AdaGrad, and may serve to accelerate Newton’s method via embedded gradient descent steps.
-As future work, we aim to establish a more rigorous theoretical analysis of the convergence properties of the enhanced algorithm.
-A more practical and application-driven direction involves systematically evaluating the performance of the enhanced method under different learning rate configurations.
+In particular, it offers a viable alternative to traditional line-search procedures and constitutes a game-changing approach to accelerating Newton’s method via embedded gradient-descent steps.
+Open directions for future work include conducting a rigorous theoretical analysis of the enhanced algorithm’s convergence behavior and performing an application-oriented evaluation across diverse learning rate configurations.